### v1

# High-dimensional Autoregressive Modeling for Time Series with Hierarchical Structures

Abstract
High-dimensional time series often exhibit hierarchical structures represented by tensors, while statistical methodologies that can effectively exploit the structural information remain limited. We propose a supervised factor modeling framework that accommodates general hierarchical structures by extracting low-dimensional features sequentially in the mode orders that respect the hierarchical structure. Our method can select a small collection of such orders to allow for impurities in the hierarchical structures, yielding interpretable loading matrices that preserve the hierarchical relationships. A practical estimation procedure is proposed, with a hyperparameter selection scheme that identifies a parsimonious set of action orders and interim ranks, thereby revealing the possibly latent hierarchical structures. Theoretically, non‑asymptotic error bounds are derived for the proposed estimators in both regression and autoregressive settings. An application to the IPIP‑NEO‑120 personality panel illustrates superior forecasting performance and clearer structural interpretation compared with existing methods based on tensor decompositions and hierarchical factor analysis.
Keywords: Autoregressive model, Hierarchical tensor data, High-dimensional time series, Non-asymptotic properties, Supervised factor modeling
1 Introduction
High-dimensional time series abound in modern scientific and econometric studies (wang2019factor; chang2024autocovariance; chen2025estimating), which often exhibit multiway (chen2022analysis; chen2022factor; wang2024high) and partially hierarchical organizations (johnson2014measuring; gao2023divide). The data are hierarchical when they contain multiple modes representing the layers of a nested structure obtained from several groups, subgroups, and items. In many applications, however, the hierarchical structure is only partial as some modes lie outside this hierarchy or even form another hierarchy. For balanced data which admits a tensor representation, methods that can adapt to such partially hierarchical tensor time series remain limited, as existing tensor-based methods or high-dimensional dynamic factor models typically assume a fully juxtaposing or hierarchical organization (montgomery2006design; stock2011dynamic; brunner2012tutorial; yuan2023two). In practice, real datasets frequently combine both: hierarchical components accompanied by cross-cutting covariates. Developing a framework that can efficiently summarize, predict, and statistically interpret such partially hierarchical structures with complex inter-mode dependencies is therefore of both theoretical and practical importance.
A representative example is the International Personality Item Pool (IPIP) NEO-120 personality inventory (johnson2014measuring; qiao2025exploratory), which contains 120 questions assessing five broad personality traits, each with 24 questions. The data are collected over time from participants of different ages and genders, resulting in a three-way tensor time series with dimensions $12\penalty 10000\ (gender\penalty 10000\ and\penalty 10000\ age\penalty 10000\ groups)\times 5\penalty 10000\ (personalities)\times 24\penalty 10000\ (questions)$. The last two dimensions reflect a hierarchical organization, while the first does not belong to this hierarchy. Given the dataset’s scale ($12\times 5\times 24=1440$ dimensions) and the limited number of time points (approximately 100 monthly observations), there is a need for an efficient and statistically interpretable model that can extract latent features from the data for prediction of future observations. The extraction process which reduces the dimensionality in a supervised manner should respect the data’s partially hierarchical structure, when both latent feature detection and trend analysis are of sociological interest.
Natural strategies to handle data of such high dimensionality include imposing sparsity or low-rankness on the model, so that a more tractable low-dimensional structure emerges. Sparsity-inducing methods (basu2015regularized; raskutti2019convex; bertsimas2020sparse) have been widely used in high-dimensional time series analysis with its well-known prediction ability (wilms2023sparse; zheng2025interpretable). However, these methods may lack structural interpretability unless the underlying sparsity pattern aligns with a meaningful structure, as is common in genetics and neuroscience (zhou2013tensor; bi2022modeling; wu2023sparse). In IPIP-NEO-120, for instance, the responses within each personality trait are highly correlated due to the experimental design, and a sparse model may fail to capture this inherent structure.
In contrast, low-rank modeling compresses the data to capture latent structure via orthogonal loading matrices, leading to a more interpretable framework of factor modeling and analysis. Supervised Factor Modeling (SFM) (huang2025supervised) is a prominent low-rank approach that uses loading matrices to extract features (or factors) from both predictors and responses, followed by a multivariate regression on these features. Specifically for hierarchical data, literature is copious with methods that extract features tailored to the hierarchical construct, such as higher-order factor models (yung1999relationship), group factor models (klami2014group), hierarchical factor models (schmid1957development; qiao2025exact). These methods either extracts factors recursively along the hierarchy, or directly from each of its levels. However, they all assign different loading matrices to different lower-level modes under the same higher-level mode, so the number of parameters exceed the dimensions of the data. These approaches are thus only feasible in the unsupervised context, where the proportion of explained variance, rather than prediction performance, is the primary concern. When otherwise applied in the supervised context, only low dimensional settings are allowed.
As traditional low-rank methods for unsupervised factor analysis of hierarchical constructs are unsuitable for high-dimensional estimation problem, SFM methods based on tensor decomposition (huang2025supervised) are preferred given their parsimony. Among these, Tucker and canonical polyadic (CP) decompositions are widely used (kolda2009tensor). Tucker decomposition (tucker1966some) is notable for its numerical stability and interpretable orthogonal loadings, which align well with supervised factor modeling (wang2024high). However, its model complexity grows exponentially with tensor order, limiting its efficiency in high-order settings (oseledets2011tensor). This issue is exacerbated in tensor autoregressive models, where coefficient tensors often have high orders. In the IPIP-NEO-120 example, a lag-one tensor autoregressive model would involve a sixth-order coefficient tensor, and estimation based on Tucker decomposition would be computationally less efficient. Additionally, Tucker loadings are confined on Kronecker product spaces, restricting its applicability to data with complex or hierarchical structures, as discussed in Sections 2.1 and 2.2. CP decomposition (harshman1970foundations) offers a more efficient and flexible alternative (lock2018tensor), with model complexity scaling linearly with tensor order and dimensions. However, CP loadings are complicated, constraining its interpretation in unsupervised settings as a PCA surrogate (han2024cp; han2022tensor), and CP-based SFM (cai2025efficient) lacks clear interpretation for structured data.
Table 1: Comparison of different methods for supervised factor modeling. For simplicity, we assume equal dimension $p$ for each mode of the data, and equal number of features, $r$, extracted from each mode, and denote $d$ as the order of the coefficient tensor.
Models SFM based on Tensor Decomp. Factor Models for SFM
Ours CP Tucker $T$ GFM HFM HOFM
Number of parameters: $O(\cdot)$ $pr^{2}d$ prdprd $prd+r^{d}$ $pr^{2}d$ $rp^{d}$ $rp^{d}$ $rp^{d}$
Interpretation: Hierarchical $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Interpretation: Crossed $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$
Meanwhile, tensor regression based on tensor train (TT) decomposition (oseledets2011tensor) of the coefficient tensor has been proposed with its efficiency highlighted for tensors with high orders (liu2020low). Under the SFM structure, qin2025computational and si2024efficient adopt a pair of TT-styled loadings for the predictor and repsonse, and their sequential structure can capture hierarchical relationships among modes with reduced number of parameters. Specifically, each loading matrix is a product of several component matrices, with each component extracting features from one mode to which the previous mode imparts extracted features. The TT-based model is hence interpretable in the regression tasks where both the predictor and response exhibit a fully hierarchical structure (zhou2022optimal; si2024efficient). However, its strict sequential structure inherited from TT decomposition limits its applicability to data with partially hierarchical structures, as in the IPIP-NEO-120 example. Another major limitation is that the order of the component matrices in TT loadings is fixed to be forward for the predictor and backward for the response. This fixed construction hinders its application to autoregressive time series models, where the predictor and response share basically the same structure. Permuting the order of the predictor’s modes for a forced structural alignment in turn deteriorates the autoregressive model into a general stochastic regression model, complicating the temporal dependence analysis. Table 1 provides a summary of the reviewed methods under the SFM framework, comparing their number of parameters and interpretability for hierarchical and crossed (non-hierarchical) structures.
In view of the aforementioned challenges, we propose a novel tensor-based supervised factor modeling framework that effectively extracts features from partially hierarchical tensor time series, while preserving the hierarchical structure information. Our model combines admissible mode orders of the partially hierarchical structure instrinc to the data, analogous to a mixture model. Then, it extracts features sequentially, thereby capturing complex inter-mode dependencies with cross-cutting covariates obscuring any exact hierarchical structure. The proposed method is computationally efficient, as information summarization and its upward impartation to the higher-level layer bear a shared low-rank structure for the lower-level layer. This parsimony enables both accurate estimation with sample size and statistical interpretability inherited from SFM framework. Specifically, our contributions are summarized as follows:
  • We are, to the best of our knowledge, the first to bridge the methodological gap between fully hierarchical structures and non-hierarchical (juxtaposing) structures. Our proposed novel tensor-based SFM method accommodates to partially hierarchical structures in both regression and autoregression tasks, with theoretical guarantees evidenced by both simulation and empirical examples.
  • Our proposed method offers a flexible and unified way to detect from data principal and lesser hierarchical structures forming a mixture, enabled by our hyperparameter selection algorithm, when the true structure is obscured.
  • The construction in our proposed method can reveal an algebraic-graphical duality of the hierarchical structures, enshrined in an algebraic proposition, which may be of independent interest in the study of tensor decomposition and hierarchical models.
The rest of the paper is organized as follows. In Section 2, we introduce the general supervised factor modeling framework, and then present our proposed method for partially hierarchical data, followed by an interpretable tensor autoregressive model. Section 3 provides the theoretical analysis of our method and Section 4 describes its algorithms along with hyperparameter selection. In Section 5, we conduct extensive simulation studies to evaluate the performance of our method against existing alternatives. In Section 6, we apply our method to the IPIP-NEO-120 dataset for analysis. Finally, we conclude with a discussion in Section 7. All technical proofs and additional numerical results are relegated to the Supplementary Material.
Notations and preliminaries.
Throughout the paper, we denote vectors by boldface small letters, e.g., $\boldsymbol{x}$; matrices by boldface capital letters, e.g., $\boldsymbol{X}$; tensors by boldface Euler capital letters, e.g., $\mathscr{X}$. For a $d$-th order tensor $\mbox{$\mathscr{X}$}\in\mathbb{R}^{p_{1}\times p_{2}\times\cdots\times p_{d}}$, denote by $\operatorname*{vec}(\mbox{$\mathscr{X}$})$, $\mbox{$\mathscr{X}$}_{(s)}$, and $\mbox{$\mathscr{X}$}]_{s}$ its vectorization, mode matricization, and sequential matricization at mode $s$ with $1\leq s\leq d$, respectively. The Frobenius norm of $\mathscr{X}$ is defined as $\left\lVert\mbox{$\mathscr{X}$}\right\rVert_{\mathrm{F}}=\sqrt{\langle\mbox{$\mathscr{X}$},\mbox{$\mathscr{X}$}\rangle}$, where $\langle\cdot,\cdot\rangle$ is the inner product. Moreover, the $(i_{1}i_{2}\ldots i_{d})$-th element of $\mathscr{X}$ is denoted by $\mbox{$\mathscr{X}$}_{i_{1}i_{2}\ldots i_{d}}$, for $i_{k}=1,2,\ldots,p_{k}$, $k=1,2,\ldots,d$. We denote $\times_{s}$ the $s$-th mode matrix product, and $\otimes$ denotes the Kronecker product. For a generic matrix $\boldsymbol{X}$, we denote by $\boldsymbol{X}^{\top}$, $\left\lVert\boldsymbol{X}\right\rVert_{\mathrm{F}}$, $\left\lVert\boldsymbol{X}\right\rVert_{2}$, and $\operatorname*{vec}(\boldsymbol{X})$ its transpose, Frobenius norm, spectral norm, and a long vector obtained by stacking all its columns, respectively. If $\boldsymbol{X}$ is further a square matrix, then denote its minimum and maximum eigenvalue by $\lambda_{\min}(\boldsymbol{X})$ and $\lambda_{\max}(\boldsymbol{X})$, respectively. Denote the orthogonal set as $\mathbb{O}^{p\times r}=\{\boldsymbol{X}\in\mathbb{R}^{p\times r}:\boldsymbol{X}^{\top}\boldsymbol{X}=\boldsymbol{I}_{r}\}$ for $p\geq r$, where $\boldsymbol{I}_{r}\in\mathbb{R}^{r\times r}$ is the identity matrix. With a little abuse of terminology, a rectangular matrix is said to be orthogonal if it belongs to $\mathbb{O}^{p\times r}$. For two real-valued sequences $x_{k}$ and $y_{k}$, $x_{k}\gtrsim y_{k}$ if there exists a $C>0$ such that $x_{k}\geq Cy_{k}$ for all $k$. In addition, we write $x_{k}\asymp y_{k}$ if $x_{k}\gtrsim y_{k}$ and $y_{k}\gtrsim x_{k}$.
2 Model settings
2.1 Tucker tensor regression and supervised factor modeling
We begin with a general regression model involving tensor-valued responses and predictors:
[…]
$\mbox{$\mathscr{A}$}=\mbox{$\mathscr{G}$}\times_{1}\boldsymbol{U}_{1}\times_{2}\cdots\times_{N}\boldsymbol{U}_{N}\times_{N+1}\boldsymbol{V}_{1}\times_{N+2}\cdots\times_{N+M}\boldsymbol{V}_{M},$ (2.2)
where $\mbox{$\mathscr{G}$}\in\mathbb{R}^{s_{1}\times\cdots\times s_{N}\times r_{1}\times\cdots\times r_{M}}$ is the core tensor, and $\boldsymbol{U}_{n}\in\mathbb{R}^{q_{n}\times s_{n}}$ with $1\leq n\leq N$ and $\boldsymbol{V}_{m}\in\mathbb{R}^{p_{m}\times r_{m}}$ with $1\leq m\leq M$ are factor matrices. Model (2.1), together with the low-rank coefficient tensor at (2.2), is referred to as the Tucker tensor regression model.
For simplicity, we denote $q_{N+m}=p_{m}$, $s_{N+m}=r_{m}$ and $\boldsymbol{U}_{N+m}=\boldsymbol{V}_{m}$ for $1\leq m\leq M$. Note that the Tucker decomposition at (2.2) is not unique, since $\mbox{$\mathscr{A}$}=\mbox{$\mathscr{G}$}\times_{1}\boldsymbol{U}_{1}\times_{2}\cdots\times_{N+M}\boldsymbol{U}_{N+M}=\left(\mbox{$\mathscr{G}$}\times_{1}\boldsymbol{O}_{1}\times_{2}\cdots\times_{N+M}\boldsymbol{O}_{N+M}\right)\times_{1}\left(\boldsymbol{U}_{1}\boldsymbol{O}_{1}^{-1}\right)\times_{2}\cdots\times_{N+M}\left(\boldsymbol{U}_{N+M}\boldsymbol{O}_{N+M}^{-1}\right)$ for any invertible matrices $\boldsymbol{O}_{i}\in\mathbb{R}^{s_{i}\times s_{i}}$ with $1\leq i\leq N+M$. Without loss of generality, we restrict these factor matrices to be orthonormal, i.e., $\boldsymbol{U}_{j}\in\mathbb{O}^{q_{j}\times s_{j}}$ for each $1\leq j\leq N+M$; see, for example, the higher-order singular value decomposition (HOSVD) in kolda2009tensor. Consequently, the Tucker tensor regression model can be rewritten into
[…]
Let $\boldsymbol{U}=\boldsymbol{U}_{N}\otimes\dots\otimes\boldsymbol{U}_{1}\in\mathbb{O}^{Q\times s}$ and $\boldsymbol{V}=\boldsymbol{V}_{M}\otimes\cdots\otimes\boldsymbol{V}_{1}\in\mathbb{O}^{P\times r}$, where $Q\vcentcolon=\prod_{n=1}^{N}q_{n}$, $P\vcentcolon=\prod_{m=1}^{M}p_{m}$, $s\vcentcolon=\prod_{n=1}^{N}s_{n}$, and $r\vcentcolon=\prod_{m=1}^{M}r_{m}$. Note that $\boldsymbol{U}_{n}$’s and $\boldsymbol{V}_{m}$’s are still not unique, and hence $\boldsymbol{U}$ and $\boldsymbol{V}$. However, the corresponding column spaces $\text{col}(\boldsymbol{U})$ and $\text{col}(\boldsymbol{V})$, as well as their projection matrices $\boldsymbol{P}_{\boldsymbol{U}}=\boldsymbol{U}\boldsymbol{U}^{\top}$ and $\boldsymbol{P}_{\boldsymbol{V}}=\boldsymbol{V}\boldsymbol{V}^{\top}$, can be uniquely defined.
Note that model (2.3) involves dimension reduction on both the responses and predictors. We first consider responses, and project $\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})$ onto $\text{col}(\boldsymbol{U})$ and its orthogonal complement, i.e., $\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})=\boldsymbol{P}_{\boldsymbol{U}}\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})+\boldsymbol{P}_{\boldsymbol{U}}^{\perp}\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})$ with $\boldsymbol{P}_{\boldsymbol{U}}^{\perp}=\boldsymbol{I}_{Q}-\boldsymbol{U}\boldsymbol{U}^{\top}$. It then holds that
[…]
where all predictable information of $\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})$ is contained in $\text{col}(\boldsymbol{U})$. Moreover,
$\mathrm{cov}\left\{\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t}),\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})|\boldsymbol{P}_{\boldsymbol{V}}\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})\right\}=\boldsymbol{0},$
indicating that $\text{col}(\boldsymbol{V})$ contains all useful information of predictors in terms of predicting responses. As a result, define $\boldsymbol{f}_{t}^{\mathrm{response}}=\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t}\times_{n=1}^{N}\boldsymbol{U}_{n}^{\top})=\boldsymbol{U}^{\top}\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})\in\mathbb{R}^{s}$ and $\boldsymbol{f}_{t}^{\mathrm{predictor}}=\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t}\times_{m=1}^{M}\boldsymbol{V}_{m}^{\top})=\boldsymbol{V}^{\top}\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})\in\mathbb{R}^{r}$ to be the response and predictor factors, respectively, and then the Tucker tensor regression can be well interpreted from a supervised factor modeling perspective (wang2022high; wang2024high; huang2025supervised).
The above supervised factor modeling interpretation also motivates us to introduce a new framework to construct a tensor regression model, especially for data with complex structures beyond the capability of Tucker decomposition, or for non-orthogonal loadings in the feature extraction process such as the CP decomposition (harshman1970foundations). Specifically, we first extract features or factors from responses and predictors by
$\boldsymbol{f}_{t}^{\mathrm{predictor}}=\mathcal{M}(\mbox{$\mathscr{X}$}_{t})\quad\text{and}\quad\boldsymbol{f}_{t}^{\mathrm{response}}=\mathcal{N}(\mbox{$\mathscr{Y}$}_{t}),$ (2.4)
respectively. Here, $\mathcal{M}(\cdot):\mathbb{R}^{p_{1}\times\cdots\times p_{M}}\to\mathbb{R}^{r}$ and $\mathcal{N}(\cdot):\mathbb{R}^{q_{1}\times\cdots\times q_{N}}\to\mathbb{R}^{s}$ are feature-extracting operators, with $r$ and $s$ being the number of extracted predictor and response features, respectively. A specific structure is then assumed to the coefficient tensor $\mathscr{A}$ of model (2.1) such that $\boldsymbol{f}_{t}^{\mathrm{response}}=\boldsymbol{\Theta}\boldsymbol{f}_{t}^{\mathrm{predictor}}+\boldsymbol{e}_{t}$, where $\boldsymbol{\Theta}\in\mathbb{R}^{s\times r}$ and $\boldsymbol{e}_{t}\in\mathbb{R}^{s}$ are the corresponding coefficient matrix and compressed error term, respectively. To this end, if operators $\mathcal{M}$ and $\mathcal{N}$ are linear, i.e., $\mathcal{M}(\mbox{$\mathscr{X}$}_{t})=\boldsymbol{\Lambda}_{x}^{\top}\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})$ and $\mathcal{N}(\mbox{$\mathscr{Y}$}_{t})=\boldsymbol{\Lambda}_{y}^{\top}\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})$, with non-degenerating loading matrices $\boldsymbol{\Lambda}_{x}\in\mathbb{R}^{P\times r}$ and $\boldsymbol{\Lambda}_{y}\in\mathbb{R}^{Q\times s}$, we then can set the coefficient tensor to
$\mbox{$\mathscr{A}$}]_{N}=\boldsymbol{\Lambda}_{y}\left(\boldsymbol{\Lambda}_{y}^{\top}\boldsymbol{\Lambda}_{y}\right)^{-1}\boldsymbol{\Theta}\boldsymbol{\Lambda}_{x}^{\top}.$ (2.5)
Moreover, if $\boldsymbol{\Lambda}_{y}$ is orthonormal, then $\boldsymbol{\Lambda}_{y}^{\top}\boldsymbol{\Lambda}_{y}=\boldsymbol{I}_{s}$ and $\mbox{$\mathscr{A}$}]_{N}=\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}\boldsymbol{\Lambda}_{x}^{\top}$.
When Tucker decomposition is applied as in (2.2), the loading matrices $\boldsymbol{\Lambda}_{y}=\boldsymbol{U}$ and $\boldsymbol{\Lambda}_{x}=\boldsymbol{V}$ extract features simultaneously from all modes of the response and predictor, respectively. Such simultaneous compression may not be suitable for more complex data, say with hierarchical structures in the next subsection, and a new methodology is required.
2.2 Purely hierarchical structure
According to technological advances, the available multi-dimensional or tensor-valued data have become more and more complex, and it is very common to observe hierarchical relationships among some modes. In order to exploit their structural information, we attempt to design a regression model using the proposed framework at (2.4) and (2.5). This subsection starts from the simple case with purely hierarchical responses and predictors.
Figure 1: Three hierarchical feature extraction procedures for a second-order tensor $\mbox{$\mathscr{X}$}_{t}\in\mathbb{R}^{p_{1}\times p_{2}}$: hierarchical factor model (HFM), higher-order factor model (HOFM) and our model.
We first consider the feature extraction on predictors, and assume for simplicity that $M=2$, i.e., $\mbox{$\mathscr{X}$}_{t}\in\mathbb{R}^{p_{1}\times p_{2}}$, and the $p_{1}$ variables of the first mode are nested within each of the $p_{2}$ variables in its second mode. There are two commonly used methods in the literature of factor analysis to extract features from such hierarchical data $\mbox{$\mathscr{X}$}_{t}$: the hierarchical factor model (HFM) in schmid1957development and higher-order factor model (HOFM) in yung1999relationship. The HFM directly extracts features from each level of the hierarchy separately and jointly. In contrast, the HOFM performs feature extraction in a sequential manner along the hierarchy, where features extracted from a lower level are used to summarize those from the upper level. The loading matrices for these two methods are illustrated in Figure 1. Note that the HFM has more than $p_{2}$ features, while the HOFM has $O(rP)$ parameters with $P=p_{1}p_{2}$, making both of them infeasible in high-dimensional settings. Moreover, Tucker decomposition method extracts features simultaneously from all modes, i.e. $\mathcal{M}(\mbox{$\mathscr{X}$}_{t})=\operatorname*{vec}(\boldsymbol{V}_{1}^{\top}\mbox{$\mathscr{X}$}_{t}\boldsymbol{V}_{2})$, and it is not suitable to handle such hierarchical relationships.
Note that the HOFM employs different loadings for each level of the hierarchy. Such flexibility can maintain its performance, while the corresponding number of parameters is also enlarged dramatically for a larger dimension of $M$. This paper alternatively proposes a sequential feature extraction method with shared loadings within each level of the hierarchy. Specifically, we define the operator $\mathcal{M}(\mbox{$\mathscr{X}$}_{t};\{\boldsymbol{G}_{m}\}_{m=1}^{2})=\boldsymbol{G}_{2}^{\top}(\boldsymbol{I}_{p_{2}}\otimes\boldsymbol{G}_{1}^{\top})\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})$, where $\boldsymbol{G}_{1}\in\mathbb{O}^{p_{1}\times r_{1}}$ and $\boldsymbol{G}_{2}\in\mathbb{O}^{p_{2}r_{1}\times r_{2}}$ are the component matrices for summarizing predictors, where $r_{1}$ and $r_{2}$ are the interim dimensions of features extracted from the corresponding mode. To interpret this construction, we liken the hierarchy of $\mbox{$\mathscr{X}$}_{t}\in\mathbb{R}^{p_{1}\times p_{2}}$ to a tree structure with $p_{2}$ branches, each branch containing $p_{1}$ leaves. For each branch, $\boldsymbol{G}_{1}^{\top}$ extracts $r_{1}$ features from its leaves to represent this branch. Overall, we have $r_{1}p_{2}$ features extracted from all leaves, given by $(\boldsymbol{I}_{p_{2}}\otimes\boldsymbol{G}_{1}^{\top})\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})$. Then the subsequent $\boldsymbol{G}_{2}^{\top}\in\mathbb{R}^{r_{2}\times p_{2}r_{1}}$ combines the extracted features from all branches to form $r_{2}$ features for the entire tree; see Figure 1(c) for the illustration of this hierarchical feature extraction procedure. Therefore, the overall loading matrix connecting the $p_{1}p_{2}$ leaves to the final $r_{2}$ features for the tree is $\boldsymbol{G}_{2}^{\top}(\boldsymbol{I}_{p_{2}}\otimes\boldsymbol{G}_{1}^{\top})$, and the column space of this loading can dwell in an arbitrary subspace of $\mathbb{R}^{r_{2}}$.
Remark 1.
Consider a special case with $p_{2}=2$, i.e., predictors $\mbox{$\mathscr{X}$}_{t}$ can be divided into two branches, $\boldsymbol{x}_{1:p_{1},1}$ and $\boldsymbol{x}_{1:p_{1},2}$. Suppose that the first branch has common features $\boldsymbol{G}^{*\top}\boldsymbol{x}_{1:p_{1},1}\in\mathbb{R}^{2}$ and branch-specific one $\boldsymbol{G}_{1}^{*\top}\boldsymbol{x}_{1:p_{1},1}\in\mathbb{R}$, while the second one has the common and branch-specific features of $\boldsymbol{G}^{*\top}\boldsymbol{x}_{1:p_{1},2}\in\mathbb{R}^{2}$ and $\boldsymbol{G}_{2}^{*\top}\boldsymbol{x}_{1:p_{1},2}\in\mathbb{R}^{2}$, respectively. Note that the HOFM uses loading matrices $\boldsymbol{G}_{1}^{(1)}\in\mathbb{R}^{p_{1}\times 3}$ and $\boldsymbol{G}_{1}^{(2)}\in\mathbb{R}^{p_{1}\times 4}$ to extract features from the first and second branches, respectively, while the proposed method employs a common loading matrix $\boldsymbol{G}_{1}\in\mathbb{R}^{p_{1}\times 5}$ to extract features from each branch. Shortly speaking, the HOFM concentrates more on branch-specific features, while our method focuses more on common features. In real applications, some weak branch-specific features may be ignored by the introduced method at the first layer of feature extraction, but they may also eventually be averaged out by the HOFM at later layers, i.e., our method does not miss globally important features.
For a general dimension of $M$, we assume that each mode of $\mbox{$\mathscr{X}$}_{t}\in\mathbb{R}^{p_{1}\times p_{2}\times\cdots\times p_{M}}$ is nested within its subsequent mode, and the feature extraction can then be iterated until the last mode of $\mbox{$\mathscr{X}$}_{t}$. Specifically, for each $1\leq m\leq M$, we define the component matrix for summarizing predictors as $\boldsymbol{G}_{m}\in\mathbb{O}^{r_{m-1}p_{m}\times r_{m}}$, where $r_{0}=1$, $r_{m}$ is the interim dimension of features extracted from the corresponding mode, and $r_{M}=r$. As a result, the feature extraction on predictors at (2.4) can be formulated below,
$\boldsymbol{f}_{t}^{\mathrm{predictor}}=\mathcal{M}(\mbox{$\mathscr{X}$}_{t};\{\boldsymbol{G}_{m}\}_{m=1}^{M})=\underbrace{\boldsymbol{G}_{M}^{\top}(\boldsymbol{I}_{p_{M}}\otimes\boldsymbol{G}_{M-1}^{\top})\cdots(\boldsymbol{I}_{p_{2}\cdots p_{M}}\otimes\boldsymbol{G}_{1}^{\top})}_{\vcentcolon=\boldsymbol{\Lambda}_{x}^{\top}}\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t}).$ (2.6)
Similarly, if the response tensor $\mbox{$\mathscr{Y}$}_{t}\in\mathbb{R}^{q_{1}\times q_{2}\times\cdots\times q_{N}}$ also has such purely hierarchical structure, the feature extraction on responses can be formulated into $\boldsymbol{f}_{t}^{\mathrm{response}}=\mathcal{N}(\mbox{$\mathscr{Y}$}_{t};\{\boldsymbol{H}_{n}\}_{n=1}^{N})=\boldsymbol{H}_{N}^{\top}(\boldsymbol{I}_{q_{N}}\otimes\boldsymbol{H}_{N-1}^{\top})\cdots(\boldsymbol{I}_{q_{2}\cdots q_{N}}\otimes\boldsymbol{H}_{1}^{\top})\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})\vcentcolon=\boldsymbol{\Lambda}_{y}^{\top}\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})$, where $\boldsymbol{H}_{n}\in\mathbb{O}^{s_{n-1}q_{n}\times s_{n}}$ with $s_{0}=1$ and $s_{N}=s$ are the component matrices for $1\leq n\leq N$. Note that $\boldsymbol{\Lambda}_{y}$ is orthonormal and, for the tensor regression model at (2.1), its coefficient tensor $\mathscr{A}$ can then be defined as $\mbox{$\mathscr{A}$}]_{N}=\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}\boldsymbol{\Lambda}_{x}^{\top}$ by adopting the method at (2.5), where $\boldsymbol{\Theta}\in\mathbb{R}^{s\times r}$ is a coefficient matrix.
Remark 2.
Consider a special case with $r=s$ and $\boldsymbol{\Theta}$ being diagonal, and then the above coefficient tensor $\mathscr{A}$ will admit a form of tensor train decomposition (oseledets2011tensor; zhou2022optimal; si2024efficient) if predictors are further arranged in a reverse hierarchical order. See Supplementary Material for the detailed illustration.
2.3 General hierarchical structure
This subsection discusses the case with partially hierarchical predictors and responses, and some of their modes may not be strictly nested within the others, possibly with unknown or even multiple hierarchical relationships among the modes.
We first consider the feature extraction on predictors. A natural way is first to assume a pseudo-full hierarchical structure to predictors, and the proposed method in the previous subsection can then be applied. The pseudo-structure should be consistent with the hierarchical structures intrinsic to the data, and the order of its modes is called an action order. For example, we consider the case with $M=3$, and mode 1 is nested within mode 2. The pseudo-structure can then be with an action order of $\{1,2,3\}$ or $\{3,1,2\}$.
In general, for the predictor tensor $\mbox{$\mathscr{X}$}_{t}\in\mathbb{R}^{p_{1}\times p_{2}\times\cdots\times p_{M}}$, an action order can be denoted by $\alpha=(\alpha_{(1)},\alpha_{(2)},\ldots,\alpha_{(M)})$, which is a permutation of set $\{1,2,\ldots,M\}$. We then introduce an operator $\mathcal{T}(\cdot;\cdot)$ to permute $\mbox{$\mathscr{X}$}_{t}$ such that its modes follow the action order $\alpha$,
$\mathcal{T}(\mbox{$\mathscr{X}$}_{t};\alpha)_{i_{1}\ldots i_{M}}=(\mbox{$\mathscr{X}$}_{t})_{i_{\alpha_{(1)}}\dots i_{\alpha_{(M)}}}\hskip 8.53581pt\text{and}\hskip 8.53581pt\operatorname*{vec}(\mathcal{T}(\mbox{$\mathscr{X}$}_{t};\alpha))=\boldsymbol{T}_{x}(\alpha)\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t}),$
where $\boldsymbol{T}_{x}(\alpha)\in\mathbb{R}^{P\times P}$ is the permutation matrix. Note that $\boldsymbol{T}_{x}(\alpha)=\boldsymbol{I}_{P}$ when $\alpha=(1,2,\ldots,M)$. As a result, by applying (2.6), we then can extract features from predictors
$\boldsymbol{f}_{t}^{\mathrm{predictor}}=\mathcal{M}(\mathcal{T}(\mbox{$\mathscr{X}$}_{t};\alpha);\{\boldsymbol{G}_{m}\in\mathbb{O}^{r_{m-1}p_{\alpha_{(m)}}\times r_{m}}\}_{m=1}^{M})\in\mathbb{R}^{r}\hskip 8.53581pt\text{with}\hskip 8.53581pt(r_{0},r_{M})=(1,r).$ (2.7)
Denote by $\mathcal{P}_{x}(M)$ the set of all possible action orders for predictors $\mbox{$\mathscr{X}$}_{t}\in\mathbb{R}^{p_{1}\times p_{2}\times\cdots\times p_{M}}$. Assume that $\mbox{$\mathscr{X}$}_{t}$ has $m$ hierarchical structures, which occupy $M^{\prime}$ modes in total, and the set $\mathcal{P}_{x}(M)$ can then be verified to have $(M-M^{\prime}+m)!$ possible action orders.
Proposition 1.
For any $\mbox{$\mathscr{X}$}\in\mathbb{R}^{p_{1}\times p_{2}\times\cdots\times p_{M}}$, an action order $\alpha=(\alpha_{(1)},\alpha_{(2)},\ldots,\alpha_{(M)})$ is used for the permutation, and factor matrices $\{\boldsymbol{G}_{m}\in\mathbb{O}^{r_{m-1}p_{\alpha_{(m)}}\times r_{m}}\}_{m=1}^{M}$ with $(r_{0},r_{M})=(1,r)$ are applied to extract features as in (2.7). It then holds that, for a different action order $\alpha^{\prime}=(\alpha_{(1)}^{\prime},\alpha_{(2)}^{\prime},\ldots,\alpha_{(M)}^{\prime})$, there exist $\{\boldsymbol{G}_{m}^{\prime}\in\mathbb{O}^{r^{\prime}_{m-1}p_{\alpha^{\prime}_{(m)}}\times r^{\prime}_{m}}\}_{m=1}^{M}$ with $(r_{0}^{\prime},r_{M}^{\prime})=(1,r)$ such that $\mathcal{M}(\mathcal{T}(\mbox{$\mathscr{X}$};\alpha);\{\boldsymbol{G}_{m}\}_{m=1}^{M})=\mathcal{M}(\mathcal{T}(\mbox{$\mathscr{X}$};\alpha^{\prime});\{\boldsymbol{G}_{m}^{\prime}\}_{m=1}^{M})$.
The above proposition proves a nice equivalence of all elements inside $\mathcal{P}_{x}(M)$, however, it is a theoretical result only. In real applications, each action order may have a totally different number of parameters, and a wrongly selected order may end up with a huge number of parameters. As a result, we may choose the one with the smallest size,
$\alpha^{*}=\operatorname*{arg\,min}_{\alpha\in\mathcal{P}_{x}(M)}d_{x}(\alpha)\hskip 8.53581pt\text{with}\hskip 8.53581ptd_{x}(\alpha)=\sum_{m=1}^{M}r_{m-1}r_{m}p_{m}.$ (2.8)
On the other hand, for a certain predictor tensor $\mbox{$\mathscr{X}$}_{t}$, more than one action orders may be more practical, while the summarized features $\boldsymbol{f}_{t}^{\mathrm{predictor}}$ have a fixed dimension of $r$ regardless which action orders are used. This motivates us to choose the best set of action orders,
$\mathcal{P}_{x}(M,K_{x})\vcentcolon=\{\alpha_{1}^{*},\ldots,\alpha_{K_{x}}^{*}\}=\operatorname*{arg\,min}_{1\leq K\leq|\mathcal{P}_{x}(M)|,\alpha_{k}\in\mathcal{P}_{x}(M)}\sum_{k=1}^{K}d_{x}(\alpha_{k})\hskip 8.53581pt\text{subject to}\hskip 8.53581pt\sum_{k=1}^{K}r_{M}(\alpha_{k})=r,$ (2.9)
where $|\mathcal{P}_{x}(M)|$ is the cardinality of $\mathcal{P}_{x}(M)$, and $r_{M}(\alpha_{k})$ is the rank of $r_{M}$ with respect to $\alpha_{k}$. Note that the number of parameters for $\mathcal{P}_{x}(M,K_{x})$ is smaller than that for $\alpha^{*}$ from (2.8).
For the selected $\alpha^{*}$ from (2.8), we can extract features from predictors as in (2.7). This paper focuses on $\mathcal{P}_{x}(M,K_{x})$, the selected set of action orders. For each $1\leq k\leq K_{x}$, denote $\alpha_{k}^{*}=(\alpha_{(1)}^{k},\alpha_{(2)}^{k},\ldots,\alpha_{(M)}^{k})$ with ranks $(r_{0,k},r_{1,k},\ldots,r_{M,k})$, and let $\boldsymbol{f}_{k,t}^{\mathrm{predictor}}=\mathcal{M}(\mathcal{T}(\mbox{$\mathscr{X}$}_{t};\alpha_{k}^{*});\{\boldsymbol{G}_{m,k}\in\mathbb{O}^{r_{m-1,k}p_{\alpha^{k}_{(m)}}\times r_{m,k}}\}_{m=1}^{M})\in\mathbb{R}^{r_{M,k}}$. As a result, the extracted features from predictors can be denoted by $\boldsymbol{f}_{t}^{\mathrm{predictor}}=\left(\boldsymbol{f}_{1,t}^{\mathrm{predictor}\top},\ldots,\boldsymbol{f}_{K_{x},t}^{\mathrm{predictor}\top}\right)^{\top}=\boldsymbol{\Lambda}_{x}^{\top}\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})\in\mathbb{R}^{r}$, where $r=\sum_{k=1}^{K_{x}}r_{M,k}$, and the loading matrix $\boldsymbol{\Lambda}_{x}$ has the form of
$\boldsymbol{\Lambda}_{x}=\begin{bmatrix}\boldsymbol{G}_{M,1}^{\top}(\boldsymbol{I}_{p_{\alpha^{1}_{(M)}}}\otimes\boldsymbol{G}_{M-1,1}^{\top})\cdots(\boldsymbol{I}_{p_{\alpha^{1}_{(2)}}\cdots p_{\alpha^{1}_{(M)}}}\otimes\boldsymbol{G}_{1,1}^{\top})\boldsymbol{T}_{x}(\alpha_{1})\\ \boldsymbol{G}_{M,2}^{\top}(\boldsymbol{I}_{p_{\alpha^{2}_{(M)}}}\otimes\boldsymbol{G}_{M-1,2}^{\top})\cdots(\boldsymbol{I}_{p_{\alpha^{2}_{(2)}}\cdots p_{\alpha^{2}_{(M)}}}\otimes\boldsymbol{G}_{1,2}^{\top})\boldsymbol{T}_{x}(\alpha_{2})\\ \cdots\\ \boldsymbol{G}_{M,K_{x}}^{\top}(\boldsymbol{I}_{p_{\alpha^{K_{x}}_{(M)}}}\otimes\boldsymbol{G}_{M-1,K_{x}}^{\top})\cdots(\boldsymbol{I}_{p_{\alpha^{K_{x}}_{(2)}}\cdots p_{\alpha^{K_{x}}_{(M)}}}\otimes\boldsymbol{G}_{1,K_{x}}^{\top})\boldsymbol{T}_{x}(\alpha_{K_{x}})\end{bmatrix}^{\top}$. (2.10)
The above matrix is no longer orthogonal in general. The philosophy of the above feature extraction is analogous to that of CP decomposition, which concatenates multiple separately extracted features. Note that, comparing with Tucker decomposition, CP decomposition requests much less parameters by sacrificing orthogonality of loading matrices.
The feature extraction from responses $\mbox{$\mathscr{Y}$}_{t}\in\mathbb{R}^{q_{1}\times q_{2}\times\cdots\times q_{N}}$ can be conducted in a similar way. Denote the corresponding action order by $\beta=(\beta_{(1)},\beta_{(2)},\ldots,\beta_{(N)})$, which is a permutation of set $\{1,2,\ldots,N\}$, and let $\mathcal{P}_{y}(N)$ be the set of all possible action orders. By a method similar to (2.8) and (2.9), we can define the best action order $\beta^{*}$ and the best set of action orders $\mathcal{P}_{y}(N,K_{y})\vcentcolon=\{\beta_{1}^{*},\ldots,\beta_{K_{y}}^{*}\}=\operatorname*{arg\,min}_{1\leq K\leq|\mathcal{P}_{y}(N)|,\beta_{k}\in\mathcal{P}_{y}(N)}\sum_{k=1}^{K}d_{y}(\beta_{k})$ with $d_{y}(\beta)=\sum_{n=1}^{N}s_{n-1}s_{n}q_{n}$, respectively. Moreover, for each $1\leq k\leq K_{y}$, let $\beta_{k}^{*}=(\beta_{(1)}^{k},\beta_{(2)}^{k},\ldots,\beta_{(N)}^{k})$ with ranks $(s_{0,k},s_{1,k},\ldots,s_{N,k})$, component matrices be $\boldsymbol{H}_{n,{k}}\in\mathbb{O}^{s_{n-1,{k}}q_{\beta^{k}_{(n)}}\times s_{n,{k}}}$ with $1\leq n\leq N$, and the corresponding extracted features be $\boldsymbol{f}_{{k},t}^{\mathrm{response}}=\mathcal{N}(\mathcal{T}(\mbox{$\mathscr{Y}$}_{t};\beta^{*}_{k});\{\boldsymbol{H}_{n,{k}}\}_{n=1}^{N})$. As a result, the total extracted features from responses can be represented as $\boldsymbol{f}_{t}^{\mathrm{response}}=\left(\boldsymbol{f}_{1,t}^{\mathrm{response}\top},\ldots,\boldsymbol{f}_{K_{y},t}^{\mathrm{response}\top}\right)^{\top}\vcentcolon=\boldsymbol{\Lambda}_{y}^{\top}\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})\in\mathbb{R}^{s}$, where $s=\sum_{k=1}^{K_{y}}s_{N,k}$, and the loading matrix $\boldsymbol{\Lambda}_{y}$ can be defined similar to (2.10) with the permutation matrix $\boldsymbol{T}_{y}(\beta)\in\mathbb{R}^{Q\times Q}$.
From (2.5), the coefficient tensor of model (2.1) can be designed to have the form of
$[\mbox{$\mathscr{A}$}_{\mathrm{LR}}]_{N}=\boldsymbol{\Lambda}_{y}\left(\boldsymbol{\Lambda}_{y}^{\top}\boldsymbol{\Lambda}_{y}\right)^{-1}\boldsymbol{\Theta}_{\mathrm{LR}}\boldsymbol{\Lambda}_{x}^{\top}$ (2.11)
with $\boldsymbol{\Theta}_{\mathrm{LR}}\in\mathbb{R}^{s\times r}$ being the coefficient matrix, leading to an efficient and interpretable tensor regression model for data with hierarchical structures.
2.4 Autoregressive model for hierarchical time series
Consider an autoregressive model for hierarchical time series $\{\mbox{$\mathscr{Y}$}_{t}\}$ with $\mbox{$\mathscr{Y}$}_{t}\in\mathbb{R}^{q_{1}\times q_{2}\times\cdots\times q_{N}}$,
[…]
Assumption 1.
The determinant of the matrix polynomial $\mbox{$\mathscr{A}$}(z)=\boldsymbol{I}_{Q}-\boldsymbol{A}_{1}z-\cdots-\boldsymbol{A}_{L}z^{L}$ is not equal to zero for all $z\in\mathbb{C}$ and $|z|<1$.
We first consider the case with $L=1$, and the coefficient tensor $\mbox{$\mathscr{A}$}_{1}$ can be directly assumed to have the decomposition in (2.11) with the same definitions of $\boldsymbol{\Lambda}_{x}$ and $\boldsymbol{\Lambda}_{y}$, where $M=N$, $p_{n}=q_{n}$ for $1\leq n\leq N$, and $\mbox{$\mathscr{X}$}_{t}=\mbox{$\mathscr{Y}$}_{t-1}$. As a result, model (2.12) can then be rewritten into an interpretable form:
[…]
Note that the high-order tensor $\mbox{$\mathscr{Y}$}_{t}$ is summarized into features $\boldsymbol{f}_{t}^{\mathrm{response}}\in\mathbb{R}^{s}$, while the information from $\mbox{$\mathscr{Y}$}_{t-1}$ is compressed into $\boldsymbol{f}^{\mathrm{predictor}}_{t}\in\mathbb{R}^{r}$. The coefficient matrix $\boldsymbol{\Theta}_{1}\in\mathbb{R}^{s\times r}$ characterizes the linear relationship between features $\boldsymbol{f}^{\mathrm{response}}_{t}$ and $\boldsymbol{f}^{\mathrm{predictor}}_{t}$.
For a general order $L$, we apply (2.5) to each coefficient tensor $\mbox{$\mathscr{A}$}_{l}$, resulting in:
$\mbox{$\mathscr{A}$}_{l}]_{N}=\boldsymbol{\Lambda}_{y}\left(\boldsymbol{\Lambda}_{y}^{\top}\boldsymbol{\Lambda}_{y}\right)^{-1}\boldsymbol{\Theta}_{l}\boldsymbol{\Lambda}_{x}^{\top}\quad\text{for}\quad 1\leq l\leq L.$ (2.13)
[…]
3 High-dimensional estimation
3.1 Ordinary least squares approach
This paper considers the ordinary least squares (OLS) method to estimate the proposed tensor autoregressive model at (2.12) and (2.13) with hierarchical time series.
Suppose that both sets of action orders $\mathcal{P}_{x}(M,K_{x})=\{\alpha_{1},\ldots,\alpha_{K_{x}}\}$ and $\mathcal{P}_{y}(N,K_{y})=\{\beta_{1},\ldots,\beta_{K_{y}}\}$ are known and, for all $1\leq k\leq K_{x}$ and $1\leq k^{\prime}\leq K_{y}$, the corresponding ranks $\{r_{m,k}\}_{m=1}^{M}$ and $\{s_{n,k^{\prime}}\}_{n=1}^{N}$ are also given. Let $\boldsymbol{p}=(p_{1},\ldots,p_{M})$, $\boldsymbol{q}=(q_{1},\ldots,q_{N})$, $\boldsymbol{r}_{k}=(r_{1,k},\ldots,r_{M,k})$ and $\boldsymbol{s}_{k^{\prime}}=(s_{1,k^{\prime}},\ldots,s_{N,k^{\prime}})$, where $1\leq k\leq K_{x}$ and $1\leq k^{\prime}\leq K_{y}$. Note that $\sum_{k=1}^{K_{x}}r_{M,k}=r$, $\sum_{k=1}^{K_{y}}s_{N,k}=s$, $M=N$ and $\boldsymbol{p}=\boldsymbol{q}$. We then denote the parameter spaces of loading matrices from predictor and response features respectively by
[…]
For autoregressive models at (2.12) and (2.13), the coefficient tensor can be denoted by $\mbox{$\mathscr{A}$}_{\mathrm{AR}}\in\mathbb{R}^{q_{1}\times\cdots\times q_{N}\times q_{1}\times\cdots\times q_{N}\times L}$ with $\mbox{$\mathscr{A}$}_{\mathrm{AR}}]_{N}=([\mbox{$\mathscr{A}$}_{1}]_{N},[\mbox{$\mathscr{A}$}_{2}]_{N},\ldots,[\mbox{$\mathscr{A}$}_{L}]_{N})$ being the mode-$N$ sequential matricization. Let $\boldsymbol{\Theta}_{\mathrm{AR}}=[\boldsymbol{\Theta}_{1},\ldots,\boldsymbol{\Theta}_{L}]\in\mathbb{R}^{s\times rL}$, and it holds that $\mbox{$\mathscr{A}$}_{\mathrm{AR}}]_{N}=\boldsymbol{\Lambda}_{y}(\boldsymbol{\Lambda}_{y}^{\top}\boldsymbol{\Lambda}_{y})^{-1}\boldsymbol{\Theta}_{\mathrm{AR}}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top})$. Moreover, by reparameterizing $(\boldsymbol{\Lambda}_{y}^{\top}\boldsymbol{\Lambda}_{y})^{-1}\boldsymbol{\Theta}_{\mathrm{AR}}$ into $\boldsymbol{\Theta}_{\mathrm{AR}}$, we have $\mbox{$\mathscr{A}$}_{\mathrm{AR}}]_{N}=\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}_{\mathrm{AR}}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top})$, and then the parameter space can be defined as
$\begin{split}\mathcal{S}_{\mathrm{AR}}=\{\mbox{$\mathscr{A}$}\in&\mathbb{R}^{q_{1}\times\cdots\times q_{N}\times p_{1}\times\cdots\times p_{M}\times L}:\penalty 10000\ [\mbox{$\mathscr{A}$}]_{N}=\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top}),\penalty 10000\ \boldsymbol{\Theta}\in\mathbb{R}^{s\times rL},\\ &\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \boldsymbol{\Lambda}_{y}\in\mathcal{R}_{y}(\boldsymbol{q},\{\boldsymbol{s}_{k}\}_{k=1}^{K_{y}},\mathcal{P}_{y}(N,K_{y})),\penalty 10000\ \boldsymbol{\Lambda}_{x}\in\mathcal{R}_{x}(\boldsymbol{p},\{\boldsymbol{r}_{k}\}_{k=1}^{K_{x}},\mathcal{P}_{x}(M,K_{x}))\}.\end{split}$
For an observed time series $\{\mbox{$\mathscr{Y}$}_{1-L},\ldots,\mbox{$\mathscr{Y}$}_{0},\mbox{$\mathscr{Y}$}_{1},\ldots,\mbox{$\mathscr{Y}$}_{T}\}$ generated by the model at (2.12) and (2.13), the predictor tensor $\mbox{$\mathscr{X}$}_{t}\in\mathbb{R}^{q_{1}\times\cdots\times q_{N}\times L}$ can be formed by stacking $\mbox{$\mathscr{Y}$}_{t-1},\dots,\mbox{$\mathscr{Y}$}_{t-L}$ together, and the OLS estimator is hence defined as
$\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{AR}}=\operatorname*{arg\,min}_{\mbox{$\mathscr{A}$}\in\mathcal{S}_{\mathrm{AR}}}\mathcal{L}_{T}(\mbox{$\mathscr{A}$})\quad\text{with}\quad\mathcal{L}_{T}(\mbox{$\mathscr{A}$})=\frac{1}{T}\sum_{t=1}^{T}\left\lVert\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})-[\mbox{$\mathscr{A}$}]_{N}\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})\right\rVert_{2}^{2}.$
Moreover, for the tensor regression model at (2.1) and (2.11), by letting $L=1$, we can similarly define its parameter space and OLS estimator, denoted by $\mathcal{S}_{\mathrm{LR}}$ and $\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{LR}}$, respectively.
Denote by $\mbox{$\mathscr{A}$}^{*}$ the true coefficient tensor of the tensor regression or autoregressive model in the previous section, and it depends on the best set of action orders, which is unknown in real applications. In fact, $\mathcal{P}_{x}(M,K_{x})$ and $K_{x}$ in this section are only the running set and numbers of action orders for predictors, respectively, and the situation is the same for $\mathcal{P}_{y}(N,K_{y})$ and $K_{y}$. Fortunately, from Proposition 1, we can project the true coefficient tensor $\mbox{$\mathscr{A}$}^{*}$ into the running sets $\mathcal{P}_{x}(M,K_{x})$ and $\mathcal{P}_{y}(N,K_{y})$, with running numbers $K_{x}$ and $K_{y}$. Such projection has no error, while the number of parameters will be enlarged. Without loss of generality, we will assume $\mathcal{P}_{x}(M,K_{x})$ and $\mathcal{P}_{y}(N,K_{y})$ to be the best sets in establishing theoretical properties, and Section 4 will discuss the selection of these hyperparameters.
3.2 Non-asymptotic properties for tensor regression
[…]
We begin with the tensor-on-tensor regression setting, where $\{(\mbox{$\mathscr{Y}$}_{t},\mbox{$\mathscr{X}$}_{t})\}_{t=1}^{T}$ are assumed to be independent and identically distributed (i.i.d.) observations. The following assumptions are imposed.
Assumption 2.
Predictors $\{\boldsymbol{x}_{t}\}$ are i.i.d. with $\mathbb{E}(\boldsymbol{x}_{t})=\boldsymbol{0}$, $\mathrm{var}(\boldsymbol{x}_{t})=\boldsymbol{\Sigma}_{x}$, and follow a $\sigma^{2}$-sub-Gaussian distribution. There exist constants $0<c_{x}<C_{x}<\infty$ such that $c_{x}\leq\lambda_{\min}(\boldsymbol{\Sigma}_{x})\leq\lambda_{\max}(\boldsymbol{\Sigma}_{x})\leq C_{x}$, where the two quantities $c_{x}$ and $C_{x}$ depend on the dimensions of $p_{j}$’s, and they may shrink to zero or diverge to infinity as the dimensions increase.
Assumption 3.
Error terms $\{\boldsymbol{e}_{t}\}$ are i.i.d. and, conditional on $\boldsymbol{x}_{t}$, $\boldsymbol{e}_{t}$ has mean zero and follows $\kappa^{2}$-sub-Gaussian distribution.
Assumption 4.
The matrix $\boldsymbol{\Theta}_{\mathrm{LR}}$ has its operator norm upper bounded by a constant $g_{1}>0$.
Assumption 2 guarantees the well-conditioning of the covariance matrix of the predictors. Assumption 3, on the other hand, controls the tail behavior of the noise by assuming sub-Gaussianity, a standard requirement in high-dimensional statistics that facilitates concentration inequalities. Both assumptions are widely adopted in the literature on high-dimensional regression and low-rank estimation; see, for example, wainwright2019high; si2024efficient; cai2025efficient. Assumption 4 imposes an upper bound on the operator norm of the coefficient matrix $\boldsymbol{\Theta}_{\mathrm{LR}}$. This is a mild technical condition, as the dimensions $s_{\mathrm{tot}}$ and $r_{\mathrm{tot}}$ of $\boldsymbol{\Theta}_{\mathrm{LR}}$ are typically small by carefully choosing the action orders.
The following theorem provides a non-asymptotic upper bound on the estimation error of estimator $\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{LR}}$.
Theorem 1.
Suppose that Assumptions 2 - 4 hold, and the true coefficient tensor $\mbox{$\mathscr{A}$}_{\mathrm{LR}}^{*}\in\mathcal{S}_{\mathrm{LR}}$. If the sample size $T\gtrsim(\sigma^{4}/c_{x}^{2})d_{\mathrm{LR}}\log((\sigma^{4}/c_{x}^{2})d_{\mathrm{LR}})$, then there exists $c_{1}>0$ such that
$\|\mbox{$\mathscr{\widehat{A}}$}_{\mathrm{LR}}-\mbox{$\mathscr{A}$}^{*}_{\mathrm{LR}}\|_{\mathrm{F}}\lesssim\frac{\kappa\sigma}{c_{x}}\sqrt{\frac{d_{\mathrm{LR}}\log T}{T}}$
with probability at least $1-2\exp\left\{-c_{1}d_{\mathrm{LR}}\log T\right\}$, where $d_{\mathrm{LR}}=\sum_{k=1}^{K_{x}}d_{x}(\alpha_{k})+\sum_{k=1}^{K_{y}}d_{y}(\beta_{k})+sr$.
We define $d_{\mathrm{LR}}$ as the model complexity of the low-rank regression estimator, which can be rewritten explicitly as $d_{\mathrm{LR}}=K_{x}r_{\max}^{2}\sum_{m=1}^{M}p_{m}+K_{y}s_{\max}^{2}\sum_{n=1}^{N}q_{n}+K_{x}K_{y}r_{\max}s_{\max}$, where $r_{\max}\vcentcolon=\max_{m,k}r_{m,k}$ and $s_{\max}\vcentcolon=\max_{n,k^{\prime}}s_{n,k^{\prime}}$. From Theorem 1, our proposed low-rank regression estimator $\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{LR}}$, up to a logarithmic factor, achieves an estimation error bound of order $O_{p}(\sqrt{d_{\mathrm{LR}}/T})$ when the sample size $T\gtrsim d_{\mathrm{LR}}/\text{SNR}^{2}$, where $\text{SNR}=c_{x}/\sigma^{2}$ is the predictor signal-to-noise ratio. This rate aligns with the general optimal rate for low-rank regression established in han2022optimal. The model complexity $d_{\mathrm{LR}}$ grows linearly with the dimensions of each mode $\{p_{m}\}$ and $\{q_{n}\}$, and quadratically with the interim ranks $\{r_{m,k}\}$ and $\{s_{n,k^{\prime}}\}$. The logarithmic factor $\log T$, negligible in most common cases, arises from the non-orthogonality of the loadings. In contrast, if we assume the coefficient tensor $\mathscr{A}$ follows a Tucker decomposition with Tucker ranks $(s_{1},s_{2},\ldots,s_{N},r_{1},r_{2},\ldots,r_{M})$, the corresponding complexity would be $d_{\mathrm{Tucker}}=\prod_{n=1}^{N}s_{n}\times\prod_{m=1}^{M}r_{m}+\sum_{n=1}^{N}q_{n}s_{n}+\sum_{m=1}^{M}p_{m}r_{m}$, which grows exponentially with the tensor orders $N$ and $M$ (han2022optimal; wang2022high).
3.3 Non-asymptotic properties for tensor autoregression
We now turn to the tensor autoregressive model specified in (2.12). This setting introduces temporal dependence, and the assumptions are adapted accordingly.
Assumption 5.
Let error terms $\boldsymbol{e}_{t}=\boldsymbol{\Sigma}_{e}^{1/2}\boldsymbol{\xi}_{t}$. Random vectors $\{\boldsymbol{\xi}_{t}\}$ are i.i.d. with mean zero and covariance $\boldsymbol{I}_{Q}$, the entries $(\boldsymbol{\xi}_{tj})_{1\leq j\leq Q}$ of $\boldsymbol{\xi}_{t}$ are mutually independent and $\kappa^{2}$-sub-Gaussian distributed, and there exist constants $0<c_{e}<C_{e}<\infty$ such that $c_{e}\leq\lambda_{\min}(\boldsymbol{\Sigma}_{e})\leq\lambda_{\max}(\boldsymbol{\Sigma}_{e})\leq C_{e}$.
Assumption 6.
The matrix $\boldsymbol{\Theta}_{\mathrm{AR}}$ has its operator norm upper bounded by a constant $g_{2}>0$.
The sub-Gaussianity assumption in Assumption 5 is commonly used in the analysis of high-dimensional time series models. It provides control over the tail behavior of the error terms and facilitates the use of concentration inequalities. This assumption is widely adopted in recent literature, including zheng2020finite; wang2024high; huang2025supervised. Assumption 6 is essential for the consistency of the estimator in the autoregressive setting, which is relatively mild and practically reasonable. In our setting, $\boldsymbol{\Lambda}_{y}$ and $\boldsymbol{\Lambda}_{x}$ may exhibit multi-collinearity, and this assumption is weaker than enforcing a lower bound on the smallest singular value of $\boldsymbol{\Lambda}_{x}$ and $\boldsymbol{\Lambda}_{y}$ for model well-behavedness.
Following basu2015regularized, we define the following quantities,
$\mu_{\mathrm{min}}(\mbox{$\mathscr{A}$})=\min_{|z|=1}\lambda_{\mathrm{min}}(\bar{\mbox{$\mathscr{A}$}}(z)\mbox{$\mathscr{A}$}(z))\quad\text{and}\quad\mu_{\mathrm{max}}(\mbox{$\mathscr{A}$})=\max_{|z|=1}\lambda_{\mathrm{max}}(\bar{\mbox{$\mathscr{A}$}}(z)\mbox{$\mathscr{A}$}(z)),$
where $\bar{\mbox{$\mathscr{A}$}}(z)$ denotes the conjugate transpose of $\mbox{$\mathscr{A}$}(z)$.
Furthermore, the AR($L$) model in (2.12) can be alternatively represented as a $QL$-dimensional AR(1) model with its transition matrix $\boldsymbol{B}$ defined in (C.3) in the Supplementary Material. We similarly define the matrix polynomial associated with the lifted AR(1) representation as $\boldsymbol{B}(z)=\boldsymbol{I}_{QL}-\boldsymbol{B}z$. We then define
$\mu_{\mathrm{min}}(\boldsymbol{B})=\min_{|z|=1}\lambda_{\mathrm{min}}(\bar{\boldsymbol{B}}(z)\boldsymbol{B}(z))\quad\text{and}\quad\mu_{\mathrm{max}}(\boldsymbol{B})=\max_{|z|=1}\lambda_{\mathrm{max}}(\bar{\boldsymbol{B}}(z)\boldsymbol{B}(z)),$
which quantify the smallest and largest eigenvalues the Hermitian matrix $\bar{\boldsymbol{B}}(z)\boldsymbol{B}(z)$ can take on the unit circle. Note that $\mu_{\mathrm{min}}(\boldsymbol{B})$, $\mu_{\mathrm{max}}(\boldsymbol{B})$ are not necessarily the same as $\mu_{\mathrm{min}}(\mbox{$\mathscr{A}$})$, $\mu_{\mathrm{max}}(\mbox{$\mathscr{A}$})$ (basu2015regularized), and we can show $\mu_{\mathrm{min}}(\boldsymbol{B}^{\ast})\leq\mu_{\mathrm{max}}(\mbox{$\mathscr{A}$}^{\ast})$.
We then state the error upper bound for the estimator $\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{AR}}$ below.
Theorem 2.
Suppose that Assumptions 1, 5, and 6 hold, and the true coefficient tensor $\mbox{$\mathscr{A}$}_{\mathrm{AR}}^{\ast}\in\mathcal{S}_{\mathrm{AR}}$. If the sample size $T\gtrsim(\kappa^{4}\kappa_{U,B}^{2}/\kappa_{L,A}^{2})d_{\mathrm{AR}}\log((\kappa^{4}\kappa_{U,B}^{2}/\kappa_{L,A}^{2})d_{\mathrm{AR}})$, then
$\|{\mbox{$\mathscr{\widehat{A}}$}_{\mathrm{AR}}-\mbox{$\mathscr{A}$}_{\mathrm{AR}}^{\ast}}\|_{\mathrm{F}}\lesssim\frac{\kappa^{2}\kappa_{U,B}^{\prime}}{\kappa_{L,A}}\sqrt{\frac{d_{\mathrm{AR}}\log T}{T}}$
with probability at least $1-2\exp\left\{-c_{1}d_{\mathrm{AR}}\log T\right\}$, where $d_{\mathrm{AR}}=\sum_{k=1}^{K_{x}}d_{x}(\alpha_{k})+\sum_{k=1}^{K_{y}}d_{y}(\beta_{k})+Lsr$, $\boldsymbol{B}^{\ast}$ is the true value of $\boldsymbol{B}$ constructed from $\mbox{$\mathscr{A}$}^{\ast}$, $\kappa_{U,B}=C_{e}/\mu_{\mathrm{min}}(\boldsymbol{B}^{\ast})$, $\kappa_{L,A}=c_{e}/\mu_{\mathrm{max}}(\mbox{$\mathscr{A}$}^{\ast})$, $\kappa^{\prime}_{U,B}=C_{e}/\mu^{1/2}_{\mathrm{min}}(\boldsymbol{B}^{\ast})$, and $c_{1}>0$ is a constant.
We define $d_{\mathrm{AR}}$ as the model complexity of the low-rank autoregressive estimator, which can be explicitly expressed as $d_{\mathrm{AR}}=(K_{x}r_{\max}^{2}+K_{y}s_{\max}^{2})\sum_{n=1}^{N}q_{n}+LK_{x}K_{y}r_{\max}s_{\max}$. When the order $N$, the number of action orders $K_{x}$, $K_{y}$, and the largest ranks $s_{\max}$ and $r_{\max}$ are held fixed, the key parameters $c_{e}$, $C_{e}$, $\kappa^{2}$, $\mu_{\mathrm{min}}(\boldsymbol{B}^{\ast})$, and $\mu_{\mathrm{max}}(\mbox{$\mathscr{A}$}^{\ast})$ are all bounded away from zero and infinity. Consequently, the estimation error of the low-rank tensor autoregressive estimator satisfies $\|{\mbox{$\mathscr{\widehat{A}}$}_{\mathrm{AR}}-\mbox{$\mathscr{A}$}_{\mathrm{AR}}^{\ast}}\|_{\mathrm{F}}=O_{p}(\sqrt{d_{\mathrm{AR}}/T})$ up to a logarithmic term, where $d_{\mathrm{AR}}$ is the effective model complexity of the autoregressive estimator.
4 Algorithms and hyperparameter selection
4.1 Algorithms
This subsection introduces an algorithm for computing the low-rank estimator defined in (3.1) under the general setting. To preserve the sequential low-rank structure of $\mbox{$\mathscr{A}$}]_{N}$, we adopt an alternating least squares (ALS) approach, which iteratively updates each component $\boldsymbol{H}_{n,k^{\prime}}$ and $\boldsymbol{G}_{m,k}$ while keeping all other components fixed.
We introduce some shorthand notations for clarity. Let $\boldsymbol{y}_{t}=\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})$ and $\boldsymbol{x}_{t}=\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})=(\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t-1})^{\top},\dots,\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t-L})^{\top})^{\top}$, and our model can be rewritten in the vectorized form as:
$\boldsymbol{y}_{t}=\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top})\boldsymbol{x}_{t}+\boldsymbol{e}_{t},$ (4.1)
[…]
We sequentially update each orthogonal matrix in each row block of $\boldsymbol{\Lambda}_{x}^{\top}$ (i.e., update $\boldsymbol{G}_{1,1}$ up to $\boldsymbol{G}_{M,1}$ for the first row block), and then switch to the next row block for another sequence of updates, and so on. Specifically, we convert the update of each component into a regression problem relegated to Supplementary Material, similar to coordinate descent. The update for $\boldsymbol{\Lambda}_{y}^{\top}$ follows a similar procedure. After updating all components in $\boldsymbol{\Lambda}_{x}^{\top}$ and $\boldsymbol{\Lambda}_{y}^{\top}$, we orthogonalize each row block of the two loading matrices and update the core tensor $\boldsymbol{\Theta}$ via an SSVD algorithm detailed in Supplementary Material. This process is repeated until convergence. For convenience, we arrange these matrices into $\boldsymbol{L}_{1},\boldsymbol{L}_{2},\ldots,\boldsymbol{L}_{K_{x}M+K_{y}N}$, where $\boldsymbol{L}_{i}$ is the $i$-th orthogonal matrix in the update procedure, and we re-define the loss function with regard to $\boldsymbol{L}_{i}$’s and $\boldsymbol{\Theta}$ as follows:
$\mathcal{L}_{T}\big(\boldsymbol{L}_{1},\dots,\boldsymbol{L}_{K_{x}M+K_{y}N},\boldsymbol{\Theta}\big)=\frac{1}{T}\sum_{t=1}^{T}\|\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})-\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top})\operatorname*{vec}{(\mbox{$\mathscr{X}$}_{t})}\|^{2},$
$\text{s.t., }\boldsymbol{G}_{m,k}=\boldsymbol{L}_{(k-1)M+m}\quad\text{and}\quad\boldsymbol{H}_{n,k^{\prime}}=\boldsymbol{L}_{K_{x}M+(k^{\prime}-1)N+n}$ (4.2)
for all $m,k,n,k^{\prime}$, and $\boldsymbol{\Lambda}_{x}$ and $\boldsymbol{\Lambda}_{y}$ are defined in (2.10), respectively. The overall algorithm is summarized below:
Algorithm 1 ALS for low-rank estimation
Input: Data $\{(\boldsymbol{y}_{t},\boldsymbol{x}_{t})\}_{t=1}^{T}$, initial coefficient $\mbox{$\mathscr{A}$}^{(0)}]_{N}=\boldsymbol{\Lambda}_{y}^{(0)}\boldsymbol{\Theta}^{(0)}\boldsymbol{\Lambda}_{x}^{(0)\top}$, and components $\{\{\boldsymbol{H}^{(0)}_{n,k^{\prime}}\}_{n=1}^{N}\}_{k^{\prime}=1}^{K_{y}}$ and $\{\{\boldsymbol{G}^{(0)}_{m,k}\}_{m=1}^{M}\}_{k=1}^{K_{x}}$ arranged into $\{\boldsymbol{L}_{i}^{(0)}\}_{i=1}^{K_{x}M+K_{y}N}$.
repeat
  for $i=1,2,\dots,K_{x}M+K_{y}N$ do
   $\boldsymbol{L}^{(u+0.5)}_{i}=\operatorname*{arg\,min}_{\boldsymbol{L}_{i}}\mathcal{L}_{T}\big(\boldsymbol{L}_{1}^{(u+0.5)},\dots,\boldsymbol{L}_{i-1}^{(u+0.5)},\boldsymbol{L}_{i},\boldsymbol{L}_{i+1}^{(u)},\dots,\boldsymbol{L}_{K_{x}M+K_{y}N}^{(u)},\boldsymbol{\Theta}^{(u)}\big)$
  end for
  $\{\boldsymbol{L}_{i}^{(u+1)}\}_{i=1}^{K_{x}M+K_{y}N},\boldsymbol{\Theta}^{(u)}=\text{SSVD}(\{\boldsymbol{L}_{i}^{(u+0.5)}\}_{i=1}^{K_{x}M+K_{y}N},\boldsymbol{\Theta}^{(u)})$
  $\boldsymbol{\Theta}^{(u+1)}=\operatorname*{arg\,min}_{\boldsymbol{\Theta}}\mathcal{L}_{T}\big(\boldsymbol{L}_{1}^{(u+1)},\dots,\boldsymbol{L}_{K_{x}M+K_{y}N}^{(u+1)},\boldsymbol{\Theta}\big)$
until convergence or $u=u_{\max}$, where $u_{\max}$ is the maximum iteration
Return $\mbox{$\mathscr{A}$}^{(u)}]_{N}=\boldsymbol{\Lambda}_{y}^{(u)}\boldsymbol{\Theta}^{(u)}\boldsymbol{\Lambda}_{x}^{(u)\top}$ via (4.2), (2.10).
4.2 Hyperparameter selection
Some hyperparameters involved in Algorithm 1 are typically unknown in most real-world applications, and their selection not only impacts the model’s performance, but also reveals the underlying structural information of the data. This subsection introduces a boosting-based procedure for determining the number of action orders $K_{y}$ and $K_{x}$, the corresponding action sets $\mathcal{P}_{y}(N,K_{y})$ and $\mathcal{P}_{x}(M,K_{x})$, and the associated rank sequences $\{\boldsymbol{s}_{k^{\prime}}\}_{k^{\prime}=1}^{K_{y}}$ and $\{\boldsymbol{r}_{k}\}_{k=1}^{K_{x}}$.
To guide the selection of these hyperparameters, we adopt an information criterion. Let $\mbox{$\mathscr{\widehat{A}}$}(R)$ denote the estimator from (3.1), where the configuration $R$ collects all hyperparameters: $R=(K_{y},K_{x},\mathcal{P}_{y}(N,K_{y}),\mathcal{P}_{x}(M,K_{x}),\{\boldsymbol{s}_{k^{\prime}}\}_{k^{\prime}=1}^{K_{y}},\{\boldsymbol{r}_{k}\}_{k=1}^{K_{x}})$. The Bayesian information criterion (BIC) is then constructed below,
$\mathrm{BIC}(R)=T\log\left\{\mathcal{L}_{T}(\mbox{$\mathscr{\widehat{A}}$}(R))\right\}+\phi d(R)\log(T),$ (4.3)
where $\phi$ is a tuning parameter, and $d(R)=\sum_{k^{\prime}=1}^{K_{y}}\sum_{n=1}^{N}q_{n}s_{n-1,k^{\prime}}s_{n,k^{\prime}}+\sum_{k=1}^{K_{x}}\sum_{n=1}^{N}q_{n}r_{n-1,k}r_{n,k}+Ls_{\mathrm{tot}}r_{\mathrm{tot}}$ is the total number of parameters.
The boosting-based hyperparameter selection procedure consists of two main stages: action order selection and rank reduction, which are detailed below. The overall procedure is summarized in Algorithm 2. Finally, the lag length $L$ is determined via a simple grid search to minimize the BIC over a predetermined candidate set $\{1,2,\ldots,L_{\max}\}$.
4.2.1 Action order selection
To construct the boosting-based algorithm, we first introduce the key notations and setup. Define $\widetilde{\boldsymbol{\Lambda}}_{x}(\alpha_{k},\boldsymbol{r}_{k})=\boldsymbol{\Lambda}_{x}(\{\alpha_{k}\})$, where the operator $\boldsymbol{\Lambda}_{x}(\cdot)$ is defined in (2.10) and $\boldsymbol{r}_{k}$ is the associated rank. It can be verified that for any two positive integer vectors $\boldsymbol{r}_{k,1}$ and $\boldsymbol{r}_{k,2}$, there exists a matrix $\widetilde{\boldsymbol{\Lambda}}_{x}(\alpha_{k},\boldsymbol{r}_{k,3})$ with $\boldsymbol{r}_{k,3}\leq\boldsymbol{r}_{k,1}+\boldsymbol{r}_{k,2}$ by elementwise comparison, satisfying
$\widetilde{\boldsymbol{\Lambda}}_{x}(\alpha_{k},\boldsymbol{r}_{k,3})=\boldsymbol{O}\left[\widetilde{\boldsymbol{\Lambda}}_{x}(\alpha_{k},\boldsymbol{r}_{k,1}),\widetilde{\boldsymbol{\Lambda}}_{x}(\alpha_{k},\boldsymbol{r}_{k,2})\right],$ (4.4)
where $\boldsymbol{O}$ is an invertible matrix that can be absorbed into $\boldsymbol{\Theta}$. The same applies to $\widetilde{\boldsymbol{\Lambda}}_{y}(\cdot)$ by defining $\widetilde{\boldsymbol{\Lambda}}_{y}(\beta_{k^{\prime}},\boldsymbol{s}_{k^{\prime}})=\boldsymbol{\Lambda}_{y}(\{\beta_{k^{\prime}}\})$, where $\boldsymbol{s}_{k^{\prime}}$ is the associated rank and $\boldsymbol{\Lambda}_{y}(\cdot)$ is defined similarly to (2.10). The proof is provided in Supplementary Material.
We refer to any model defined by (4.1) with $\boldsymbol{\Lambda}_{y}=\widetilde{\boldsymbol{\Lambda}}_{y}(\beta_{k^{\prime}},\boldsymbol{1}_{N})$ and $\boldsymbol{\Lambda}_{x}=\widetilde{\boldsymbol{\Lambda}}_{x}(\alpha_{k},\boldsymbol{1}_{N})$ as a weak learner, denoted by $WL(\beta_{k^{\prime}},\alpha_{k})$, where $\boldsymbol{1}_{N}\in\mathbb{R}^{N}$ is a vector of ones. This weak learner serves as the base component in our boosting framework. The additive property introduced in (4.4) implies that the proposed model is additive with respect to rank increments: increasing the rank of a selected action order is equivalent to concatenating additional weak learners that share the same action order.
To reduce computational complexity, we predefine candidate sets of action orders, denoted by $\mathcal{P}_{y}^{(0)}$ and $\mathcal{P}_{x}^{(0)}$, based on prior knowledge of the data. These sets contain all feasible action orders to be considered for responses and predictors, respectively. For notational convenience, we abuse $K_{y}$ and $K_{x}$ to denote the number of action orders in $\mathcal{P}_{y}^{(0)}$ and $\mathcal{P}_{x}^{(0)}$ in this section.
To track how many times each action order is selected during the boosting process, we define the evolving action sets at iteration $v$ as $\mathcal{P}_{y}^{(v)}=\{\beta_{1}(s_{1}^{(v)}),\ldots,\beta_{K_{y}}(s_{K_{y}}^{(v)})\}$ and $\mathcal{P}_{x}^{(v)}=\{\alpha_{1}(r_{1}^{(v)}),\ldots,\alpha_{K_{x}}(r_{K_{x}}^{(v)})\}$, where $s_{k^{\prime}}^{(v)}$ and $r_{k}^{(v)}$ indicate the number of times action orders $\beta_{k^{\prime}}$ and $\alpha_{k}$ have been selected up to iteration $v$. We set $\mathcal{P}_{y}^{(0)}=\{\beta_{1}(0),\ldots,\beta_{K_{y}}(0)\}$ and $\mathcal{P}_{x}^{(0)}=\{\alpha_{1}(0),\ldots,\alpha_{K_{x}}(0)\}$ for initialization. For each selected action order $\beta_{k^{\prime}}(s_{k^{\prime}}^{(v)})\in\mathcal{P}_{y}^{(v)}$ and $\alpha_{k}(r_{k}^{(v)})\in\mathcal{P}_{x}^{(v)}$, we let $\boldsymbol{s}_{k^{\prime}}^{(v)}=s_{k^{\prime}}^{(v)}\boldsymbol{1}_{N}$ and $\boldsymbol{r}_{k}^{(v)}=r_{k}^{(v)}\boldsymbol{1}_{N}$. These sets and ranks are then used in Algorithm 1 to estimate $\boldsymbol{\Lambda}_{y}^{(v)}$, $\boldsymbol{\Lambda}_{x}^{(v)}$ and $\boldsymbol{\Theta}^{(v)}$, yielding the prediction $\widehat{\boldsymbol{y}}_{t}^{(v)}$.
[…]
$s_{\tilde{k}^{\prime}}^{(v+1)}=s_{\tilde{k}^{\prime}}^{(v)}+1,\quad r_{\tilde{k}}^{(v+1)}=r_{\tilde{k}}^{(v)}+1$
while keeping all other counts unchanged. By iteratively adding weak learners structured according to selected action orders, the model progressively refines its representational capacity in a data-adaptive and interpretable way.
4.2.2 Rank reduction
Algorithm 2 Boosting-Based Algorithm for Hyperparameter Selection
Input: Data $\{(\boldsymbol{y}_{t},\boldsymbol{x}_{t})\}_{t=1}^{T}$, initial sets $\mathcal{P}_{y}^{(0)}$ and $\mathcal{P}_{x}^{(0)}$, and maximum iterations $V_{\max}$.
Initialize: Set $\widehat{\boldsymbol{y}}_{t}^{(0)}=\boldsymbol{0}$ for all $t$ and $v=0$.
repeat
  Evaluate all candidate action order pairs $(\beta_{k^{\prime}},\alpha_{k})\in\{\beta_{1},\ldots,\beta_{K_{y}^{0}}\}\times\{\alpha_{1},\ldots,\alpha_{K_{x}^{0}}\}$:
  for all $(\beta_{k^{\prime}},\alpha_{k})$ do
   Fit weak learner $WL(\beta_{k^{\prime}},\alpha_{k})$ to regress $\boldsymbol{y}_{t}-\widehat{\boldsymbol{y}}_{t}^{(v)}$ on $\boldsymbol{x}_{t}$.
  end for
[…]
  Update action counts: $s_{k^{\prime}}^{(v+1)}=s_{k^{\prime}}^{(v)}+\mathbb{I}\{k^{\prime}=\tilde{k}^{\prime}\}$, $r_{k}^{(v+1)}=r_{k}^{(v)}+\mathbb{I}\{k=\tilde{k}\}$
  Update action sets: $\mathcal{P}_{y}^{(v+1)}=\{\beta_{k^{\prime}}(s_{k^{\prime}}^{(v+1)})\}_{k^{\prime}=1}^{K_{y}}$, $\mathcal{P}_{x}^{(v+1)}=\{\alpha_{k}(r_{k}^{(v+1)})\}_{k=1}^{K_{x}}$
  Refit the full model using Algorithm 1 with $\mathcal{P}_{y}^{(v+1)}$ and $\mathcal{P}_{x}^{(v+1)}$ to obtain $\widehat{\boldsymbol{y}}_{t}^{(v+1)}$.
  $v\leftarrow v+1$
until BIC increases or $v=V_{\max}$
 Perform rank reduction: $(\mathcal{P}_{y},\mathcal{P}_{x})=\text{RR}(\mathcal{P}_{y}^{(v)},\mathcal{P}_{x}^{(v)})$
Return: $\mbox{$\mathscr{A}$}]_{N}$ by applying Algorithm 1 with action sets $\mathcal{P}_{y}$ and $\mathcal{P}_{x}$
According to Proposition 1, multiple combinations of action order sets $(\mathcal{P}_{y}^{(v)},\mathcal{P}_{x}^{(v)})$ may yield equivalent representations of the same model. Among them, we prefer the representation that achieves the lowest model complexity, as measured by BIC defined in (4.3). To this end, we introduce a procedure for selecting the optimal pair $(\mathcal{P}_{y},\mathcal{P}_{x})$ after the action order selection stage. We refer to this procedure as the rank reduction step, denoted by $\text{RR}(\cdot)$.
The rank reduction procedure is applied to both $\mathcal{P}_{x}$ and $\mathcal{P}_{y}$. We illustrate the process using $\mathcal{P}_{x}$ as an example. The procedure iterates sequentially over action orders $\alpha_{k}$ with $k=1,\ldots,K_{x}$. For a given action order $\alpha_{k}$ with associated rank $\boldsymbol{r}_{k}=(r_{1,k},\ldots,r_{N,k})$, we apply Proposition 1 to search for an equivalent but more efficient representation by exploring all alternative action orders within $\mathcal{P}_{x}$.
[…]
An identical procedure is then applied to $\mathcal{P}_{y}$, ensuring both $\boldsymbol{\Lambda}_{x}$ and $\boldsymbol{\Lambda}_{y}$ are represented in their most parsimonious forms. The final output of the rank reduction step is the optimal pair $(\mathcal{P}_{y},\mathcal{P}_{x})$ corresponding to the most efficient, yet equivalent, representation of the model structure. We summarize the overall hyperparameter selection procedure into Algorithm 2.
5 Simulation studies
This section conducts three simulation studies to evaluate the finite-sample performance of our OLS estimator proposed in Algorithm 1, investigate the effects of action order misspecification, as discussed in Proposition 1, and compare the performance of different feature extraction methods for hierarchical data under the framework of supervised factor modeling.
In the first experiment, the data generating process follows (2.12) with $L=2$ and $N=3$ and equal dimension sizes, i.e., $\mbox{$\mathscr{A}$}_{\mathrm{AR}}]_{3}=([\mbox{$\mathscr{A}$}_{1}]_{3},[\mbox{$\mathscr{A}$}_{2}]_{3})$ and $\mbox{$\mathscr{Y}$}_{t}\in\mathbb{R}^{q\times q\times q}$. For simplicity, let intrinsic ranks $r_{n,1}=r_{n,2}=r$, $s_{n,1}=s_{n,2}=s$, where $r=s+1$. We consider three types of sub-Gaussian noise: (i) i.i.d. uniform noise, (ii) i.i.d. standard normal noise, and (iii) correlated gaussian noise. The detailed data generation process is relegated to Supplementary Material.
From Theorem 2, the estimation error converges at a rate of $\sqrt{d_{\mathrm{AR}}/T}$, where $d_{\mathrm{AR}}=2(K_{x}r^{2}+K_{y}s^{2})Nq+2LK_{x}K_{y}rs$. To numerically verify this convergence rate, we consider three settings: (a) Fix $(r,T)=(3,2500)$ and vary $q\in\{10,11,12,13,14\}$; (b) Fix $(q,T)=(10,2500)$ and vary $r\in\{3,4,5,6,7\}$; (c) Fix $(q,r)=(10,3)$ and vary $T\in\{833,1000,1250,1670,2500\}$. In all cases, we apply Algorithm 1 to estimate $\mbox{$\mathscr{A}$}_{\mathrm{AR}}$. Figure 2 presents the estimation error $\|\mbox{$\mathscr{\widehat{A}}$}_{\mathrm{LR}}-\mbox{$\mathscr{A}$}^{*}_{\mathrm{LR}}\|_{\mathrm{F}}$ for these settings, averaged over 200 replications. The observed linearity in Figure 2 confirms the theoretical convergence rate in Theorem 2. It verifies that the estimation error depends on $q$ linearly, and $r$ quadratically, thereby demonstrating the efficiency of our proposed method.
[…]
In the second experiment, we investigate the impact of action order misspecification, as discussed in Proposition 1. We set $M=3$, then the action set is $\mathcal{P}(3)$, which consists of all permutations of $\{1,2,3\}$ with cardinality $|\mathcal{P}(3)|=6$. The data are generated by $\boldsymbol{y}_{t}=\mathcal{M}(\mathcal{T}(\mbox{$\mathscr{X}$}_{t};\alpha_{1});\{\boldsymbol{G}_{m,1}\}_{m=1}^{3})$, where the true action order is $\alpha_{1}=123$. For simplicity, let $r_{1,k}=r_{2,k}=r_{3,k}=r_{k}$ for $k=1,\ldots,6$, and $\boldsymbol{G}_{m,1}$’s are generated similarly as in the first experiment. We consider the i.i.d. normal setting for $\mbox{$\mathscr{X}$}_{t}\in\mathbb{R}^{r_{k}\times r_{k}\times r_{k}}$ with $\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})\sim N(\boldsymbol{0},\boldsymbol{C})$, where $\boldsymbol{C}$ is a random matrix.
By Proposition 1, for each $\alpha_{k}\in\mathcal{P}(3)$, there exists a corresponding set of $\{\boldsymbol{G}_{m,k}\}_{m=1}^{3}$ such that $\mathcal{M}(\mathcal{T}(\mbox{$\mathscr{X}$}_{t};\alpha_{1});\{\boldsymbol{G}_{m,1}\}_{m=1}^{3})=\mathcal{M}(\mathcal{T}(\mbox{$\mathscr{X}$}_{t};\alpha_{k});\{\boldsymbol{G}_{m,k}\}_{m=1}^{3})$, with the cost of potentially larger intrinsic rank $r_{k}$. We apply Algorithm 1 to estimate $\boldsymbol{G}_{m,k}$’s for each $k$, and compute the mean squared error (MSE) of the predictions. Figure 3(a) presents the MSEs for different values of $r_{k}$’s averaged over 50 replications. For both settings of $\mbox{$\mathscr{X}$}_{t}$, we observe that when $r_{k}$ is small, the MSEs for the misspecified action orders are significantly larger than that of the true action order $\alpha_{1}$. However, as $r_{k}$ increases, MSEs for all action orders converge to the same level as that of the true action order $\alpha_{1}$. This aligns with our theoretical findings in Proposition 1.
Figure 3: (a) Average MSEs under different action orders with varying intrinsic rank $r_{k}$. (b) Average rolling forecast MSEs for HOFM, HFM, and our proposed method under supervised factor modeling framework.
In the last experiment, we compare our proposed method with two hierarchical feature extraction methods: HFM (schmid1957development) and HOFM (yung1999relationship), under the framework of SFM. The data generating process follows (2.12) with the coefficient tensor $\mbox{$\mathscr{A}$}_{\mathrm{AR}}$ constructed by both HOFM and our proposed method. Details of the data generation and training procedure are relegated to Supplementary Material.
To compare the one-step ahead forecast performance of HOFM, HFM, and our proposed method, we apply Algorithm 1 to estimate the coefficient tensor $\mbox{$\mathscr{A}$}_{\mathrm{AR}}$ for our model, and use the alternating least squares algorithm to estimate the coefficient tensor for HFM and HOFM. Figure 3(b) presents the average rolling forecast MSEs over 50 replications. We observe that our method consistently outperforms both HFM and HOFM, regardless of whether the data are generated by HOFM or our method. This demonstrating the efficiency of our model in capturing hierarchical structures with shared components within each mode in loading matrices for feature extractions.
6 Real data analysis
In this section, we demonstrate the application of our proposed model to the International Personality Item Pool (IPIP) NEO-120 personality inventory (johnson2014measuring), using data collected from June 1, 2015 to April 30, 2025. The dataset is available in the Hugging Face Datasets repository: https://huggingface.co/datasets/ecorbari/IPIP120-SCORES. It includes scores for the five major personality traits (Neuroticism, Extraversion, Conscientiousness, Agreeableness, and Openness to Experience), assessed by 24 questions. Participants are categorized by gender (two groups) and age (six groups: 10-19, 20-29, 30-39, 40-49, 50-59, and 60+), resulting in 12 demographic groups. To analyze trends in personality traits over time, we compute monthly average scores for each question (recorded on a 1-5 Likert scale) within each demographic group, forming a tensor time series $\{\mbox{$\mathscr{Y}$}_{t}\in\mathbb{R}^{12\times 5\times 24}\}_{t=1}^{T}$ with $T=115$ observations. There is one missing entry in the dataset, which we impute using spline interpolation. For preprocessing, the data are first detrended by first-order differencing, and then standardized to have zero means.
The tensor autoregression model (2.12) is employed with each coefficient tensor $\mbox{$\mathscr{A}$}_{l}\in\mathbb{R}^{12\times 5\times 24\times 12\times 5\times 24}$. Observations are divided into a training set with the first 103 months and a testing set with the last 12 months. To enable a feasible estimation, we consider our proposed model, SFM based on CP and TT decompositions, and HOFM under SFM frameworks for feature extraction with the lag order fixed at $L=1$ by jointly consider the PACFs of all entries in data. We omit the comparison with Tucker decomposition and HFM based on SFM frameworks due to thier large number of parameters in the coefficient matrix, which makes them infeasible for this data analysis. For our proposed model, from prior knowledge of the data, we set the candidate action order sets as $\mathcal{P}_{x}^{(0)}=\mathcal{P}_{y}^{(0)}=\{(3,2,1),(1,3,2)\}$, i.e., the features are extracted by the order of questions, personalities, and demographic groups, or the order of demographic groups, questions, and personalities. We then apply Algorithm 2 to get the selected intrinsic ranks $\boldsymbol{r}_{1}=(1,1,2)$, $\boldsymbol{r}_{2}=(9,1,2)$, $\boldsymbol{s}_{1}=(3,1,2)$, and $\boldsymbol{s}_{2}=(3,2,2)$. For the two tensor decomposition based SFMs, we adopt the ALS methods in zhou2013tensor for CP-based tensor autoregression, and wang2022high for Tucker-based tensor autoregression. For TT-based SFM, we follow si2024efficient and qin2025computational to perform Riemannian gradient descent for estimation. The intrinsic ranks for these methods are selected by their respective methods.
We next evaluate the rolling forecast performance on the testing set of these four methods, i.e., fit the model to historical data up to month $t$ to forecast the data at month $t+1$ for $t=103,\ldots,114$. Specifically, we use mean squared forecast errors (MSFE) and mean absolute forecast errors (MAFE) to assess the one-step ahead forecast accuracy. Table 2 summarizes the results, where our proposed method outperforms all competing methods in both MSFE and MAFE metrics. This result supports the existence of hierarchical structures in the IPIP NEO-120 data, which can be effectively captured by our proposed model. Moreover, the superior performance of our method over TT decomposition based SFM highlights the advantage of our model by considering multiple action orders for general hierarchical data. And the better performance compared to HOFM based SFM demonstrates the benefit of improved efficiency in feature extraction by sharing components within each mode in loading matrices. In conclusion, our proposed model provides an efficient and interpretable approach for modeling tensor time series data with hierarchical structures.
Table 2: Mean squared forecast errors (MSFE) and mean absolute forecast errors (MAFE) of our model and three competing methods on the testing set of the IPIP NEO-120 data. The smallest numbers in each row are in bold.
Ours CP $T$ HOFM
MSFE 138.78 142.28 145.73 142.15
MSAE 279.28 280.63 283.43 280.58
Finally, we zoom in on the estimated coefficient tensor from our proposed model to interpret the hierarchical structures in the IPIP NEO-120 data and the importance of different action orders. We discuss the component matrices involved in feature extraction for responses under action order $(1,3,2)$ as an example. Figure 4(a) presents the heatmap of the component matrix $\boldsymbol{H}_{1,1}^{\top}$, which correspond to the first step of feature extraction, summarizing the information across the 24 questions for each personality trait. In addition, Figure 4(b) displays $\boldsymbol{H}_{2,1}^{\top}$ that represents the combined effects of personality traits and questions by a single extracted feature. Lastly, Figure 4(c) illustrates $\boldsymbol{H}_{3,1}^{\top}$, capturing the interactions between demographic groups and the combined features from questions and personality traits. The response tensor $\mbox{$\mathscr{Y}$}_{t}$ is then summarized into two factors as its low-dimensional representation. Component matrices with the second action order $(1,3,2)$ for responses and those for predictors with both action orders can be found in Supplementary Material, which have similar interpretations as those for predictors. Finally, Figure 4(d) presents the heatmap of $\boldsymbol{\Theta}$, the estimated coefficient matrix that captures the relationships between the extracted factors from predictors and responses by two action orders. We observe that features extracted by action order $(3,2,1)$ have a more significant impact on extracted features from responses under both action orders compared to those extracted by action order $(1,3,2)$, as indicated by the larger absolute values in the corresponding blocks of $\boldsymbol{\Theta}$.
Figure 4: Heatmaps of (a) $\boldsymbol{H}_{1,1}^{\top}$, (b) $\boldsymbol{H}_{2,1}^{\top}$, (c) $\boldsymbol{H}_{3,1}^{\top}$, and (d) $\boldsymbol{\Theta}$.
7 Conclusion and discussion
In this paper, we introduce a novel autoregressive model under SFM framework for tensor time series data that effectively captures hierarchical structures through the incorporation of multiple action orders in the construction of the coefficient tensor. We employ a structure of shared components within each mode in loading matrices for efficient feature extraction. An alternating least squares algorithm is developed for parameter estimation, and a boosting-based method is proposed for hyperparameter selection, which aims to reveal the underlying hierarchical structures present in the data. Theoretical analyses demonstrate the consistency of our estimator under sub-Gaussian noise with misspecification of action orders. Simulation studies and real data analysis illustrate the superior performance of our method compared to existing approaches.
For future research, we may extend the proposed model to accommodate more complex time series structures, such as moving average (MA) or ARMA models. We can also develop robust estimation techniques for handling heavy-tailed noise distributions, so that our methods can be deployed to a wider range of real-world applications. Additionally, investigating statistical inference procedures for the selected action orders and intrinsic ranks would provide valuable insights into the interpretability and reliability of the model in practical settings.

### v2

@@
-High-dimensional Autoregressive Modeling for Time Series with Hierarchical Structures
+High-dimensional Autoregressive Modeling for Time Series Data with Hierarchical Structures
 Abstract
-High-dimensional time series often exhibit hierarchical structures represented by tensors, while statistical methodologies that can effectively exploit the structural information remain limited.
-We propose a supervised factor modeling framework that accommodates general hierarchical structures by extracting low-dimensional features sequentially in the mode orders that respect the hierarchical structure.
-Our method can select a small collection of such orders to allow for impurities in the hierarchical structures, yielding interpretable loading matrices that preserve the hierarchical relationships.
-A practical estimation procedure is proposed, with a hyperparameter selection scheme that identifies a parsimonious set of action orders and interim ranks, thereby revealing the possibly latent hierarchical structures.
-Theoretically, non‑asymptotic error bounds are derived for the proposed estimators in both regression and autoregressive settings.
-An application to the IPIP‑NEO‑120 personality panel illustrates superior forecasting performance and clearer structural interpretation compared with existing methods based on tensor decompositions and hierarchical factor analysis.
-Keywords: Autoregressive model, Hierarchical tensor data, High-dimensional time series, Non-asymptotic properties, Supervised factor modeling
+Modern applications have made ubiquitous high-dimensional data, especially time-dependent data, with more and more complicated structures, and it also has become more frequent to encounter the scenario of hierarchical relationships among variables.
+However, there is still a lack of supervised learning tool in the literature for them.
+To fill this gap, we introduce a new model-designing framework, and it then combines with unsupervised factor modeling tools to form an efficient and interpretable autoregressive model for high-dimensional time series with hierarchical structures.
+An ordinary least squares estimation is considered, and its non-asymptotic properties are established.
+Moreover, we propose an algorithm to search for estimates, and a boosting method is also suggested for hyperparameter selection.
+Simulation experiments are conducted to evaluate finite-sample performance of the proposed methodology, and its usefulness is demonstrated by an application to the Personality‑120 dataset.
+Abstract
+This supplementary material provides technical details and additional empirical evidence for the main paper.
+We first introduce the tensor notation and algebra used throughout the paper in Section A, and then present additional results on model properties in Section B, including proofs of propositions, derivations of the closed-form ALS update rules, and the SSVD procedure used for orthogonalization.
+We also provide complete proofs of the main theoretical guarantees in Section C, together with auxiliary lemmas in Section D.
+Finally, we include additional details for the simulation studies in Section E and further empirical results for the Personality-120 data analyses in Section F, which complement the findings reported in the main text.
+Keywords: Autoregressive model, Hierarchical tensor data, High-dimensional time series, Non-asymptotic properties, Tensor decomposition
 1 Introduction
-High-dimensional time series abound in modern scientific and econometric studies (wang2019factor; chang2024autocovariance; chen2025estimating), which often exhibit multiway (chen2022analysis; chen2022factor; wang2024high) and partially hierarchical organizations (johnson2014measuring; gao2023divide).
-The data are hierarchical when they contain multiple modes representing the layers of a nested structure obtained from several groups, subgroups, and items.
-In many applications, however, the hierarchical structure is only partial as some modes lie outside this hierarchy or even form another hierarchy.
-For balanced data which admits a tensor representation, methods that can adapt to such partially hierarchical tensor time series remain limited, as existing tensor-based methods or high-dimensional dynamic factor models typically assume a fully juxtaposing or hierarchical organization (montgomery2006design; stock2011dynamic; brunner2012tutorial; yuan2023two).
-In practice, real datasets frequently combine both: hierarchical components accompanied by cross-cutting covariates.
-Developing a framework that can efficiently summarize, predict, and statistically interpret such partially hierarchical structures with complex inter-mode dependencies is therefore of both theoretical and practical importance.
-A representative example is the International Personality Item Pool (IPIP) NEO-120 personality inventory (johnson2014measuring; qiao2025exploratory), which contains 120 questions assessing five broad personality traits, each with 24 questions.
-The data are collected over time from participants of different ages and genders, resulting in a three-way tensor time series with dimensions $12\penalty 10000\ (gender\penalty 10000\ and\penalty 10000\ age\penalty 10000\ groups)\times 5\penalty 10000\ (personalities)\times 24\penalty 10000\ (questions)$.
-The last two dimensions reflect a hierarchical organization, while the first does not belong to this hierarchy.
-Given the dataset’s scale ($12\times 5\times 24=1440$ dimensions) and the limited number of time points (approximately 100 monthly observations), there is a need for an efficient and statistically interpretable model that can extract latent features from the data for prediction of future observations.
-The extraction process which reduces the dimensionality in a supervised manner should respect the data’s partially hierarchical structure, when both latent feature detection and trend analysis are of sociological interest.
-Natural strategies to handle data of such high dimensionality include imposing sparsity or low-rankness on the model, so that a more tractable low-dimensional structure emerges.
-Sparsity-inducing methods (basu2015regularized; raskutti2019convex; bertsimas2020sparse) have been widely used in high-dimensional time series analysis with its well-known prediction ability (wilms2023sparse; zheng2025interpretable).
-However, these methods may lack structural interpretability unless the underlying sparsity pattern aligns with a meaningful structure, as is common in genetics and neuroscience (zhou2013tensor; bi2022modeling; wu2023sparse).
-In IPIP-NEO-120, for instance, the responses within each personality trait are highly correlated due to the experimental design, and a sparse model may fail to capture this inherent structure.
-In contrast, low-rank modeling compresses the data to capture latent structure via orthogonal loading matrices, leading to a more interpretable framework of factor modeling and analysis.
-Supervised Factor Modeling (SFM) (huang2025supervised) is a prominent low-rank approach that uses loading matrices to extract features (or factors) from both predictors and responses, followed by a multivariate regression on these features.
-Specifically for hierarchical data, literature is copious with methods that extract features tailored to the hierarchical construct, such as higher-order factor models (yung1999relationship), group factor models (klami2014group), hierarchical factor models (schmid1957development; qiao2025exact).
-These methods either extracts factors recursively along the hierarchy, or directly from each of its levels.
-However, they all assign different loading matrices to different lower-level modes under the same higher-level mode, so the number of parameters exceed the dimensions of the data.
-These approaches are thus only feasible in the unsupervised context, where the proportion of explained variance, rather than prediction performance, is the primary concern.
-When otherwise applied in the supervised context, only low dimensional settings are allowed.
-As traditional low-rank methods for unsupervised factor analysis of hierarchical constructs are unsuitable for high-dimensional estimation problem, SFM methods based on tensor decomposition (huang2025supervised) are preferred given their parsimony.
-Among these, Tucker and canonical polyadic (CP) decompositions are widely used (kolda2009tensor).
-Tucker decomposition (tucker1966some) is notable for its numerical stability and interpretable orthogonal loadings, which align well with supervised factor modeling (wang2024high).
-However, its model complexity grows exponentially with tensor order, limiting its efficiency in high-order settings (oseledets2011tensor).
-This issue is exacerbated in tensor autoregressive models, where coefficient tensors often have high orders.
-In the IPIP-NEO-120 example, a lag-one tensor autoregressive model would involve a sixth-order coefficient tensor, and estimation based on Tucker decomposition would be computationally less efficient.
-Additionally, Tucker loadings are confined on Kronecker product spaces, restricting its applicability to data with complex or hierarchical structures, as discussed in Sections 2.1 and 2.2.
-CP decomposition (harshman1970foundations) offers a more efficient and flexible alternative (lock2018tensor), with model complexity scaling linearly with tensor order and dimensions.
-However, CP loadings are complicated, constraining its interpretation in unsupervised settings as a PCA surrogate (han2024cp; han2022tensor), and CP-based SFM (cai2025efficient) lacks clear interpretation for structured data.
-Table 1: Comparison of different methods for supervised factor modeling.
-For simplicity, we assume equal dimension $p$ for each mode of the data, and equal number of features, $r$, extracted from each mode, and denote $d$ as the order of the coefficient tensor.
-Models SFM based on Tensor Decomp. Factor Models for SFM
-Ours CP Tucker $T$ GFM HFM HOFM
-Number of parameters: $O(\cdot)$ $pr^{2}d$ prdprd $prd+r^{d}$ $pr^{2}d$ $rp^{d}$ $rp^{d}$ $rp^{d}$
-Interpretation: Hierarchical $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-Interpretation: Crossed $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$
-Meanwhile, tensor regression based on tensor train (TT) decomposition (oseledets2011tensor) of the coefficient tensor has been proposed with its efficiency highlighted for tensors with high orders (liu2020low).
-Under the SFM structure, qin2025computational and si2024efficient adopt a pair of TT-styled loadings for the predictor and repsonse, and their sequential structure can capture hierarchical relationships among modes with reduced number of parameters.
-Specifically, each loading matrix is a product of several component matrices, with each component extracting features from one mode to which the previous mode imparts extracted features.
-The TT-based model is hence interpretable in the regression tasks where both the predictor and response exhibit a fully hierarchical structure (zhou2022optimal; si2024efficient).
-However, its strict sequential structure inherited from TT decomposition limits its applicability to data with partially hierarchical structures, as in the IPIP-NEO-120 example.
-Another major limitation is that the order of the component matrices in TT loadings is fixed to be forward for the predictor and backward for the response.
-This fixed construction hinders its application to autoregressive time series models, where the predictor and response share basically the same structure.
-Permuting the order of the predictor’s modes for a forced structural alignment in turn deteriorates the autoregressive model into a general stochastic regression model, complicating the temporal dependence analysis.
-Table 1 provides a summary of the reviewed methods under the SFM framework, comparing their number of parameters and interpretability for hierarchical and crossed (non-hierarchical) structures.
-In view of the aforementioned challenges, we propose a novel tensor-based supervised factor modeling framework that effectively extracts features from partially hierarchical tensor time series, while preserving the hierarchical structure information.
-Our model combines admissible mode orders of the partially hierarchical structure instrinc to the data, analogous to a mixture model.
-Then, it extracts features sequentially, thereby capturing complex inter-mode dependencies with cross-cutting covariates obscuring any exact hierarchical structure.
-The proposed method is computationally efficient, as information summarization and its upward impartation to the higher-level layer bear a shared low-rank structure for the lower-level layer.
-This parsimony enables both accurate estimation with sample size and statistical interpretability inherited from SFM framework.
-Specifically, our contributions are summarized as follows:
-  • We are, to the best of our knowledge, the first to bridge the methodological gap between fully hierarchical structures and non-hierarchical (juxtaposing) structures.
-Our proposed novel tensor-based SFM method accommodates to partially hierarchical structures in both regression and autoregression tasks, with theoretical guarantees evidenced by both simulation and empirical examples.
-  • Our proposed method offers a flexible and unified way to detect from data principal and lesser hierarchical structures forming a mixture, enabled by our hyperparameter selection algorithm, when the true structure is obscured.
-  • The construction in our proposed method can reveal an algebraic-graphical duality of the hierarchical structures, enshrined in an algebraic proposition, which may be of independent interest in the study of tensor decomposition and hierarchical models.
-The rest of the paper is organized as follows.
-In Section 2, we introduce the general supervised factor modeling framework, and then present our proposed method for partially hierarchical data, followed by an interpretable tensor autoregressive model.
-Section 3 provides the theoretical analysis of our method and Section 4 describes its algorithms along with hyperparameter selection.
-In Section 5, we conduct extensive simulation studies to evaluate the performance of our method against existing alternatives.
-In Section 6, we apply our method to the IPIP-NEO-120 dataset for analysis.
-Finally, we conclude with a discussion in Section 7.
+With the rapid advancement of information technology, multi-dimensional or tensor-valued data are now ubiquitous, and particularly many of them are time-dependent.
+Examples can be found in medical imaging (Zhou et al., 2013; Li and Zhang, 2021), digital marketing (Bi et al., 2018; Hao et al., 2021), economics and finance (Chen et al., 2022; Wang et al., 2024) and psychology (Tang et al., 2018).
+It is arguably the most important task to conduct supervised learning on these data, say tensor regression (Lock, 2018; Raskutti et al., 2019) and tensor autoregression (Li and Xiao, 2021; Wang et al., 2024), while existing methods typically treat all modes in an equal manner, thus mainly for the data with fully juxtaposing structures.
+However, for these observed tensors, some modes may be nested within others, forming hierarchical relationships, and they can be encountered more readily for the data with a larger number of modes (Gao and Tsay, 2023; Qiao et al., 2025a).
+Meanwhile, it is still an open problem to model and then to predict these high-dimensional tensor-valued data, especially time-dependent data, with hierarchical structures.
+As an illustrative example, we consider the Personality-120 dataset, which contains 24 questions for each of five broad personality traits, ending up with 120 questions in total.
+For any personality trait, each of its 24 questions is designed to quantitatively assess a specific psychological aspect, and the answers to those 24 questions jointly can be used to roughly portray an individual’s personality under this trait; see more empirical examples in psychometrics (Johnson, 2014; Qiao et al., 2025b).
+In addition, the dataset is collected over time from participants of different ages and genders, forming 12 demographic groups.
+As a result, at each time point, there is an observed three-mode tensor with dimensions of $12\penalty 10000\ (\text{demographic groups})\times 5\penalty 10000\ (\text{personality traits})\times 24\penalty 10000\ (\text{questions})$, where the last two modes reflect a hierarchical structure (questions nested within traits), while the 24 questions under each personality trait are arranged in a similar way.
+Moreover, the first mode (demographic groups) represents a cross-cutting covariate that does not belong to this hierarchy.
+It is challenging to construct an efficient time series model for the above example since it has $12\times 5\times 24=1440$ variables yet with only around 100 time points.
+On the other hand, to achieve an appropriate physical interpretation, another layer of difficulties is imposed due to the inherent hierarchical structure between personality traits and questions.
+As arguably the most widely used model, autoregression (AR) has been a primary workhorse for many high-dimensional tasks on time series with the observed vector, matrix or tensor at each time point (Basu and Michailidis, 2015; Chen et al., 2021; Wang et al., 2024).
+The general AR model has $O(P^{2})$ parameters, where $P$ is the number of variables, say $P=1440$ in the Personality-120 example, and it hence is necessary to restrict the parameter space such that a reliable estimation can be achieved.
+A direct approach in the literature is to assume that the coefficient matrices or tensors are sparse and then to apply sparsity-inducing regularized estimation (Basu and Michailidis, 2015; Wilms et al., 2023; Zheng, 2025).
+The sparse AR model has been widely used, say in genetics and neuroscience (Zhou et al., 2013; Bi et al., 2022; Wu and Feng, 2023), where dependence among variables is believed to be sparse.
+However, for the Personality-120 example, there exists strong dependence among all $P=1440$ variables, especially each 24 variables from the same personality trait, and it is more reasonable to assume that these variables are driven by a small number of common latent factors (Qiao et al., 2025a).
+Similar scenario can be found for macroeconomic variables, where the factor modeling is usually considered (Lam and Yao, 2012; Bai and Wang, 2016).
+Another important method for dimension reduction is to impose low-rank structures to coefficient matrices (Negahban and Wainwright, 2011) or tensors (Han et al., 2022) and, as a most commonly used tools for compressing tensors (Kolda and Bader, 2009), Tucker decomposition (Tucker, 1966) can be directly applied to the coefficient tensor of AR models (Wang et al., 2022, 2024; Huang et al., 2025).
+On the one hand, although the corresponding dimension of parameter spaces can be reduced dramatically, Tucker decomposition is well known not to be able to effectively compress higher-mode tensors, and it loses the efficiency very quickly as the mode of tensor increases, say greater than four or five; see Oseledets (2011).
+Unfortunately, consider an AR model for tensor-valued time series with mode three, and the mode of its coefficient tensor will be six.
+On the other hand, together with the regression setting, Tucker decomposition will lead to a nice interpretation of supervised factor modeling (Wang et al., 2024; Huang et al., 2025).
+Specifically, it is equivalent to first projecting responses and predictors into a small number of latent factors, termed response and predictor factors, respectively, and then there exists a regression relationship between the two types of factors.
+The response factors are used to summarize all predictable components, while the predictor ones contain all useful information of predictors in terms of prediction.
+However, the above physical interpretation will disappear for the data with hierarchical structures, such as the Personality-120 example; see Section 2.1 for details.
+CP (Harshman and others, 1970) and tensor train (Oseledets, 2011) decompositions are another two popular methods for compressing tensors, and both of them are more efficient than Tucker decomposition (Kolda and Bader, 2009).
+However, it is difficult to interpret fitted models when CP decomposition is applied to coefficient tensors of tensor-on-tensor regression (Lock, 2018) and tensor autoregression.
+Tensor train (TT) decomposition may still lead to a supervised factor modeling interpretation when it is applied to tensor regression (Liu et al., 2020; Qin and Zhu, 2025) and autoregression (Si et al., 2024), while it is limited to data with a very special structure; see Remark 2 for details.
+To the best of our knowledge, for tensor-valued data, especially time-dependent data, with hierarchical structures, it is infeasible to achieve an efficient model with physical interpretations by directly applying an existing tensor decomposition method to coefficient tensors of the corresponding model.
+Motivated by the supervised factor modeling interpretation of Tucker tensor regression, this paper introduces a new framework for constructing tensor regression and autoregressive models.
+Specifically, two factor modeling approaches are first applied to responses and predictors to extract features, respectively, and a low-rank structure is then imposed to the coefficient tensor so that it admits a regression relationship between response and predictor features.
+We may even employ different factor modeling methods for responses and predictors; see Section 2.1 for more details.
+Note that, for the data with general hierarchical structures, traditional factor modeling methods can be used to extract features or factors; see, for example, higher-order factor models (Yung et al., 1999), group factor models (Klami et al., 2014), and hierarchical factor models (Schmid and Leiman, 1957; Qiao et al., 2025a).
+These tools are unsupervised learning methods, and they admit various hierarchical interpretations by either extracting features recursively along the hierarchy, or directly from different groups of variables.
+We can apply them to construct tensor regression and autoregressive models within the new framework, while the loading matrices assigned to each lower-level mode are all different, leading to a huge number of parameters and hence excluding high-dimensional settings; see Section 2.2 for details.
+On the other hand, for many recent hierarchical data such as the Personality-120 example, they are arranged in tensor forms, i.e., each mode has the same number of variables.
+Moreover, the variables at each mode are also arranged in similar ways, and the corresponding factors may be formed with same loading matrices.
+Table 1: Comparison of new-framework models with the higher-order factor model (HOFM), group factor model (GFM) or hierarchical factor model (HFM) being used to extract features, in terms of model complexity and interpretability for hierarchical data.
+The traditional framework refers to the case with tensor decomposition methods being directly applied.
+The coefficient tensor has $d$ modes, and each has the dimension of $p$ and rank of $r$.
+Models New framework Traditional framework
+Ours HOFM GFM HFM CP Tucker $T$
+Number of parameters: $O(\cdot)$ $pr^{2}d$ $rp^{d}$ $rp^{d}$ $rp^{d}$ prdprd $prd+r^{d}$ $pr^{2}d$
+Interpretation: Hierarchical $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$
+Interpretation: Crossed $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+The first contribution of this paper is to overcome the difficulty of traditional factor modeling tools for hierarchical data by considering shared loading matrices within each layer of the hierarchy; see Table 1 for a comparison of their model complexity and interpretability for hierarchical data.
+Moreover, the real-world hierarchical data most likely have a mixture structure, i.e., some modes are nested, while others are crossed.
+We hence apply the new modeling framework to construct an efficient and interpretable regression model for these data with high dimensions in Section 2.3, and it is further extended to AR models for time-dependent data in Section 2.4.
+The second contribution is to consider an ordinary least squares estimation in Section 3, and its non-asymptotic properties are established.
+This is non-trivial since the corresponding loading matrices are non-orthogonal and, as a cost, an additional term of $\log T$ is added to the corresponding estimation error bound.
+Finally, Section 4 proposes an algorithm to search for the estimates, and a boosting-based method is introduced for hyperparameter selection due to the presence of hierarchical structures.
+In addition, Section 5 conducts simulation experiments to evaluate numerical performance of the proposed methodology, and it is further applied to the Personality-120 dataset in Section 6.
+A short conclusion and discussion is given in Section 7.
 All technical proofs and additional numerical results are relegated to the Supplementary Material.
-Notations and preliminaries.
 Throughout the paper, we denote vectors by boldface small letters, e.g., $\boldsymbol{x}$; matrices by boldface capital letters, e.g., $\boldsymbol{X}$; tensors by boldface Euler capital letters, e.g., $\mathscr{X}$.
@@
 The Frobenius norm of $\mathscr{X}$ is defined as $\left\lVert\mbox{$\mathscr{X}$}\right\rVert_{\mathrm{F}}=\sqrt{\langle\mbox{$\mathscr{X}$},\mbox{$\mathscr{X}$}\rangle}$, where $\langle\cdot,\cdot\rangle$ is the inner product.
-Moreover, the $(i_{1}i_{2}\ldots i_{d})$-th element of $\mathscr{X}$ is denoted by $\mbox{$\mathscr{X}$}_{i_{1}i_{2}\ldots i_{d}}$, for $i_{k}=1,2,\ldots,p_{k}$, $k=1,2,\ldots,d$.
-We denote $\times_{s}$ the $s$-th mode matrix product, and $\otimes$ denotes the Kronecker product.
+Moreover, the $(i_{1}i_{2}\ldots i_{d})$-th element of $\mathscr{X}$ is denoted by $\mbox{$\mathscr{X}$}_{i_{1}i_{2}\ldots i_{d}}$, with $1\leq k\leq d$ and $1\leq i_{k}\leq p_{k}$.
+We denote $\times_{s}$ the $s$-th mode matrix product, and $\otimes$ is the Kronecker product.
 For a generic matrix $\boldsymbol{X}$, we denote by $\boldsymbol{X}^{\top}$, $\left\lVert\boldsymbol{X}\right\rVert_{\mathrm{F}}$, $\left\lVert\boldsymbol{X}\right\rVert_{2}$, and $\operatorname*{vec}(\boldsymbol{X})$ its transpose, Frobenius norm, spectral norm, and a long vector obtained by stacking all its columns, respectively.
-If $\boldsymbol{X}$ is further a square matrix, then denote its minimum and maximum eigenvalue by $\lambda_{\min}(\boldsymbol{X})$ and $\lambda_{\max}(\boldsymbol{X})$, respectively.
-Denote the orthogonal set as $\mathbb{O}^{p\times r}=\{\boldsymbol{X}\in\mathbb{R}^{p\times r}:\boldsymbol{X}^{\top}\boldsymbol{X}=\boldsymbol{I}_{r}\}$ for $p\geq r$, where $\boldsymbol{I}_{r}\in\mathbb{R}^{r\times r}$ is the identity matrix.
-With a little abuse of terminology, a rectangular matrix is said to be orthogonal if it belongs to $\mathbb{O}^{p\times r}$.
-For two real-valued sequences $x_{k}$ and $y_{k}$, $x_{k}\gtrsim y_{k}$ if there exists a $C>0$ such that $x_{k}\geq Cy_{k}$ for all $k$.
-In addition, we write $x_{k}\asymp y_{k}$ if $x_{k}\gtrsim y_{k}$ and $y_{k}\gtrsim x_{k}$.
+If $\boldsymbol{X}$ is further a square matrix, we then denote its minimum and maximum eigenvalue by $\lambda_{\min}(\boldsymbol{X})$ and $\lambda_{\max}(\boldsymbol{X})$, respectively.
+Denote a set of orthonormal matrices by $\mathbb{O}^{p\times r}=\{\boldsymbol{X}\in\mathbb{R}^{p\times r}:\boldsymbol{X}^{\top}\boldsymbol{X}=\boldsymbol{I}_{r}\}$ for $p\geq r$, where $\boldsymbol{I}_{r}\in\mathbb{R}^{r\times r}$ is the identity matrix.
+For two real-valued sequences $x_{k}$ and $y_{k}$, we denote $x_{k}\gtrsim y_{k}$ or $x_{k}\lesssim y_{k}$ if there exists a $C>0$ such that $x_{k}\geq Cy_{k}$ or $x_{k}\leq Cy_{k}$ for all $k$, respectively.
+In addition, $x_{k}\asymp y_{k}$ if $x_{k}\gtrsim y_{k}$ and $x_{k}\lesssim y_{k}$.
 2 Model settings
@@
 Note that the Tucker decomposition at (2.2) is not unique, since $\mbox{$\mathscr{A}$}=\mbox{$\mathscr{G}$}\times_{1}\boldsymbol{U}_{1}\times_{2}\cdots\times_{N+M}\boldsymbol{U}_{N+M}=\left(\mbox{$\mathscr{G}$}\times_{1}\boldsymbol{O}_{1}\times_{2}\cdots\times_{N+M}\boldsymbol{O}_{N+M}\right)\times_{1}\left(\boldsymbol{U}_{1}\boldsymbol{O}_{1}^{-1}\right)\times_{2}\cdots\times_{N+M}\left(\boldsymbol{U}_{N+M}\boldsymbol{O}_{N+M}^{-1}\right)$ for any invertible matrices $\boldsymbol{O}_{i}\in\mathbb{R}^{s_{i}\times s_{i}}$ with $1\leq i\leq N+M$.
-Without loss of generality, we restrict these factor matrices to be orthonormal, i.e., $\boldsymbol{U}_{j}\in\mathbb{O}^{q_{j}\times s_{j}}$ for each $1\leq j\leq N+M$; see, for example, the higher-order singular value decomposition (HOSVD) in kolda2009tensor.
+Without loss of generality, we restrict these factor matrices to be orthonormal, i.e., $\boldsymbol{U}_{j}\in\mathbb{O}^{q_{j}\times s_{j}}$ for each $1\leq j\leq N+M$; see, for example, the higher-order singular value decomposition (HOSVD) in Kolda and Bader (2009).
 Consequently, the Tucker tensor regression model can be rewritten into
@@
 indicating that $\text{col}(\boldsymbol{V})$ contains all useful information of predictors in terms of predicting responses.
-As a result, define $\boldsymbol{f}_{t}^{\mathrm{response}}=\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t}\times_{n=1}^{N}\boldsymbol{U}_{n}^{\top})=\boldsymbol{U}^{\top}\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})\in\mathbb{R}^{s}$ and $\boldsymbol{f}_{t}^{\mathrm{predictor}}=\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t}\times_{m=1}^{M}\boldsymbol{V}_{m}^{\top})=\boldsymbol{V}^{\top}\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})\in\mathbb{R}^{r}$ to be the response and predictor factors, respectively, and then the Tucker tensor regression can be well interpreted from a supervised factor modeling perspective (wang2022high; wang2024high; huang2025supervised).
-The above supervised factor modeling interpretation also motivates us to introduce a new framework to construct a tensor regression model, especially for data with complex structures beyond the capability of Tucker decomposition, or for non-orthogonal loadings in the feature extraction process such as the CP decomposition (harshman1970foundations).
+As a result, define $\boldsymbol{f}_{t}^{\mathrm{response}}=\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t}\times_{n=1}^{N}\boldsymbol{U}_{n}^{\top})=\boldsymbol{U}^{\top}\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})\in\mathbb{R}^{s}$ and $\boldsymbol{f}_{t}^{\mathrm{predictor}}=\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t}\times_{m=1}^{M}\boldsymbol{V}_{m}^{\top})=\boldsymbol{V}^{\top}\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})\in\mathbb{R}^{r}$ to be the response and predictor factors, respectively, and then the Tucker tensor regression can be well interpreted from a supervised factor modeling perspective (Wang et al., 2022, 2024; Huang et al., 2025).
+The above supervised factor modeling interpretation also motivates us to introduce a new framework to construct a tensor regression model, especially for data with complex structures beyond the capability of Tucker decomposition, or for non-orthogonal loadings in the feature extraction process such as the CP decomposition (Harshman and others, 1970).
 Specifically, we first extract features or factors from responses and predictors by
@@
 respectively.
-Here, $\mathcal{M}(\cdot):\mathbb{R}^{p_{1}\times\cdots\times p_{M}}\to\mathbb{R}^{r}$ and $\mathcal{N}(\cdot):\mathbb{R}^{q_{1}\times\cdots\times q_{N}}\to\mathbb{R}^{s}$ are feature-extracting operators, with $r$ and $s$ being the number of extracted predictor and response features, respectively.
-A specific structure is then assumed to the coefficient tensor $\mathscr{A}$ of model (2.1) such that $\boldsymbol{f}_{t}^{\mathrm{response}}=\boldsymbol{\Theta}\boldsymbol{f}_{t}^{\mathrm{predictor}}+\boldsymbol{e}_{t}$, where $\boldsymbol{\Theta}\in\mathbb{R}^{s\times r}$ and $\boldsymbol{e}_{t}\in\mathbb{R}^{s}$ are the corresponding coefficient matrix and compressed error term, respectively.
+Here, $\mathcal{M}(\cdot):\mathbb{R}^{p_{1}\times\cdots\times p_{M}}\to\mathbb{R}^{r}$ and $\mathcal{N}(\cdot):\mathbb{R}^{q_{1}\times\cdots\times q_{N}}\to\mathbb{R}^{s}$ are feature-extracting operators, with $r$ and $s$ being the number of extracted predictor and response features, respectively, and they do not necessarily belong to the same type.
+A specific structure is then assumed to the coefficient tensor $\mathscr{A}$ of model (2.1) such that $\boldsymbol{f}_{t}^{\mathrm{response}}=\boldsymbol{\Theta}\boldsymbol{f}_{t}^{\mathrm{predictor}}+\widetilde{\boldsymbol{e}}_{t}$, where $\boldsymbol{\Theta}\in\mathbb{R}^{s\times r}$ and $\widetilde{\boldsymbol{e}}_{t}\in\mathbb{R}^{s}$ are the corresponding coefficient matrix and compressed error term, respectively.
 To this end, if operators $\mathcal{M}$ and $\mathcal{N}$ are linear, i.e., $\mathcal{M}(\mbox{$\mathscr{X}$}_{t})=\boldsymbol{\Lambda}_{x}^{\top}\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})$ and $\mathcal{N}(\mbox{$\mathscr{Y}$}_{t})=\boldsymbol{\Lambda}_{y}^{\top}\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})$, with non-degenerating loading matrices $\boldsymbol{\Lambda}_{x}\in\mathbb{R}^{P\times r}$ and $\boldsymbol{\Lambda}_{y}\in\mathbb{R}^{Q\times s}$, we then can set the coefficient tensor to
@@
 2.2 Purely hierarchical structure
-According to technological advances, the available multi-dimensional or tensor-valued data have become more and more complex, and it is very common to observe hierarchical relationships among some modes.
+According to technological advances, the available multi-dimensional or tensor-valued data have become more and more complex, and it is very common to observe hierarchical relationships among some modes with their variables being arranged in a similar way.
 In order to exploit their structural information, we attempt to design a regression model using the proposed framework at (2.4) and (2.5).
@@
 We first consider the feature extraction on predictors, and assume for simplicity that $M=2$, i.e., $\mbox{$\mathscr{X}$}_{t}\in\mathbb{R}^{p_{1}\times p_{2}}$, and the $p_{1}$ variables of the first mode are nested within each of the $p_{2}$ variables in its second mode.
-There are two commonly used methods in the literature of factor analysis to extract features from such hierarchical data $\mbox{$\mathscr{X}$}_{t}$: the hierarchical factor model (HFM) in schmid1957development and higher-order factor model (HOFM) in yung1999relationship.
+There are two commonly used methods in the literature of factor analysis to extract features from such hierarchical data $\mbox{$\mathscr{X}$}_{t}$: the hierarchical factor model (HFM) in Schmid and Leiman (1957) and higher-order factor model (HOFM) in Yung et al.
+(1999).
 The HFM directly extracts features from each level of the hierarchy separately and jointly.
@@
 Such flexibility can maintain its performance, while the corresponding number of parameters is also enlarged dramatically for a larger dimension of $M$.
-This paper alternatively proposes a sequential feature extraction method with shared loadings within each level of the hierarchy.
-Specifically, we define the operator $\mathcal{M}(\mbox{$\mathscr{X}$}_{t};\{\boldsymbol{G}_{m}\}_{m=1}^{2})=\boldsymbol{G}_{2}^{\top}(\boldsymbol{I}_{p_{2}}\otimes\boldsymbol{G}_{1}^{\top})\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})$, where $\boldsymbol{G}_{1}\in\mathbb{O}^{p_{1}\times r_{1}}$ and $\boldsymbol{G}_{2}\in\mathbb{O}^{p_{2}r_{1}\times r_{2}}$ are the component matrices for summarizing predictors, where $r_{1}$ and $r_{2}$ are the interim dimensions of features extracted from the corresponding mode.
+However, many recent hierarchical data may have a structure similar to the factorial data, and the variables at each level are also arranged in a similar way.
+This motivates us to alternatively proposes a sequential feature extraction method with shared loadings within each level of the hierarchy.
+Specifically, we define the operator $\mathcal{M}(\mbox{$\mathscr{X}$}_{t};\{\boldsymbol{G}_{m}\}_{m=1}^{2})=\boldsymbol{G}_{2}^{\top}(\boldsymbol{I}_{p_{2}}\otimes\boldsymbol{G}_{1}^{\top})\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})$, where $\boldsymbol{G}_{1}\in\mathbb{O}^{p_{1}\times r_{1}}$ and $\boldsymbol{G}_{2}\in\mathbb{O}^{p_{2}r_{1}\times r_{2}}$ are the component matrices for summarizing predictors, and $r_{1}$ and $r_{2}$ are the interim dimensions of features extracted from the corresponding mode.
 To interpret this construction, we liken the hierarchy of $\mbox{$\mathscr{X}$}_{t}\in\mathbb{R}^{p_{1}\times p_{2}}$ to a tree structure with $p_{2}$ branches, each branch containing $p_{1}$ leaves.
@@
 Shortly speaking, the HOFM concentrates more on branch-specific features, while our method focuses more on common features.
-In real applications, some weak branch-specific features may be ignored by the introduced method at the first layer of feature extraction, but they may also eventually be averaged out by the HOFM at later layers, i.e., our method does not miss globally important features.
+When the data have an irregular structure or one artificially switches the orders of variables at one level, our method may lead to a misleading interpretation, and we may then have to consider the HOFM or even HFM for the feature extracting.
+As a result, a huge number of parameters will be involved, making it impossible for high-dimensional data.
 For a general dimension of $M$, we assume that each mode of $\mbox{$\mathscr{X}$}_{t}\in\mathbb{R}^{p_{1}\times p_{2}\times\cdots\times p_{M}}$ is nested within its subsequent mode, and the feature extraction can then be iterated until the last mode of $\mbox{$\mathscr{X}$}_{t}$.
@@
 Remark 2.
-Consider a special case with $r=s$ and $\boldsymbol{\Theta}$ being diagonal, and then the above coefficient tensor $\mathscr{A}$ will admit a form of tensor train decomposition (oseledets2011tensor; zhou2022optimal; si2024efficient) if predictors are further arranged in a reverse hierarchical order.
-See Supplementary Material for the detailed illustration.
+Consider a special case with $r=s$ and $\boldsymbol{\Theta}$ being diagonal, and then the above coefficient tensor $\mathscr{A}$ will admit a form of tensor train decomposition (Oseledets, 2011; Zhou et al., 2022; Si et al., 2024) if predictors are further arranged in a reverse hierarchical order.
+See Section B.6 of the Supplementary Material for the detailed illustration.
 2.3 General hierarchical structure
-This subsection discusses the case with partially hierarchical predictors and responses, and some of their modes may not be strictly nested within the others, possibly with unknown or even multiple hierarchical relationships among the modes.
+This subsection discusses a more practical case with partially hierarchical predictors and responses, and some of their modes may not be strictly nested within the others, possibly with unknown or even multiple hierarchical relationships among the modes.
 We first consider the feature extraction on predictors.
@@
 For example, we consider the case with $M=3$, and mode 1 is nested within mode 2.
-The pseudo-structure can then be with an action order of $\{1,2,3\}$ or $\{3,1,2\}$.
+The pseudo-structure can then be represented by action order $(1,2,3)$ or $(3,1,2)$.
 In general, for the predictor tensor $\mbox{$\mathscr{X}$}_{t}\in\mathbb{R}^{p_{1}\times p_{2}\times\cdots\times p_{M}}$, an action order can be denoted by $\alpha=(\alpha_{(1)},\alpha_{(2)},\ldots,\alpha_{(M)})$, which is a permutation of set $\{1,2,\ldots,M\}$.
@@
 It then holds that, for a different action order $\alpha^{\prime}=(\alpha_{(1)}^{\prime},\alpha_{(2)}^{\prime},\ldots,\alpha_{(M)}^{\prime})$, there exist $\{\boldsymbol{G}_{m}^{\prime}\in\mathbb{O}^{r^{\prime}_{m-1}p_{\alpha^{\prime}_{(m)}}\times r^{\prime}_{m}}\}_{m=1}^{M}$ with $(r_{0}^{\prime},r_{M}^{\prime})=(1,r)$ such that $\mathcal{M}(\mathcal{T}(\mbox{$\mathscr{X}$};\alpha);\{\boldsymbol{G}_{m}\}_{m=1}^{M})=\mathcal{M}(\mathcal{T}(\mbox{$\mathscr{X}$};\alpha^{\prime});\{\boldsymbol{G}_{m}^{\prime}\}_{m=1}^{M})$.
-The above proposition proves a nice equivalence of all elements inside $\mathcal{P}_{x}(M)$, however, it is a theoretical result only.
+The above proposition proves a nice equivalence of all elements inside $\mathcal{P}_{x}(M)$; however, it is a theoretical result only.
 In real applications, each action order may have a totally different number of parameters, and a wrongly selected order may end up with a huge number of parameters.
@@
 The philosophy of the above feature extraction is analogous to that of CP decomposition, which concatenates multiple separately extracted features.
-Note that, comparing with Tucker decomposition, CP decomposition requests much less parameters by sacrificing orthogonality of loading matrices.
+Note that, comparing with Tucker decomposition, CP decomposition requires much less parameters by sacrificing orthogonality of loading matrices.
 The feature extraction from responses $\mbox{$\mathscr{Y}$}_{t}\in\mathbb{R}^{q_{1}\times q_{2}\times\cdots\times q_{N}}$ can be conducted in a similar way.
@@
 For autoregressive models at (2.12) and (2.13), the coefficient tensor can be denoted by $\mbox{$\mathscr{A}$}_{\mathrm{AR}}\in\mathbb{R}^{q_{1}\times\cdots\times q_{N}\times q_{1}\times\cdots\times q_{N}\times L}$ with $\mbox{$\mathscr{A}$}_{\mathrm{AR}}]_{N}=([\mbox{$\mathscr{A}$}_{1}]_{N},[\mbox{$\mathscr{A}$}_{2}]_{N},\ldots,[\mbox{$\mathscr{A}$}_{L}]_{N})$ being the mode-$N$ sequential matricization.
-Let $\boldsymbol{\Theta}_{\mathrm{AR}}=[\boldsymbol{\Theta}_{1},\ldots,\boldsymbol{\Theta}_{L}]\in\mathbb{R}^{s\times rL}$, and it holds that $\mbox{$\mathscr{A}$}_{\mathrm{AR}}]_{N}=\boldsymbol{\Lambda}_{y}(\boldsymbol{\Lambda}_{y}^{\top}\boldsymbol{\Lambda}_{y})^{-1}\boldsymbol{\Theta}_{\mathrm{AR}}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top})$.
+Let $\boldsymbol{\Theta}_{\mathrm{AR}}=(\boldsymbol{\Theta}_{1},\ldots,\boldsymbol{\Theta}_{L})\in\mathbb{R}^{s\times rL}$, and it holds that $\mbox{$\mathscr{A}$}_{\mathrm{AR}}]_{N}=\boldsymbol{\Lambda}_{y}(\boldsymbol{\Lambda}_{y}^{\top}\boldsymbol{\Lambda}_{y})^{-1}\boldsymbol{\Theta}_{\mathrm{AR}}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top})$.
 Moreover, by reparameterizing $(\boldsymbol{\Lambda}_{y}^{\top}\boldsymbol{\Lambda}_{y})^{-1}\boldsymbol{\Theta}_{\mathrm{AR}}$ into $\boldsymbol{\Theta}_{\mathrm{AR}}$, we have $\mbox{$\mathscr{A}$}_{\mathrm{AR}}]_{N}=\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}_{\mathrm{AR}}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top})$, and then the parameter space can be defined as
@@
 For an observed time series $\{\mbox{$\mathscr{Y}$}_{1-L},\ldots,\mbox{$\mathscr{Y}$}_{0},\mbox{$\mathscr{Y}$}_{1},\ldots,\mbox{$\mathscr{Y}$}_{T}\}$ generated by the model at (2.12) and (2.13), the predictor tensor $\mbox{$\mathscr{X}$}_{t}\in\mathbb{R}^{q_{1}\times\cdots\times q_{N}\times L}$ can be formed by stacking $\mbox{$\mathscr{Y}$}_{t-1},\dots,\mbox{$\mathscr{Y}$}_{t-L}$ together, and the OLS estimator is hence defined as
-$\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{AR}}=\operatorname*{arg\,min}_{\mbox{$\mathscr{A}$}\in\mathcal{S}_{\mathrm{AR}}}\mathcal{L}_{T}(\mbox{$\mathscr{A}$})\quad\text{with}\quad\mathcal{L}_{T}(\mbox{$\mathscr{A}$})=\frac{1}{T}\sum_{t=1}^{T}\left\lVert\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})-[\mbox{$\mathscr{A}$}]_{N}\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})\right\rVert_{2}^{2}.$
+$\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{AR}}=\operatorname*{arg\,min}_{\mbox{$\mathscr{A}$}\in\mathcal{S}_{\mathrm{AR}}}\mathcal{L}_{T}(\mbox{$\mathscr{A}$})\quad\text{with}\quad\mathcal{L}_{T}(\mbox{$\mathscr{A}$})=\frac{1}{T}\sum_{t=1}^{T}\left\lVert\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})-[\mbox{$\mathscr{A}$}]_{N}\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})\right\rVert_{2}^{2}.$ (3.1)
 Moreover, for the tensor regression model at (2.1) and (2.11), by letting $L=1$, we can similarly define its parameter space and OLS estimator, denoted by $\mathcal{S}_{\mathrm{LR}}$ and $\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{LR}}$, respectively.
@@
 Such projection has no error, while the number of parameters will be enlarged.
-Without loss of generality, we will assume $\mathcal{P}_{x}(M,K_{x})$ and $\mathcal{P}_{y}(N,K_{y})$ to be the best sets in establishing theoretical properties, and Section 4 will discuss the selection of these hyperparameters.
-3.2 Non-asymptotic properties for tensor regression
+Without loss of generality, we will assume $\mathcal{P}_{x}(M,K_{x})$ and $\mathcal{P}_{y}(N,K_{y})$ to be known in establishing theoretical properties, and Section 4 will discuss the selection of these hyperparameters.
+3.2 Non-asymptotic properties for tensor regression settings
 […]
-We begin with the tensor-on-tensor regression setting, where $\{(\mbox{$\mathscr{Y}$}_{t},\mbox{$\mathscr{X}$}_{t})\}_{t=1}^{T}$ are assumed to be independent and identically distributed (i.i.d.) observations.
-The following assumptions are imposed.
 Assumption 2.
-Predictors $\{\boldsymbol{x}_{t}\}$ are i.i.d.
-with $\mathbb{E}(\boldsymbol{x}_{t})=\boldsymbol{0}$, $\mathrm{var}(\boldsymbol{x}_{t})=\boldsymbol{\Sigma}_{x}$, and follow a $\sigma^{2}$-sub-Gaussian distribution.
+Predictor $\boldsymbol{x}_{t}$ follows a $\sigma^{2}$-sub-Gaussian distribution with $\mathbb{E}(\boldsymbol{x}_{t})=\boldsymbol{0}$ and $\mathrm{var}(\boldsymbol{x}_{t})=\boldsymbol{\Sigma}_{x}$.
 There exist constants $0<c_{x}<C_{x}<\infty$ such that $c_{x}\leq\lambda_{\min}(\boldsymbol{\Sigma}_{x})\leq\lambda_{\max}(\boldsymbol{\Sigma}_{x})\leq C_{x}$, where the two quantities $c_{x}$ and $C_{x}$ depend on the dimensions of $p_{j}$’s, and they may shrink to zero or diverge to infinity as the dimensions increase.
 Assumption 3.
-Error terms $\{\boldsymbol{e}_{t}\}$ are i.i.d. and, conditional on $\boldsymbol{x}_{t}$, $\boldsymbol{e}_{t}$ has mean zero and follows $\kappa^{2}$-sub-Gaussian distribution.
+Error term $\boldsymbol{e}_{t}$, conditional on predictor $\boldsymbol{x}_{t}$, follows $\kappa^{2}$-sub-Gaussian distribution with mean zero.
 Assumption 4.
-The matrix $\boldsymbol{\Theta}_{\mathrm{LR}}$ has its operator norm upper bounded by a constant $g_{1}>0$.
-Assumption 2 guarantees the well-conditioning of the covariance matrix of the predictors.
-Assumption 3, on the other hand, controls the tail behavior of the noise by assuming sub-Gaussianity, a standard requirement in high-dimensional statistics that facilitates concentration inequalities.
-Both assumptions are widely adopted in the literature on high-dimensional regression and low-rank estimation; see, for example, wainwright2019high; si2024efficient; cai2025efficient.
-Assumption 4 imposes an upper bound on the operator norm of the coefficient matrix $\boldsymbol{\Theta}_{\mathrm{LR}}$.
-This is a mild technical condition, as the dimensions $s_{\mathrm{tot}}$ and $r_{\mathrm{tot}}$ of $\boldsymbol{\Theta}_{\mathrm{LR}}$ are typically small by carefully choosing the action orders.
-The following theorem provides a non-asymptotic upper bound on the estimation error of estimator $\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{LR}}$.
+It holds that $\|\boldsymbol{\Theta}_{\mathrm{LR}}\|_{\mathrm{op}}<g_{1}$, where $\|\cdot\|_{\mathrm{op}}$ is the operator norm.
+Assumption 2 guarantees the well-conditioning of predictors, while Assumption 3 controls the tail behavior of noises by the sub-Gaussianity.
+The two assumptions are standard high-dimensional settings, and they have been widely used in the literature (Wainwright, 2019; Si et al., 2024; Cai et al., 2025).
+Assumption 4 imposes an upper bound on the operator norm of the coefficient matrix $\boldsymbol{\Theta}_{\mathrm{LR}}\in\mathbb{R}^{s\times r}$.
+It is used to handle the non-orthogonality of loading matrices $\boldsymbol{\Lambda}_{x}$ and $\boldsymbol{\Lambda}_{y}$, and we can remove it with the cost of a large model complexity.
+This assumption is mild since both dimensions $s$ and $r$ are typically small by carefully choosing the action orders.
+Denote the model complexity by $d_{\mathrm{LR}}=\sum_{k=1}^{K_{x}}d_{x}(\alpha_{k})+\sum_{k=1}^{K_{y}}d_{y}(\beta_{k})+sr$, and the estimation error bound is stated below.
 Theorem 1.
 Suppose that Assumptions 2 - 4 hold, and the true coefficient tensor $\mbox{$\mathscr{A}$}_{\mathrm{LR}}^{*}\in\mathcal{S}_{\mathrm{LR}}$.
-If the sample size $T\gtrsim(\sigma^{4}/c_{x}^{2})d_{\mathrm{LR}}\log((\sigma^{4}/c_{x}^{2})d_{\mathrm{LR}})$, then there exists $c_{1}>0$ such that
+If the sample size $T\gtrsim(\sigma^{4}/c_{x}^{2})d_{\mathrm{LR}}\log\{(\sigma^{4}/c_{x}^{2})d_{\mathrm{LR}}\}$, then,
 $\|\mbox{$\mathscr{\widehat{A}}$}_{\mathrm{LR}}-\mbox{$\mathscr{A}$}^{*}_{\mathrm{LR}}\|_{\mathrm{F}}\lesssim\frac{\kappa\sigma}{c_{x}}\sqrt{\frac{d_{\mathrm{LR}}\log T}{T}}$
-with probability at least $1-2\exp\left\{-c_{1}d_{\mathrm{LR}}\log T\right\}$, where $d_{\mathrm{LR}}=\sum_{k=1}^{K_{x}}d_{x}(\alpha_{k})+\sum_{k=1}^{K_{y}}d_{y}(\beta_{k})+sr$.
-We define $d_{\mathrm{LR}}$ as the model complexity of the low-rank regression estimator, which can be rewritten explicitly as $d_{\mathrm{LR}}=K_{x}r_{\max}^{2}\sum_{m=1}^{M}p_{m}+K_{y}s_{\max}^{2}\sum_{n=1}^{N}q_{n}+K_{x}K_{y}r_{\max}s_{\max}$, where $r_{\max}\vcentcolon=\max_{m,k}r_{m,k}$ and $s_{\max}\vcentcolon=\max_{n,k^{\prime}}s_{n,k^{\prime}}$.
-From Theorem 1, our proposed low-rank regression estimator $\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{LR}}$, up to a logarithmic factor, achieves an estimation error bound of order $O_{p}(\sqrt{d_{\mathrm{LR}}/T})$ when the sample size $T\gtrsim d_{\mathrm{LR}}/\text{SNR}^{2}$, where $\text{SNR}=c_{x}/\sigma^{2}$ is the predictor signal-to-noise ratio.
-This rate aligns with the general optimal rate for low-rank regression established in han2022optimal.
+with probability at least $1-2\exp\left\{-C_{1}d_{\mathrm{LR}}\log T\right\}$, where $C_{1}$ is a positive constant.
+Let $r_{\max}\vcentcolon=\max_{m,k}r_{m,k}$ and $s_{\max}\vcentcolon=\max_{n,k^{\prime}}s_{n,k^{\prime}}$, and the model complexity can be simplified as $d_{\mathrm{LR}}\leq K_{x}r_{\max}^{2}\sum_{m=1}^{M}p_{m}+K_{y}s_{\max}^{2}\sum_{n=1}^{N}q_{n}+K_{x}K_{y}r_{\max}s_{\max}$.
+From the above theorem, the OLS estimator $\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{LR}}$ achieves an estimation error bound of order $O_{p}(\sqrt{d_{\mathrm{LR}}/T})$, up to a logarithmic factor, when the sample size $T\gtrsim d_{\mathrm{LR}}/(c_{x}/\sigma^{2})^{2}$.
+This rate aligns with the general optimal rate for low-rank regression established in Han et al.
+(2022).
 The model complexity $d_{\mathrm{LR}}$ grows linearly with the dimensions of each mode $\{p_{m}\}$ and $\{q_{n}\}$, and quadratically with the interim ranks $\{r_{m,k}\}$ and $\{s_{n,k^{\prime}}\}$.
-The logarithmic factor $\log T$, negligible in most common cases, arises from the non-orthogonality of the loadings.
-In contrast, if we assume the coefficient tensor $\mathscr{A}$ follows a Tucker decomposition with Tucker ranks $(s_{1},s_{2},\ldots,s_{N},r_{1},r_{2},\ldots,r_{M})$, the corresponding complexity would be $d_{\mathrm{Tucker}}=\prod_{n=1}^{N}s_{n}\times\prod_{m=1}^{M}r_{m}+\sum_{n=1}^{N}q_{n}s_{n}+\sum_{m=1}^{M}p_{m}r_{m}$, which grows exponentially with the tensor orders $N$ and $M$ (han2022optimal; wang2022high).
-3.3 Non-asymptotic properties for tensor autoregression
-We now turn to the tensor autoregressive model specified in (2.12). This setting introduces temporal dependence, and the assumptions are adapted accordingly.
+3.3 Non-asymptotic properties for tensor autoregressive models
+This subsection states non-asymptotic properties of the OLS estimator $\mbox{$\mathscr{\widehat{A}}$}_{\mathrm{AR}}$ for tensor autoregressive models at (2.12) and (2.13), and we first introduce some conditions below.
 Assumption 5.
-Let error terms $\boldsymbol{e}_{t}=\boldsymbol{\Sigma}_{e}^{1/2}\boldsymbol{\xi}_{t}$.
-Random vectors $\{\boldsymbol{\xi}_{t}\}$ are i.i.d.
-with mean zero and covariance $\boldsymbol{I}_{Q}$, the entries $(\boldsymbol{\xi}_{tj})_{1\leq j\leq Q}$ of $\boldsymbol{\xi}_{t}$ are mutually independent and $\kappa^{2}$-sub-Gaussian distributed, and there exist constants $0<c_{e}<C_{e}<\infty$ such that $c_{e}\leq\lambda_{\min}(\boldsymbol{\Sigma}_{e})\leq\lambda_{\max}(\boldsymbol{\Sigma}_{e})\leq C_{e}$.
+Let $\boldsymbol{e}_{t}=\boldsymbol{\Sigma}_{e}^{1/2}\boldsymbol{\xi}_{t}$ with $\boldsymbol{\Sigma}_{e}=\mathrm{var}(\boldsymbol{e}_{t})$.
+Random vectors $\{\boldsymbol{\xi}_{t}\}$ are $i.i.d.$ with mean zero and covariance matrix $\boldsymbol{I}_{Q}$, and the entries of $\boldsymbol{\xi}_{t}\in\mathbb{R}^{Q}$ are independent and $\kappa^{2}$-sub-Gaussian distributed.
+Moreover, there exist constants $0<c_{e}<C_{e}<\infty$ such that $c_{e}\leq\lambda_{\min}(\boldsymbol{\Sigma}_{e})\leq\lambda_{\max}(\boldsymbol{\Sigma}_{e})\leq C_{e}$, where the two quantities depend on the dimensions of $q_{j}$’s, and they may shrink to zero or diverge to infinity as the dimensions increase.
 Assumption 6.
-The matrix $\boldsymbol{\Theta}_{\mathrm{AR}}$ has its operator norm upper bounded by a constant $g_{2}>0$.
-The sub-Gaussianity assumption in Assumption 5 is commonly used in the analysis of high-dimensional time series models.
-It provides control over the tail behavior of the error terms and facilitates the use of concentration inequalities.
-This assumption is widely adopted in recent literature, including zheng2020finite; wang2024high; huang2025supervised.
-Assumption 6 is essential for the consistency of the estimator in the autoregressive setting, which is relatively mild and practically reasonable.
-In our setting, $\boldsymbol{\Lambda}_{y}$ and $\boldsymbol{\Lambda}_{x}$ may exhibit multi-collinearity, and this assumption is weaker than enforcing a lower bound on the smallest singular value of $\boldsymbol{\Lambda}_{x}$ and $\boldsymbol{\Lambda}_{y}$ for model well-behavedness.
-Following basu2015regularized, we define the following quantities,
+It holds that $\|\boldsymbol{\Theta}_{\mathrm{AR}}\|_{\mathrm{op}}<g_{2}$, where $\|\cdot\|_{\mathrm{op}}$ is the operator norm.
+The sub-Gaussian assumption has also been commonly used in the literature of high-dimensional time series (Zheng and Cheng, 2020; Wang et al., 2024; Huang et al., 2025).
+We next establish non-asymptotic properties of the OLS estimator $\mbox{$\mathscr{\widehat{A}}$}_{\mathrm{AR}}$, and they rely on the temporal and cross-sectional dependence of $\{\mbox{$\mathscr{Y}$}_{t}\}$ (Basu and Michailidis, 2015).
+To this end, for the matrix polynomial of model (2.12), we define the quantities below,
 $\mu_{\mathrm{min}}(\mbox{$\mathscr{A}$})=\min_{|z|=1}\lambda_{\mathrm{min}}(\bar{\mbox{$\mathscr{A}$}}(z)\mbox{$\mathscr{A}$}(z))\quad\text{and}\quad\mu_{\mathrm{max}}(\mbox{$\mathscr{A}$})=\max_{|z|=1}\lambda_{\mathrm{max}}(\bar{\mbox{$\mathscr{A}$}}(z)\mbox{$\mathscr{A}$}(z)),$
-where $\bar{\mbox{$\mathscr{A}$}}(z)$ denotes the conjugate transpose of $\mbox{$\mathscr{A}$}(z)$.
-Furthermore, the AR($L$) model in (2.12) can be alternatively represented as a $QL$-dimensional AR(1) model with its transition matrix $\boldsymbol{B}$ defined in (C.3) in the Supplementary Material.
-We similarly define the matrix polynomial associated with the lifted AR(1) representation as $\boldsymbol{B}(z)=\boldsymbol{I}_{QL}-\boldsymbol{B}z$.
-We then define
-$\mu_{\mathrm{min}}(\boldsymbol{B})=\min_{|z|=1}\lambda_{\mathrm{min}}(\bar{\boldsymbol{B}}(z)\boldsymbol{B}(z))\quad\text{and}\quad\mu_{\mathrm{max}}(\boldsymbol{B})=\max_{|z|=1}\lambda_{\mathrm{max}}(\bar{\boldsymbol{B}}(z)\boldsymbol{B}(z)),$
-which quantify the smallest and largest eigenvalues the Hermitian matrix $\bar{\boldsymbol{B}}(z)\boldsymbol{B}(z)$ can take on the unit circle.
-Note that $\mu_{\mathrm{min}}(\boldsymbol{B})$, $\mu_{\mathrm{max}}(\boldsymbol{B})$ are not necessarily the same as $\mu_{\mathrm{min}}(\mbox{$\mathscr{A}$})$, $\mu_{\mathrm{max}}(\mbox{$\mathscr{A}$})$ (basu2015regularized), and we can show $\mu_{\mathrm{min}}(\boldsymbol{B}^{\ast})\leq\mu_{\mathrm{max}}(\mbox{$\mathscr{A}$}^{\ast})$.
-We then state the error upper bound for the estimator $\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{AR}}$ below.
+[…]
 Theorem 2.
-Suppose that Assumptions 1, 5, and 6 hold, and the true coefficient tensor $\mbox{$\mathscr{A}$}_{\mathrm{AR}}^{\ast}\in\mathcal{S}_{\mathrm{AR}}$.
-If the sample size $T\gtrsim(\kappa^{4}\kappa_{U,B}^{2}/\kappa_{L,A}^{2})d_{\mathrm{AR}}\log((\kappa^{4}\kappa_{U,B}^{2}/\kappa_{L,A}^{2})d_{\mathrm{AR}})$, then
+Suppose that Assumptions 1, 5 and 6 hold, and the true coefficient tensor $\mbox{$\mathscr{A}$}_{\mathrm{AR}}^{\ast}\in\mathcal{S}_{\mathrm{AR}}$.
+If the sample size $T\gtrsim(\kappa^{4}\kappa_{U,B}^{2}/\kappa_{L,A}^{2})d_{\mathrm{AR}}\log\{\kappa^{4}\kappa_{U,B}^{2}/\kappa_{L,A}^{2})d_{\mathrm{AR}}\}$, then
 $\|{\mbox{$\mathscr{\widehat{A}}$}_{\mathrm{AR}}-\mbox{$\mathscr{A}$}_{\mathrm{AR}}^{\ast}}\|_{\mathrm{F}}\lesssim\frac{\kappa^{2}\kappa_{U,B}^{\prime}}{\kappa_{L,A}}\sqrt{\frac{d_{\mathrm{AR}}\log T}{T}}$
-with probability at least $1-2\exp\left\{-c_{1}d_{\mathrm{AR}}\log T\right\}$, where $d_{\mathrm{AR}}=\sum_{k=1}^{K_{x}}d_{x}(\alpha_{k})+\sum_{k=1}^{K_{y}}d_{y}(\beta_{k})+Lsr$, $\boldsymbol{B}^{\ast}$ is the true value of $\boldsymbol{B}$ constructed from $\mbox{$\mathscr{A}$}^{\ast}$, $\kappa_{U,B}=C_{e}/\mu_{\mathrm{min}}(\boldsymbol{B}^{\ast})$, $\kappa_{L,A}=c_{e}/\mu_{\mathrm{max}}(\mbox{$\mathscr{A}$}^{\ast})$, $\kappa^{\prime}_{U,B}=C_{e}/\mu^{1/2}_{\mathrm{min}}(\boldsymbol{B}^{\ast})$, and $c_{1}>0$ is a constant.
-We define $d_{\mathrm{AR}}$ as the model complexity of the low-rank autoregressive estimator, which can be explicitly expressed as $d_{\mathrm{AR}}=(K_{x}r_{\max}^{2}+K_{y}s_{\max}^{2})\sum_{n=1}^{N}q_{n}+LK_{x}K_{y}r_{\max}s_{\max}$.
-When the order $N$, the number of action orders $K_{x}$, $K_{y}$, and the largest ranks $s_{\max}$ and $r_{\max}$ are held fixed, the key parameters $c_{e}$, $C_{e}$, $\kappa^{2}$, $\mu_{\mathrm{min}}(\boldsymbol{B}^{\ast})$, and $\mu_{\mathrm{max}}(\mbox{$\mathscr{A}$}^{\ast})$ are all bounded away from zero and infinity.
-Consequently, the estimation error of the low-rank tensor autoregressive estimator satisfies $\|{\mbox{$\mathscr{\widehat{A}}$}_{\mathrm{AR}}-\mbox{$\mathscr{A}$}_{\mathrm{AR}}^{\ast}}\|_{\mathrm{F}}=O_{p}(\sqrt{d_{\mathrm{AR}}/T})$ up to a logarithmic term, where $d_{\mathrm{AR}}$ is the effective model complexity of the autoregressive estimator.
+with probability at least $1-2\exp\left\{-C_{1}d_{\mathrm{AR}}\log T\right\}$, where $C_{1}>0$ is a positive constant.
+[…]
 4 Algorithms and hyperparameter selection
 4.1 Algorithms
-This subsection introduces an algorithm for computing the low-rank estimator defined in (3.1) under the general setting.
-To preserve the sequential low-rank structure of $\mbox{$\mathscr{A}$}]_{N}$, we adopt an alternating least squares (ALS) approach, which iteratively updates each component $\boldsymbol{H}_{n,k^{\prime}}$ and $\boldsymbol{G}_{m,k}$ while keeping all other components fixed.
-We introduce some shorthand notations for clarity.
-Let $\boldsymbol{y}_{t}=\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})$ and $\boldsymbol{x}_{t}=\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})=(\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t-1})^{\top},\dots,\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t-L})^{\top})^{\top}$, and our model can be rewritten in the vectorized form as:
-$\boldsymbol{y}_{t}=\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top})\boldsymbol{x}_{t}+\boldsymbol{e}_{t},$ (4.1)
+This subsection introduces an alternating least squares (ALS) method to search for the estimate $\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{AR}}$ in Section 3.1.
+Specifically, for each $\mbox{$\mathscr{A}$}\in\mathcal{S}_{\mathrm{AR}}$ at (3.1), its mode-$N$ sequential matricization has the form of $\mbox{$\mathscr{A}$}]_{N}=\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top})$, where $\boldsymbol{\Theta}=(\boldsymbol{\Theta}_{1},\ldots,\boldsymbol{\Theta}_{L})$.
+Let
+$\boldsymbol{\Lambda}_{x,k}$ $=\boldsymbol{T}_{x}^{\top}(\alpha_{k})(\boldsymbol{I}_{p_{\alpha^{k}_{(2)}}\cdots p_{\alpha^{k}_{(M)}}}\otimes\boldsymbol{G}_{1,k})\cdots(\boldsymbol{I}_{p_{\alpha^{k}_{(M)}}}\otimes\boldsymbol{G}_{M-1,k})\boldsymbol{G}_{M,k}\in\mathbb{R}^{P\times r_{M,k}}\hskip 8.53581pt\text{and}$
+$\boldsymbol{\Lambda}_{y,k^{\prime}}$ $=\boldsymbol{T}_{y}^{\top}(\beta_{k^{\prime}})(\boldsymbol{I}_{q_{\beta^{k^{\prime}}_{(2)}}\cdots q_{\beta^{k^{\prime}}_{(N)}}}\otimes\boldsymbol{H}_{1,k^{\prime}})\cdots(\boldsymbol{I}_{q_{\beta^{k^{\prime}}_{(N)}}}\otimes\boldsymbol{H}_{N-1,k^{\prime}})\boldsymbol{H}_{N,k^{\prime}}\in\mathbb{R}^{Q\times s_{N,k^{\prime}}}$
+for $1\leq k\leq K_{x}$ and $1\leq k^{\prime}\leq K_{y}$, and it then holds that $\boldsymbol{\Lambda}_{x}=(\boldsymbol{\Lambda}_{x,1},\ldots,\boldsymbol{\Lambda}_{x,K_{x}})\in\mathbb{R}^{P\times r}$ and $\boldsymbol{\Lambda}_{y}=(\boldsymbol{\Lambda}_{y,1},\ldots,\boldsymbol{\Lambda}_{y,K_{y}})\in\mathbb{R}^{Q\times s}$; see (2.10) for details.
+Moreover, for each $1\leq l\leq L$, $\mbox{$\mathscr{A}$}_{l}]_{N}=\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}_{l}\boldsymbol{\Lambda}_{x}^{\top}=\sum_{k^{\prime}=1}^{K_{y}}\sum_{k=1}^{K_{x}}\boldsymbol{\Lambda}_{y,k^{\prime}}\boldsymbol{\Theta}_{k^{\prime},k,l}\boldsymbol{\Lambda}_{x,k}^{\top}$, where $\boldsymbol{\Theta}_{k^{\prime},k,l}\in\mathbb{R}^{s_{N,k^{\prime}}\times r_{M,k}}$ with $1\leq k\leq K_{x}$ and $1\leq k^{\prime}\leq K_{y}$ are the corresponding submatrices of $\boldsymbol{\Theta}_{l}$.
+As a result, it is equivalent to minimizing the loss function of $\mathcal{L}_{T}(\mbox{$\mathscr{A}$})$ at (3.1) with respect to two blocks of parameters: $\{\boldsymbol{\Lambda}_{x,k}(\boldsymbol{G}_{m,k},1\leq m\leq M),1\leq k\leq K_{x}\}$, $\{\boldsymbol{\Lambda}_{y,k}(\boldsymbol{H}_{n,k},1\leq n\leq N),1\leq k^{\prime}\leq K_{y}\}$, and $\boldsymbol{\Theta}$, and the first two blocks of parameters correspond to $\boldsymbol{\Lambda}_{x}$ and $\boldsymbol{\Lambda}_{y}$, respectively.
+Algorithm 1 ALS for tensor autoregression
+Input: Time series $\{\boldsymbol{y}_{t},-L+1\leq t\leq T\}$, initial coefficients $\mbox{$\mathscr{A}$}^{(0)}]_{N}=\boldsymbol{\Lambda}_{y}^{(0)}\boldsymbol{\Theta}^{(0)}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{(0)\top})$ with component matrices $\{\{\boldsymbol{H}^{(0)}_{n,k}\}_{n=1}^{N}\}_{k=1}^{K_{y}}$ and $\{\{\boldsymbol{G}^{(0)}_{m,k}\}_{m=1}^{M}\}_{k=1}^{K_{x}}$.
+repeat
+  for $k=1,2,\ldots,K_{x}$ do
+   Sequentially update $\boldsymbol{G}_{m,k}$ for $1\leq m\leq M$ while keeping all other components being fixed at the values from the last iteration.
+   Standardize each $\boldsymbol{G}_{m,k}$ for $1\leq m\leq M$ such that each column has unit norm.
+  end for
+  for $k=1,2,\ldots,K_{y}$ do
+   Sequentially update $\boldsymbol{H}_{n,k}$ for $1\leq n\leq N$ while keeping all other components being fixed at the values from the last iteration.
+   Standardize each $\boldsymbol{H}_{n,k}$ for $1\leq n\leq N$ such that each column has unit norm.
+  end for
+  Update $\boldsymbol{\Theta}=(\boldsymbol{\Theta}_{1},\ldots,\boldsymbol{\Theta}_{L})$.
+until convergence
+ Call the SSVD algorithm to orthogonalize all $\boldsymbol{G}_{m,k}$’s and $\boldsymbol{H}_{n,k}$’s, and then update $\boldsymbol{\Theta}$.
+Return $\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{AR}}$ such that $\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{AR}}]_{N}=\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top})$.
+Let $\boldsymbol{y}_{t}=\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})$, and the tensor autoregressive model at (2.12) and (2.13) has the form of $\boldsymbol{y}_{t}=\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}_{1}\boldsymbol{\Lambda}_{x}^{\top}\boldsymbol{y}_{t-1}+\cdots+\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}_{L}\boldsymbol{\Lambda}_{x}^{\top}\boldsymbol{y}_{t-L}+\boldsymbol{e}_{t}$ or
+$\begin{split}\boldsymbol{y}_{t}=\sum_{k=1}^{K_{x}}\sum_{k^{\prime}=1}^{K_{y}}\boldsymbol{\Lambda}_{y,k^{\prime}}\boldsymbol{\Theta}_{k,k^{\prime},1}\boldsymbol{\Lambda}_{x,k}^{\top}\boldsymbol{y}_{t-1}+\cdots+\sum_{k=1}^{K_{x}}\sum_{k^{\prime}=1}^{K_{y}}\boldsymbol{\Lambda}_{y,k^{\prime}}\boldsymbol{\Theta}_{k,k^{\prime},L}\boldsymbol{\Lambda}_{x,k}^{\top}\boldsymbol{y}_{t-L}+\boldsymbol{e}_{t},\end{split}$ (4.1)
 […]
-We sequentially update each orthogonal matrix in each row block of $\boldsymbol{\Lambda}_{x}^{\top}$ (i.e., update $\boldsymbol{G}_{1,1}$ up to $\boldsymbol{G}_{M,1}$ for the first row block), and then switch to the next row block for another sequence of updates, and so on.
-Specifically, we convert the update of each component into a regression problem relegated to Supplementary Material, similar to coordinate descent.
-The update for $\boldsymbol{\Lambda}_{y}^{\top}$ follows a similar procedure.
-After updating all components in $\boldsymbol{\Lambda}_{x}^{\top}$ and $\boldsymbol{\Lambda}_{y}^{\top}$, we orthogonalize each row block of the two loading matrices and update the core tensor $\boldsymbol{\Theta}$ via an SSVD algorithm detailed in Supplementary Material.
-This process is repeated until convergence.
-For convenience, we arrange these matrices into $\boldsymbol{L}_{1},\boldsymbol{L}_{2},\ldots,\boldsymbol{L}_{K_{x}M+K_{y}N}$, where $\boldsymbol{L}_{i}$ is the $i$-th orthogonal matrix in the update procedure, and we re-define the loss function with regard to $\boldsymbol{L}_{i}$’s and $\boldsymbol{\Theta}$ as follows:
-$\mathcal{L}_{T}\big(\boldsymbol{L}_{1},\dots,\boldsymbol{L}_{K_{x}M+K_{y}N},\boldsymbol{\Theta}\big)=\frac{1}{T}\sum_{t=1}^{T}\|\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})-\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top})\operatorname*{vec}{(\mbox{$\mathscr{X}$}_{t})}\|^{2},$
-$\text{s.t., }\boldsymbol{G}_{m,k}=\boldsymbol{L}_{(k-1)M+m}\quad\text{and}\quad\boldsymbol{H}_{n,k^{\prime}}=\boldsymbol{L}_{K_{x}M+(k^{\prime}-1)N+n}$ (4.2)
-for all $m,k,n,k^{\prime}$, and $\boldsymbol{\Lambda}_{x}$ and $\boldsymbol{\Lambda}_{y}$ are defined in (2.10), respectively. The overall algorithm is summarized below:
-Algorithm 1 ALS for low-rank estimation
-Input: Data $\{(\boldsymbol{y}_{t},\boldsymbol{x}_{t})\}_{t=1}^{T}$, initial coefficient $\mbox{$\mathscr{A}$}^{(0)}]_{N}=\boldsymbol{\Lambda}_{y}^{(0)}\boldsymbol{\Theta}^{(0)}\boldsymbol{\Lambda}_{x}^{(0)\top}$, and components $\{\{\boldsymbol{H}^{(0)}_{n,k^{\prime}}\}_{n=1}^{N}\}_{k^{\prime}=1}^{K_{y}}$ and $\{\{\boldsymbol{G}^{(0)}_{m,k}\}_{m=1}^{M}\}_{k=1}^{K_{x}}$ arranged into $\{\boldsymbol{L}_{i}^{(0)}\}_{i=1}^{K_{x}M+K_{y}N}$.
-repeat
-  for $i=1,2,\dots,K_{x}M+K_{y}N$ do
-   $\boldsymbol{L}^{(u+0.5)}_{i}=\operatorname*{arg\,min}_{\boldsymbol{L}_{i}}\mathcal{L}_{T}\big(\boldsymbol{L}_{1}^{(u+0.5)},\dots,\boldsymbol{L}_{i-1}^{(u+0.5)},\boldsymbol{L}_{i},\boldsymbol{L}_{i+1}^{(u)},\dots,\boldsymbol{L}_{K_{x}M+K_{y}N}^{(u)},\boldsymbol{\Theta}^{(u)}\big)$
-  end for
-  $\{\boldsymbol{L}_{i}^{(u+1)}\}_{i=1}^{K_{x}M+K_{y}N},\boldsymbol{\Theta}^{(u)}=\text{SSVD}(\{\boldsymbol{L}_{i}^{(u+0.5)}\}_{i=1}^{K_{x}M+K_{y}N},\boldsymbol{\Theta}^{(u)})$
-  $\boldsymbol{\Theta}^{(u+1)}=\operatorname*{arg\,min}_{\boldsymbol{\Theta}}\mathcal{L}_{T}\big(\boldsymbol{L}_{1}^{(u+1)},\dots,\boldsymbol{L}_{K_{x}M+K_{y}N}^{(u+1)},\boldsymbol{\Theta}\big)$
-until convergence or $u=u_{\max}$, where $u_{\max}$ is the maximum iteration
-Return $\mbox{$\mathscr{A}$}^{(u)}]_{N}=\boldsymbol{\Lambda}_{y}^{(u)}\boldsymbol{\Theta}^{(u)}\boldsymbol{\Lambda}_{x}^{(u)\top}$ via (4.2), (2.10).
+In addition, when updating $\boldsymbol{\Lambda}_{x,k}(\boldsymbol{G}_{m,k},1\leq m\leq M)$ for each $1\leq k\leq K_{x}$, we sequentially update each of its component matrices $\boldsymbol{G}_{m,k}$ with $1\leq m\leq M$.
+Interestingly, the update of each $\boldsymbol{G}_{m,k}$ can be formulated as a least squares problem by fixing all other components, and it can be verified to have a solution with closed form; see Section B.2 of the supplementary file for details.
+The situation is the same when sequentially updating the component matrices of $\boldsymbol{\Lambda}_{y,k}(\boldsymbol{H}_{n,k},1\leq n\leq N)$ for each $1\leq k\leq K_{y}$, and hence all updating steps at Algorithm 1 have solutions with closed forms.
+Moreover, these component matrices are all orthogonal with unit norm, i.e., they are orthonormal.
+To guarantee the convergence, we give up the orthogonality during iterations, and these component matrices will be orthogonalized at the end by a SSVD algorithm detailed at Section B.4 of the supplementary file.
+For the initialization, we may simply consider random numbers around zero.
+In addition, the algorithm to search for the estimate $\widehat{\mbox{$\mathscr{A}$}}_{\mathrm{LR}}$ in Section 3.1 can be designed similarly.
 4.2 Hyperparameter selection
-Some hyperparameters involved in Algorithm 1 are typically unknown in most real-world applications, and their selection not only impacts the model’s performance, but also reveals the underlying structural information of the data.
-This subsection introduces a boosting-based procedure for determining the number of action orders $K_{y}$ and $K_{x}$, the corresponding action sets $\mathcal{P}_{y}(N,K_{y})$ and $\mathcal{P}_{x}(M,K_{x})$, and the associated rank sequences $\{\boldsymbol{s}_{k^{\prime}}\}_{k^{\prime}=1}^{K_{y}}$ and $\{\boldsymbol{r}_{k}\}_{k=1}^{K_{x}}$.
-To guide the selection of these hyperparameters, we adopt an information criterion.
-Let $\mbox{$\mathscr{\widehat{A}}$}(R)$ denote the estimator from (3.1), where the configuration $R$ collects all hyperparameters: $R=(K_{y},K_{x},\mathcal{P}_{y}(N,K_{y}),\mathcal{P}_{x}(M,K_{x}),\{\boldsymbol{s}_{k^{\prime}}\}_{k^{\prime}=1}^{K_{y}},\{\boldsymbol{r}_{k}\}_{k=1}^{K_{x}})$.
-The Bayesian information criterion (BIC) is then constructed below,
-$\mathrm{BIC}(R)=T\log\left\{\mathcal{L}_{T}(\mbox{$\mathscr{\widehat{A}}$}(R))\right\}+\phi d(R)\log(T),$ (4.3)
-where $\phi$ is a tuning parameter, and $d(R)=\sum_{k^{\prime}=1}^{K_{y}}\sum_{n=1}^{N}q_{n}s_{n-1,k^{\prime}}s_{n,k^{\prime}}+\sum_{k=1}^{K_{x}}\sum_{n=1}^{N}q_{n}r_{n-1,k}r_{n,k}+Ls_{\mathrm{tot}}r_{\mathrm{tot}}$ is the total number of parameters.
-The boosting-based hyperparameter selection procedure consists of two main stages: action order selection and rank reduction, which are detailed below.
-The overall procedure is summarized in Algorithm 2.
-Finally, the lag length $L$ is determined via a simple grid search to minimize the BIC over a predetermined candidate set $\{1,2,\ldots,L_{\max}\}$.
-4.2.1 Action order selection
-To construct the boosting-based algorithm, we first introduce the key notations and setup.
-Define $\widetilde{\boldsymbol{\Lambda}}_{x}(\alpha_{k},\boldsymbol{r}_{k})=\boldsymbol{\Lambda}_{x}(\{\alpha_{k}\})$, where the operator $\boldsymbol{\Lambda}_{x}(\cdot)$ is defined in (2.10) and $\boldsymbol{r}_{k}$ is the associated rank.
-It can be verified that for any two positive integer vectors $\boldsymbol{r}_{k,1}$ and $\boldsymbol{r}_{k,2}$, there exists a matrix $\widetilde{\boldsymbol{\Lambda}}_{x}(\alpha_{k},\boldsymbol{r}_{k,3})$ with $\boldsymbol{r}_{k,3}\leq\boldsymbol{r}_{k,1}+\boldsymbol{r}_{k,2}$ by elementwise comparison, satisfying
-$\widetilde{\boldsymbol{\Lambda}}_{x}(\alpha_{k},\boldsymbol{r}_{k,3})=\boldsymbol{O}\left[\widetilde{\boldsymbol{\Lambda}}_{x}(\alpha_{k},\boldsymbol{r}_{k,1}),\widetilde{\boldsymbol{\Lambda}}_{x}(\alpha_{k},\boldsymbol{r}_{k,2})\right],$ (4.4)
-where $\boldsymbol{O}$ is an invertible matrix that can be absorbed into $\boldsymbol{\Theta}$.
-The same applies to $\widetilde{\boldsymbol{\Lambda}}_{y}(\cdot)$ by defining $\widetilde{\boldsymbol{\Lambda}}_{y}(\beta_{k^{\prime}},\boldsymbol{s}_{k^{\prime}})=\boldsymbol{\Lambda}_{y}(\{\beta_{k^{\prime}}\})$, where $\boldsymbol{s}_{k^{\prime}}$ is the associated rank and $\boldsymbol{\Lambda}_{y}(\cdot)$ is defined similarly to (2.10).
-The proof is provided in Supplementary Material.
-We refer to any model defined by (4.1) with $\boldsymbol{\Lambda}_{y}=\widetilde{\boldsymbol{\Lambda}}_{y}(\beta_{k^{\prime}},\boldsymbol{1}_{N})$ and $\boldsymbol{\Lambda}_{x}=\widetilde{\boldsymbol{\Lambda}}_{x}(\alpha_{k},\boldsymbol{1}_{N})$ as a weak learner, denoted by $WL(\beta_{k^{\prime}},\alpha_{k})$, where $\boldsymbol{1}_{N}\in\mathbb{R}^{N}$ is a vector of ones.
-This weak learner serves as the base component in our boosting framework.
-The additive property introduced in (4.4) implies that the proposed model is additive with respect to rank increments: increasing the rank of a selected action order is equivalent to concatenating additional weak learners that share the same action order.
-To reduce computational complexity, we predefine candidate sets of action orders, denoted by $\mathcal{P}_{y}^{(0)}$ and $\mathcal{P}_{x}^{(0)}$, based on prior knowledge of the data.
-These sets contain all feasible action orders to be considered for responses and predictors, respectively.
-For notational convenience, we abuse $K_{y}$ and $K_{x}$ to denote the number of action orders in $\mathcal{P}_{y}^{(0)}$ and $\mathcal{P}_{x}^{(0)}$ in this section.
-To track how many times each action order is selected during the boosting process, we define the evolving action sets at iteration $v$ as $\mathcal{P}_{y}^{(v)}=\{\beta_{1}(s_{1}^{(v)}),\ldots,\beta_{K_{y}}(s_{K_{y}}^{(v)})\}$ and $\mathcal{P}_{x}^{(v)}=\{\alpha_{1}(r_{1}^{(v)}),\ldots,\alpha_{K_{x}}(r_{K_{x}}^{(v)})\}$, where $s_{k^{\prime}}^{(v)}$ and $r_{k}^{(v)}$ indicate the number of times action orders $\beta_{k^{\prime}}$ and $\alpha_{k}$ have been selected up to iteration $v$.
-We set $\mathcal{P}_{y}^{(0)}=\{\beta_{1}(0),\ldots,\beta_{K_{y}}(0)\}$ and $\mathcal{P}_{x}^{(0)}=\{\alpha_{1}(0),\ldots,\alpha_{K_{x}}(0)\}$ for initialization.
-For each selected action order $\beta_{k^{\prime}}(s_{k^{\prime}}^{(v)})\in\mathcal{P}_{y}^{(v)}$ and $\alpha_{k}(r_{k}^{(v)})\in\mathcal{P}_{x}^{(v)}$, we let $\boldsymbol{s}_{k^{\prime}}^{(v)}=s_{k^{\prime}}^{(v)}\boldsymbol{1}_{N}$ and $\boldsymbol{r}_{k}^{(v)}=r_{k}^{(v)}\boldsymbol{1}_{N}$.
-These sets and ranks are then used in Algorithm 1 to estimate $\boldsymbol{\Lambda}_{y}^{(v)}$, $\boldsymbol{\Lambda}_{x}^{(v)}$ and $\boldsymbol{\Theta}^{(v)}$, yielding the prediction $\widehat{\boldsymbol{y}}_{t}^{(v)}$.
+We first consider the case that the sets of action orders $\mathcal{P}_{x}(M,K_{x})$ and $\mathcal{P}_{y}(N,K_{y})$ are known, and their cardinalities, $K_{x}$ and $K_{y}$, are small; see, for example, the number of all possible action orders is not too large.
+A validation method can then be used to select the ranks $\{r_{m,k},1\leq m\leq M,1\leq k\leq K_{x}\}$ and $\{s_{n,k^{\prime}},1\leq n\leq N,1\leq k^{\prime}\leq K_{y}\}$.
+Specifically, for an observed time series $\{\mbox{$\mathscr{Y}$}_{t},1-L\leq t\leq T\}$, we set the training and validation sets to be the first $T_{0}+L$ and last $T-T_{0}$ observations, respectively.
+A rolling forecasting method with the fixed starting point is then employed, and we will choose the ranks with minimum validation errors.
+Note that, for an action order from $\mathcal{P}_{x}(M,K_{x})$ or $\mathcal{P}_{y}(N,K_{y})$, its associated ranks will be set to zero if it is not involved in extracting features.
+Moreover, it is impossible to search among all possible ranks since there are many interim ranks.
+We suggest selecting these ranks sequentially, i.e., we choose a rank among $1\leq r_{m,k}\leq R_{\max}$ or $1\leq s_{n,k^{\prime}}\leq S_{\max}$, while all the other ranks are kept unchanged.
+This will depend on the order of ranks, and the selected ranks are also not optimal globally.
+Practically, we may iterate the procedure several times to alleviate this effect to some extent.
+We next introduce a boosting method to select the sets of action orders, $\mathcal{P}_{x}(M,K_{x})$ and $\mathcal{P}_{y}(N,K_{y})$, as well as the maximum possible ranks, $R_{\max}$ and $S_{\max}$, in a data-driven manner.
+To this end, the Bayesian information criterion (BIC) is given below,
+$\mathrm{BIC}=T\log\left\{\mathcal{L}_{T}(\mbox{$\mathscr{\widehat{A}}$})\right\}+d_{\mathrm{AR}}\log(T),$
+where $d_{\mathrm{AR}}=\sum_{k=1}^{K_{x}}d_{x}(\alpha_{k})+\sum_{k=1}^{K_{y}}d_{y}(\beta_{k})+Lsr$ is the model complexity.
+Moreover, for an action order $\alpha=(\alpha_{(1)},\alpha_{(2)},\ldots,\alpha_{(M)})\in\mathcal{P}_{x}(M)$ with ranks $\boldsymbol{r}=(r_{1},\ldots,r_{M})$, we denote $\boldsymbol{\Lambda}_{x,\alpha}(\boldsymbol{r})=\boldsymbol{T}_{x}^{\top}(\alpha)(\boldsymbol{I}_{p_{\alpha_{(2)}}\cdots p_{\alpha_{(M)}}}\otimes\boldsymbol{G}_{1})\cdots(\boldsymbol{I}_{p_{\alpha_{(M)}}}\otimes\boldsymbol{G}_{M-1})\boldsymbol{G}_{M}\in\mathbb{R}^{P\times r_{M}}$; see (2.10) for more details.
+Consider the action order $\alpha$ with two groups of ranks, $\boldsymbol{r}_{1}$ and $\boldsymbol{r}_{2}$, and it is equivalent to the case with ranks $\boldsymbol{r}_{3}=\boldsymbol{r}_{1}+\boldsymbol{r}_{2}$, up to a linear transformation; see the proposition below.
+Proposition 2.
+Consider two vectors of ranks $\boldsymbol{r}_{1}=(r_{1,1},\ldots,r_{M,1})$ and $\boldsymbol{r}_{2}=(r_{1,2},\ldots,r_{M,2})$, and there exist a matrix $\boldsymbol{O}\in\mathbb{R}^{r_{M,3}\times(r_{M,1}+r_{M,2})}$ and the corresponding component matrices such that ${\boldsymbol{\Lambda}}_{x,\alpha}(\boldsymbol{r}_{3})=\boldsymbol{O}({\boldsymbol{\Lambda}}_{x,\alpha}(\boldsymbol{r}_{1}),{\boldsymbol{\Lambda}}_{x,\alpha}(\boldsymbol{r}_{2}))$, where $\boldsymbol{r}_{3}=\boldsymbol{r}_{1}+\boldsymbol{r}_{2}=(r_{1,3},\ldots,r_{M,3})$.
+The results are similar to $\boldsymbol{\Lambda}_{y}$.
+Let $\mathcal{P}_{x}(M,K_{x})=\{\alpha_{1}(i_{1}),\ldots,\alpha_{K_{x}}(i_{K_{x}})\}$ and $\mathcal{P}_{y}(N,K_{y})=\{\beta_{1}(j_{1}),\ldots,\beta_{K_{y}}(j_{K_{y}})\}$ be the running sets of action orders, where $\alpha_{k}(i_{k})$ refers to the action order $\alpha_{k}$ with ranks $(r_{1,k},\ldots,r_{M,k})=(i_{k},\ldots,i_{k})$, and $\beta_{k}(j_{k})$ to the action order $\beta_{k}$ with ranks $(s_{1,k},\ldots,s_{N,k})=(j_{k},\ldots,j_{k})$.
+We next introduce a procedure to update $\mathcal{P}_{x}(M,K_{x})$ and $\mathcal{P}_{y}(N,K_{y})$ by considering action orders with all ranks being one:
+  • (i.)
+    Fit a tensor autoregressive model at (2.12) and (2.13) with the running sets of $\mathcal{P}_{x}(M,K_{x})$ and $\mathcal{P}_{y}(N,K_{y})$ to the observed time series, and denote the fitted values by $\{\widehat{\mbox{$\mathscr{Y}$}}_{t}\}$.
+  • (ii.)
+    Consider a regression model with the response and predictors being $\mbox{$\mathscr{Y}$}_{t}-\widehat{\mbox{$\mathscr{Y}$}}_{t}$ and $\mbox{$\mathscr{Y}$}_{t-1},\cdots,\mbox{$\mathscr{Y}$}_{t-L}$, respectively, and we try each possible action order $\alpha\in\mathcal{P}_{x}(M)$ with ranks $(r_{1},\ldots,r_{M})=(1,\ldots,1)$ and $\beta\in\mathcal{P}_{y}(N)$ with ranks $(s_{1},\ldots,s_{N})=(1,\ldots,1)$.
+Denote by $\alpha^{*}$ and $\beta^{*}$ the action orders with minimum mean square errors.
+  • (iii.)
+    Update the sets of $\mathcal{P}_{x}(M,K_{x})$ and $\mathcal{P}_{y}(N,K_{y})$.
+Specifically, we add $\alpha^{*}(1)$ to $\mathcal{P}_{x}(M,K_{x})$ and replace $K_{x}$ by $K_{x}+1$ if $\alpha^{*}$ is a new action order.
+Otherwise, we simply replace $\alpha_{k}(i_{k})$ by $\alpha_{k}(i_{k}+1)$ if $\alpha^{*}=\alpha_{k}$.
+$\mathcal{P}_{y}(N,K_{y})$ can be updated in the same way.
+The above procedure stops when the BIC no longer decreases, and it ends up with the sets of action orders, as well as the associated maximum ranks.
+5 Simulation studies
+This section conducts two simulation experiments to evaluate action order misspecification, and to check finite-sample performance of the proposed methodology, respectively.
+The first experiment is to evaluate the impact of action order misspecification.
+Consider a true action order $\alpha^{*}$ with loading matrices $\{\boldsymbol{G}_{m}^{*}\}_{m=1}^{M}$.
+For another action order $\alpha_{k}$, by Proposition 1, there exists a set of loading matrices $\{\boldsymbol{G}_{m,k}\}_{m=1}^{M}$ such that the feature extraction mappings under both action orders are equivalent, i.e., $\mathcal{M}(\mathcal{T}(\mbox{$\mathscr{X}$}_{t};\alpha^{*});\{\boldsymbol{G}_{m}^{*}\}_{m=1}^{M})=\mathcal{M}(\mathcal{T}(\mbox{$\mathscr{X}$}_{t};\alpha_{k});\{\boldsymbol{G}_{m,k}\}_{m=1}^{M})$, but it may lead to larger ranks and hence large model complexity when $\alpha_{k}\neq\alpha^{*}$.
+To numerically verify this point, we fix $M=3$ and specify the tensor dimensions as $(p_{1},p_{2},p_{3})=(8,9,10)$.
+Note that there are six possible action orders in total, and the true action order is given by $\alpha^{*}=(1,2,3)$, with true ranks $r_{1}^{*}=r_{2}^{*}=r_{3}^{*}=2$.
+The data are generated by $\boldsymbol{f}_{t}=\mathcal{M}(\mathcal{T}(\mbox{$\mathscr{X}$}_{t};\alpha^{*});\{\boldsymbol{G}_{m}^{*}\}_{m=1}^{3})\in\mathbb{R}^{r_{3}^{*}}$, where entries of $\mbox{$\mathscr{X}$}_{t}\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}$ are $i.i.d.$ standard normal, and $\boldsymbol{G}_{m}^{*}$’s come from QR decomposition on random matrices with $i.i.d$ standard normal entries.
+We then directly regress $\boldsymbol{f}_{t}$ on $\mbox{$\mathscr{X}$}_{t}$, leading to a model without noise and $\boldsymbol{\Lambda}_{y}$ and $\boldsymbol{\Theta}$ being fixed as identity matrices.
+The sample size is set to $T=1000$, and we compare the accuracy of feature extraction with each action order $\alpha_{k}$, where $\alpha_{1}=\alpha^{*}$.
+When fitting the models with different action orders, we fix $r_{1,k}=r_{2,k}=r_{k}$ and $r_{3,k}=2$ for all $k$ for simplicity.
+Algorithm 1 is applied to estimate $\boldsymbol{G}_{m,k}$’s, and the rank $r_{k}$ varies from 2 to 8.
+Figure 2(a) gives mean squared errors (MSE), averaged over 50 replications.
 […]
-$s_{\tilde{k}^{\prime}}^{(v+1)}=s_{\tilde{k}^{\prime}}^{(v)}+1,\quad r_{\tilde{k}}^{(v+1)}=r_{\tilde{k}}^{(v)}+1$
-while keeping all other counts unchanged.
-By iteratively adding weak learners structured according to selected action orders, the model progressively refines its representational capacity in a data-adaptive and interpretable way.
-4.2.2 Rank reduction
-Algorithm 2 Boosting-Based Algorithm for Hyperparameter Selection
-Input: Data $\{(\boldsymbol{y}_{t},\boldsymbol{x}_{t})\}_{t=1}^{T}$, initial sets $\mathcal{P}_{y}^{(0)}$ and $\mathcal{P}_{x}^{(0)}$, and maximum iterations $V_{\max}$.
-Initialize: Set $\widehat{\boldsymbol{y}}_{t}^{(0)}=\boldsymbol{0}$ for all $t$ and $v=0$.
-repeat
-  Evaluate all candidate action order pairs $(\beta_{k^{\prime}},\alpha_{k})\in\{\beta_{1},\ldots,\beta_{K_{y}^{0}}\}\times\{\alpha_{1},\ldots,\alpha_{K_{x}^{0}}\}$:
-  for all $(\beta_{k^{\prime}},\alpha_{k})$ do
-   Fit weak learner $WL(\beta_{k^{\prime}},\alpha_{k})$ to regress $\boldsymbol{y}_{t}-\widehat{\boldsymbol{y}}_{t}^{(v)}$ on $\boldsymbol{x}_{t}$.
-  end for
+It can be seen that, for the true action order $\alpha_{1}$, the MSE is zero when $r_{k}$ is as small as two, and this is as expected since the model is correctly specified without noise.
+Moreover, for the misspecified action orders, the MSEs are substantially higher when $r_{k}$ is small, but decrease gradually to zero as $r_{k}$ increases.
+This is consistent with theoretical results at Proposition 1.
+On one hand, it confirms that loading matrices with different action orders can represent the same feature extraction mapping, given sufficiently large ranks.
+On the other hand, since the model complexity $d=\sum_{i=1}^{3}p_{i}r_{i}^{2}$ grows quadratically with respect to $r_{k}$, the misspecification may lead to larger model complexity and computational burden.
 […]
-  Update action counts: $s_{k^{\prime}}^{(v+1)}=s_{k^{\prime}}^{(v)}+\mathbb{I}\{k^{\prime}=\tilde{k}^{\prime}\}$, $r_{k}^{(v+1)}=r_{k}^{(v)}+\mathbb{I}\{k=\tilde{k}\}$
-  Update action sets: $\mathcal{P}_{y}^{(v+1)}=\{\beta_{k^{\prime}}(s_{k^{\prime}}^{(v+1)})\}_{k^{\prime}=1}^{K_{y}}$, $\mathcal{P}_{x}^{(v+1)}=\{\alpha_{k}(r_{k}^{(v+1)})\}_{k=1}^{K_{x}}$
-  Refit the full model using Algorithm 1 with $\mathcal{P}_{y}^{(v+1)}$ and $\mathcal{P}_{x}^{(v+1)}$ to obtain $\widehat{\boldsymbol{y}}_{t}^{(v+1)}$.
-  $v\leftarrow v+1$
-until BIC increases or $v=V_{\max}$
- Perform rank reduction: $(\mathcal{P}_{y},\mathcal{P}_{x})=\text{RR}(\mathcal{P}_{y}^{(v)},\mathcal{P}_{x}^{(v)})$
-Return: $\mbox{$\mathscr{A}$}]_{N}$ by applying Algorithm 1 with action sets $\mathcal{P}_{y}$ and $\mathcal{P}_{x}$
-According to Proposition 1, multiple combinations of action order sets $(\mathcal{P}_{y}^{(v)},\mathcal{P}_{x}^{(v)})$ may yield equivalent representations of the same model.
-Among them, we prefer the representation that achieves the lowest model complexity, as measured by BIC defined in (4.3).
-To this end, we introduce a procedure for selecting the optimal pair $(\mathcal{P}_{y},\mathcal{P}_{x})$ after the action order selection stage.
-We refer to this procedure as the rank reduction step, denoted by $\text{RR}(\cdot)$.
-The rank reduction procedure is applied to both $\mathcal{P}_{x}$ and $\mathcal{P}_{y}$.
-We illustrate the process using $\mathcal{P}_{x}$ as an example.
-The procedure iterates sequentially over action orders $\alpha_{k}$ with $k=1,\ldots,K_{x}$.
-For a given action order $\alpha_{k}$ with associated rank $\boldsymbol{r}_{k}=(r_{1,k},\ldots,r_{N,k})$, we apply Proposition 1 to search for an equivalent but more efficient representation by exploring all alternative action orders within $\mathcal{P}_{x}$.
+Finally, for the model complexity of $d_{\mathrm{AR}}$, the three hyperparameters, $r$, $q$ and $T$, may vary, while all the others have already been fixed in the above.
+As a result, we (i) vary $q\in\{10,11,12,13,14\}$ with $(r,T)=(3,2500)$ being fixed, (ii) vary $r\in\{3,4,5,6,7\}$ with $(q,T)=(10,2500)$ being fixed, or (iii) vary $T\in\{833,1000,1250,1670,2500\}$ with $(q,r)=(10,3)$ being fixed.
+Algorithm 1 is applied to search for $\mbox{$\mathscr{\widehat{A}}$}_{\mathrm{AR}}$, and Figure 2(b) presents the estimation errors $\|\mbox{$\mathscr{\widehat{A}}$}_{\mathrm{LR}}-\mbox{$\mathscr{A}$}^{*}_{\mathrm{LR}}\|_{\mathrm{F}}$, averaged over 200 replications, for each combination of error distributions and hyperparameter settings.
+The linearity observed in Figure 2(b) confirms the convergence rate, i.e., the squared estimation error depends linearly on $q$ and quadratically on $r$.
+It thereby demonstrates the efficiency of our proposed methodology.
+This paper also conducts a simulation experiment to compare our feature extracting methods with two traditional ones, hierarchical factor models (HFM) and higher-order factor models (HOFM), under the new model-designing framework.
+The results are provided in the supplementary file due to the limited space, and the consistently better performance can be observed.
+It perhaps is due to the fact that both HFM and HOFM involve a large number of parameters, demonstrating the necessity of shared components within each mode.
+6 Real data analysis
+This section applies the proposed methodology to the Personality-120 example, i.e.
+the International Personality Item Pool (IPIP) NEO-120 personality inventory (Johnson, 2014).
+The data span from 2015 to 2025, and they are available at the Hugging Face Datasets repository (https://huggingface.co/datasets/ecorbari/IPIP120-SCORES).
+There are five major personality traits (Neuroticism, Extraversion, Conscientiousness, Agreeableness, and Openness to Experience), and each trait is assessed by 24 questions arranged according to psychometrical conventions (Johnson, 2014; Kajonius and Johnson, 2019).
+The answers to all questions are recorded on a 1-5 Likert scale, and the reversely worded questions are reversely coded to ensure that higher scores indicate higher levels of the corresponding personality traits.
+Participants are categorized by gender (two groups) and age (six groups: 10–19, 20–29, 30–39, 40–49, 50–59, and 60+), resulting in 12 demographic groups.
+The monthly averaged scores are calculated, forming a tensor-valued time series $\{\mbox{$\mathscr{Y}$}_{t}\in\mathbb{R}^{12\times 5\times 24},1\leq t\leq T=115\}$.
+There is only one missing entry in the data, and it is imputed by using spline interpolation.
+For each sequence, it is first detrended by the first-order differencing to ensure stationarity, and we then standardize it to have mean zero and variance one.
+Tensor autoregressive model at (2.12) is fitted to the sequence, and we simply set the order $L=1$, ending up with the coefficient tensor $\mbox{$\mathscr{A}$}\in\mathbb{R}^{12\times 5\times 24\times 12\times 5\times 24}$.
+Note that there are only two possible action orders $\{(3,2,1),(1,3,2)\}$.
+The estimating procedure in Section 4 is employed, and the selected interim ranks are $\boldsymbol{r}_{1}=(6,5,1)$, $\boldsymbol{r}_{2}=(1,1,1)$, $\boldsymbol{s}_{1}=(6,1,1)$, and $\boldsymbol{s}_{2}=(1,1,1)$, i.e., there is one predictor and one response factors for each of two action orders.
+We focus on the two predictor factors, and Figure 3 gives their loading matrices.
+There are different interpretations for the two action orders.
+For $\alpha_{1}=(3,2,1)$, it first summarizes the 24 questions into six factors within each trait, and then extracts five features from $5\times 6=30$ factors for each demographic group, i.e., the five features can represent specific characteristics of each demographic group.
+Finally, the predictor factor is formed from $12\times 5=60$ features; see also Figure 4 for the illustration.
+On the other hand, for action order $\alpha_{2}=(1,3,2)$, it first summarizes the demographic groups, and then the two modes in the hierarchical structure.
+Since the effect of demographic groups on the questions is integrated in the first step, this order thus captures the common trend across demographic groups, and this trend concentrates on two personality traits, Agreeableness and Extraversion; see Figure 3(b) for details.
+Note that interpretations on two action orders can be well illustrated by the difference between the two loading matrices in Figure 3.
+Figure 3: Heatmaps of loading matrices for two predictor factors. The upper and lower panels correspond to action orders $\alpha_{1}=(3,2,1)$ and $\alpha_{2}=(1,3,2)$, respectively.
+We next give more details on the feature extracting progress, and the promax rotation, a type of oblique rotation, is applied to the estimated component matrices.
+This method allows for correlated factors, enhancing interpretability by simplifying factor structures, and it is commonly used in psychometrics (Hendrickson and White, 1964).
+Figure 4(d) gives the estimated coefficient matrix of $\boldsymbol{\Theta}$, and we may argue that the action order $(3,2,1)$ dominates the feature extracting process.
+As a result, we focus on the predictor factor corresponding to $\alpha_{1}=(3,2,1)$ again, and its component matrices are presented in Figure 4(a)–(c).
+Figure 4: Heatmaps of reshaped estimated component matrices of the predictor factor for action order $\alpha_{1}=(3,2,1)$ at (a)–(c) and that of the estimated coefficient matrix of $\boldsymbol{\Theta}$ at (d).
+It can be seen from Figure 4(c) that two factors, namely Item 46 (Factor $g_{1}$ for females aged 40-49) and Item 58 (Factor $g_{3}$ for females aged 60+), have relatively large absolute loadings, indicating their importance in the final feature extraction.
+For Item 46, Factor $f_{4}$ of Extraversion trait and $f_{3}$ of Openness trait exhibit large absolute loadings; see Figure 4(b).
+By identifying the significant cells corresponding to $f_{3}$ and $f_{4}$ in Figure 4(a), we can pinpoint important questions, i.e.
+Question 3 and 11 of Extraversion trait and Question 4, 7 and 13 of Openness trait.
+For the five questions, positive loadings are on “Avoid contacts with others,” “Like to get lost in thought,” “Prefer variety to routine,” and “Do not like poetry”, while a negative one is on “Take control of things”.
+These questions highlight a blend of social introversion (low extraversion in friendliness and assertiveness) with cognitive openness (high imagination and adventurousness) for middle-aged females.
+We can similar interpret Item 58.
+Refer to the supplementary file for more details on model interpretations.
+Finally, we compare the proposed methodology with existing methods in the literature in terms of forecasting performance; see Table 1.
+Note that models with Tucker decomposition, GFM and HFM will involve a large number of parameters, which makes them infeasible for this dataset.
+Moreover, Tucker decomposition cannot capture the nested structures presented in the data.
+As a result, the three models are removed from our comparison.
+For the two tensor decomposition based models, we adopt the ALS method in Zhou et al.
+(2013) for CP-based tensor autoregression and the Riemannian gradient descent method in Si et al.
+(2024) and Qin and Zhu (2025) for TT-based model.
+For the HOFM under supervised learning frameworks, we use an ALS method similar to Algorithm 1 for estimation and a coordinate descent algorithm for rank selection.
+Ranks are selected by their respective methods, and it leads to $r_{\mathrm{CP}}=1$ for the CP, $\boldsymbol{r}_{\mathrm{TT}}=(1,2,2,3,2,1,1)$ for the TT, and $\boldsymbol{r}_{\mathrm{HOFM}}=(3,1,6)$ for predictors and $\boldsymbol{s}_{\mathrm{HOFM}}=(4,1,1)$ for responses in the HOFM.
+Table 2: Mean squared forecast errors (MSFE) and mean absolute forecast errors (MAFE) of our model and three competing methods. The smallest numbers in each row are in bold.
+Models New framework Traditional framework
+Ours HOFM CP $T$
+MSFE 1530.4 1644.6 1577.1 1571.1
+MAFE 1153.0 1185.8 1156.0 1164.9
+We divide the 115 observations into a training set with the first 103 months and a testing set with the last 12 months.
+The rolling forecast method is employed, i.e., we fit models to historical data up to month $t$ and then forecast the data at month $t+1$ for $t=103,\ldots,114$.
+The one-step-ahead forecast accuracy is assessed by the mean squared forecast errors (MSFE) and mean absolute forecast errors (MAFE); see Table 2 for the results.
+It can be seen that our model outperforms all competing methods in both MSFE and MAFE metrics.
+This may be due to the existence of hierarchical structures in the data, which can be effectively captured by our model.
+Moreover, the superior performance of our method against TT-based model highlights the advantage of our model by considering multiple action orders for partially hierarchical data.
+The better performance compared to HOFM-based method demonstrates the benefit of improved efficiency in feature extraction by sharing components within each mode in loading matrices.
+In sum, the proposed model provides an efficient and interpretable approach for modeling tensor-valued time series with hierarchical structures.
+7 Conclusion and discussion
+With technological advancements, high-dimensional data with hierarchical structures have been observed more frequently, while it is usually infeasible to construct inference tools for them due to a huge number of parameters.
+Fortunately, most of these data come with regular structures, say in tensor forms, and the variables within each mode are arranged in a similar way, leading to a seemingly factorial but hierarchical structure.
+This makes it possible to design efficient and interpretable inference tools for them, and this paper focuses on regression and autoregressive models for supervised learning problems.
+Theoretical properties have been established for the corresponding ordinary least squares estimation, and an alternating algorithm is introduced to search for estimates.
+Applications to the Personality-120 example have demonstrated usefulness of the proposed methodology, and a nice interpretation is given to the use of multiply action orders for partially hierarchical structures.
+This paper can be extended in three possible directions.
+First, missing data will be presented when we consider more detailed demographic groups, and this actually is common in psychometrics (Johnson, 2014).
+It is an important task to adjust the proposed methodology to handle the case with missing data.
+Secondly, due to the availability of more and more large-size datasets, it becomes urgent to consider more sophisticated models for better performance, and neural networks have played a key role in the literature (Chen et al., 2025).
+This motivates us to design a suitable neural network with theoretical justifications for these hierarchical data.
+Finally, it is well known that, comparing with AR models, the autoregressive moving average (ARMA) model can achieve more accurate prediction, and it is interesting to extend the AR model here to an ARMA-type one; see, e.g., Zheng (2025).
+Supplementary Material for “High-dimensional Autoregressive Modeling for Time Series Data with Hierarchical Structures”
+Lan Li, Shibo Yu, Yingzhou Wang and Guodong Li
+Department of Statistics and Actuarial Science, University of Hong Kong April 2, 2026
+Appendix A Tensor Notation and Algebra
+Tensors, a.k.a.
+multidimensional arrays, are natural high order extensions of matrices, and the order of a tensor is known as the dimension, way or mode; see Kolda and Bader (2009) for a review of basic tensor algebra.
+A $d$-th order tensor $\mbox{$\mathscr{X}$}\in\mathbb{R}^{p_{1}\times p_{2}\times\cdots\times p_{d}}$ is defined as a $d$-dimensional array, where each order is called a mode, and its element is denoted by $\mbox{$\mathscr{X}$}_{i_{1}i_{2}\ldots i_{d}}$ or $\mbox{$\mathscr{X}$}_{i_{1},i_{2},\ldots,i_{d}}$ for $1\leq i_{l}\leq p_{l}$ with $1\leq l\leq d$.
+For a $d$-th order tensor $\mbox{$\mathscr{X}$}\in\mathbb{R}^{p_{1}\times p_{2}\times\cdots\times p_{d}}$ and a matrix $\boldsymbol{Y}\in\mathbb{R}^{q_{k}\times p_{k}}$, their mode-$k$ multiplication is defined as $\mbox{$\mathscr{X}$}\times_{k}\boldsymbol{Y}\in\mathbb{R}^{p_{1}\times\cdots\times p_{k-1}\times q_{k}\times p_{k+1}\times\cdots\times p_{d}}$ with elements of
+$(\mbox{$\mathscr{X}$}\times_{k}\boldsymbol{Y})_{i_{1},\ldots,i_{k-1},j,i_{k+1},\ldots,i_{d}}=\sum_{i_{k}=1}^{p_{k}}\mbox{$\mathscr{X}$}_{i_{1},\ldots,i_{d}}\boldsymbol{Y}_{ji_{k}},$
+where $1\leq j\leq q_{k}$, and $1\leq i_{l}\leq p_{l}$ with $1\leq l\leq d$ and $l\neq k$.
+Given another (dm)(d-m)-th order tensor $\mbox{$\mathscr{Z}$}\in\mathbb{R}^{p_{m+1}\times p_{m+2}\times\cdots\times p_{d}}$, the generalized inner product of $\mathscr{X}$ and $\mathscr{Z}$ is defined as $\langle\mbox{$\mathscr{X}$},\mbox{$\mathscr{Z}$}\rangle\in\mathbb{R}^{p_{1}\times\cdots\times p_{m}}$ with elements of
+$\langle\mbox{$\mathscr{X}$},\mbox{$\mathscr{Z}$}\rangle_{i_{1},\ldots,i_{m}}=\sum_{i_{m+1}=1}^{p_{m+1}}\sum_{i_{m+2}=1}^{p_{m+2}}\cdots\sum_{i_{d}=1}^{p_{d}}\mbox{$\mathscr{X}$}_{i_{1},i_{2},\ldots,i_{d}}\mbox{$\mathscr{Z}$}_{i_{m+1},i_{m+2},\ldots,i_{d}}\text{ for $1\leq i_{l}\leq p_{l}$ with $1\leq l\leq m$}.$
+When $m=0$, it will reduce to a scalar, and then the generalized inner product becomes an inner product.
+Moreover, the Frobenius norm of $\mathscr{X}$ is defined as $\left\lVert\mbox{$\mathscr{X}$}\right\rVert_{\mathrm{F}}=\sqrt{\langle\mbox{$\mathscr{X}$},\mbox{$\mathscr{X}$}\rangle}$.
+Matricization or unfolding is an operation to reshape a tensor into matrices of different sizes, and this paper will involve two types of matricization: mode matricization and sequential matricization.
+Mode-$s$ matricization sets the $s$-th mode as rows, and columns enumerate the rest modes.
+Specifically, let $p_{-s}=\prod_{i=1,i\neq s}^{d}p_{i}$, and the tensor $\mathscr{X}$ is reshaped into a $p_{s}$-by-$p_{-s}$ matrix, denoted by $\mbox{$\mathscr{X}$}_{(s)}$, where the element $\mbox{$\mathscr{X}$}_{i_{1},i_{2},\ldots,i_{d}}$ is mapped to $(i_{s},j)$-th element of $\mbox{$\mathscr{X}$}_{(s)}$ with
+$j=1+\sum_{k=1,k\neq s}^{d}(i_{k}-1)J_{k}\hskip 5.69054pt\text{and}\hskip 5.69054ptJ_{k}=\prod_{l=1,l\neq s}^{k-1}p_{l}.$
+Sequential matricization reshapes tensor $\mathscr{X}$ into a $\prod_{i=1}^{s}p_{i}$-by-$\prod_{i=s+1}^{d}p_{i}$ matrix, denoted by $\mbox{$\mathscr{X}$}]_{s}$, where rows enumerate all indices from modes 1 to $s$, and columns enumerate modes $s+1$ to $d$.
+Specifically, $\mbox{$\mathscr{X}$}_{i_{1},i_{2},\ldots,i_{d}}$ is mapped to $(i,j)$-th element of $\mbox{$\mathscr{X}$}]_{s}$ with
+$i=(i_{s}-1)p_{1}\cdots p_{s-1}+(i_{s-1}-1)p_{1}\cdots p_{s-2}+\cdots+i_{1}\hskip 5.69054pt\text{and}\hskip 5.69054ptj=(i_{d}-1)p_{s+1}\cdots p_{d-1}+\cdots+i_{s+1}.$
+We also define the Reshape(;)($\cdot$$\cdot$) operation such that $\text{Reshape}(\boldsymbol{a};r,p)\in\mathbb{R}^{r\times p}$ for $\boldsymbol{a}\in\mathbb{R}^{rp}$ and $\operatorname*{vec}(\text{Reshape}(\boldsymbol{a};r,p))=\boldsymbol{a}$.
+Appendix B Additional Details on Model Properties and Algorithm
+This section provides additional details supporting the results presented in Section 3.
+Specifically, we elaborate on the derivation of certain conclusions and describe the Sequential Singular Salue Decomposition (SSVD) procedure used to orthogonalize the components in our proposed model, as implemented in Algorithm 1.
+B.1 Proof of Proposition 1
+We restate Proposition 1 in a more convenient way:
+Proposition 3.
+Specify any action order $\alpha=(\alpha_{(1)},\alpha_{(2)},\dots,\alpha_{(M)})\in\mathcal{P}(M)$ and another target action order $\alpha^{\prime}=(\alpha^{\prime}_{(1)},\alpha^{\prime}_{(2)},\dots,\alpha^{\prime}_{(M)})\in\mathcal{P}(M)$.
+Then, to any set of core matrices $\{\boldsymbol{G}_{m}\in\mathbb{O}^{r_{m-1}p_{m}\times r_{m}}\}_{m=1}^{M}$ there corresponds a set of core matrices $\{\boldsymbol{G}^{\prime}_{m}\in\mathbb{O}^{r_{m-1}^{\prime}p^{\prime}_{m}\times r_{m}^{\prime}}\}_{m=1}^{M}$ such that $\mathcal{M}(\mathcal{T}(\cdot;\alpha);\{\boldsymbol{G}_{m}\}_{m=1}^{M})=\mathcal{M}(\mathcal{T}(\cdot;\alpha^{\prime});\{\boldsymbol{G}^{\prime}_{m}\}_{m=1}^{M})$.
+Proof.
+We begin by noting that the group action of switching the order of two adjacent elements in the sequence of actions $\alpha$ makes the permutation set $\mathcal{P}(M)$ a permutation group.
+This means that any permutation of the action order can be achieved by a series of adjacent swaps and thus we may assume $\alpha=(1,2,3,\dots,M)$ and $\alpha^{\prime}=(1,2,\dots,k+1,k,k+2,\dots,M)$ without loss of generality.
+We abbreviate the operator $\mathcal{T}(\cdot;\alpha^{\prime})$ as $\mathcal{T}$.
+Let $\boldsymbol{G}_{M}^{\top}(\boldsymbol{I}_{p_{M}}\otimes\boldsymbol{G}_{M-1}^{\top})(\boldsymbol{I}_{p_{M}p_{M-1}}\otimes\boldsymbol{G}_{M-2}^{\top})\cdots(\boldsymbol{I}_{p_{M}p_{M-1}\cdots p_{2}}\otimes\boldsymbol{G}_{1}^{\top})$ be the matrix corresponding to the operator $\mathcal{M}(\mathcal{T}(\cdot;\alpha);\{\boldsymbol{G}_{m}\}_{m=1}^{M})$, where each $\boldsymbol{G}_{m}\in\mathbb{O}^{p_{m}r_{m-1}\times r_{m}}$ is the orthogonal matrix corresponding to the $m$-th action in the order $\alpha$, and $r_{m}$ is the extraction size for the $m$-th interim feature.
+By assumption, $p_{1}\times p_{2}\times\cdots\times p_{M}$ is the size of the input tensor to be processed.
+We are left to show that there exist orthogonal matrices $\boldsymbol{G}_{m}^{\prime}\in\mathbb{O}^{p^{\prime}_{m}r_{m-1}^{\prime}\times r_{m}^{\prime}},m=1,2,\dots,M$ with $r_{m}^{\prime}\leq p^{\prime}_{m}r_{m-1}^{\prime}$ such that
+$\begin{split}\underbrace{\boldsymbol{G}_{M}^{\prime\top}(\boldsymbol{I}_{p_{M}}\otimes\boldsymbol{G}_{M-1}^{\prime\top})(\boldsymbol{I}_{p_{M}p_{M-1}}\otimes\boldsymbol{G}_{M-2}^{\prime\top})\dots}_{=:\boldsymbol{A}^{\prime}}(\boldsymbol{I}_{p_{M}p_{M-1}\dots p_{k+2}}\otimes\boldsymbol{H}_{k}^{\top})(\boldsymbol{I}_{p_{M}p_{M-1}\dots p_{k+2}p_{k}}\otimes\boldsymbol{H}_{k+1}^{\top})\\ \underbrace{(\boldsymbol{I}_{p_{M}p_{M-1}\dots p_{k+1}p_{k}}\otimes\boldsymbol{G}_{k-1}^{\prime\top})\cdots(\boldsymbol{I}_{p_{M}p_{M-1}\cdots p_{2}}\otimes\boldsymbol{G}_{1}^{\prime\top})}_{=:\boldsymbol{C}^{\prime}}\boldsymbol{T}\\ =\underbrace{\boldsymbol{G}_{M}^{\top}(\boldsymbol{I}_{p_{M}}\otimes\boldsymbol{G}_{M-1}^{\top})(\boldsymbol{I}_{p_{M}p_{M-1}}\otimes\boldsymbol{G}_{M-2}^{\top})\dots}_{=:\boldsymbol{A}}(\boldsymbol{I}_{p_{M}p_{M-1}\dots p_{k+2}}\otimes\boldsymbol{G}_{k+1}^{\top})(\boldsymbol{I}_{p_{M}p_{M-1}\dots p_{k+1}}\otimes\boldsymbol{G}_{k}^{\top})\\ \underbrace{(\boldsymbol{I}_{p_{M}p_{M-1}\dots p_{k+1}p_{k}}\otimes\boldsymbol{G}_{k-1}^{\top})\cdots(\boldsymbol{I}_{p_{M}p_{M-1}\cdots p_{2}}\otimes\boldsymbol{G}_{1}^{\top})}_{=:\boldsymbol{C}}\end{split}$ (B.1)
+where $\boldsymbol{T}$ is the matrix corresponding to the operator $\mathcal{T}$, and $\boldsymbol{H}_{k}=\boldsymbol{G}^{\prime}_{k+1},\boldsymbol{H}_{k+1}=\boldsymbol{G}^{\prime}_{k}$ for notational convenience.
+Denote $\mathcal{C}$ and $\mathcal{C}^{\prime}$ as the operators corresponding to the matrices $\boldsymbol{C}$ and $\boldsymbol{C}^{\prime}$, respectively, i.e., $\text{vec}(\mathcal{C}\mbox{$\mathscr{X}$})=\boldsymbol{C}\text{vec}(\mbox{$\mathscr{X}$})$ and $\text{vec}(\mathcal{C}^{\prime}\mbox{$\mathscr{X}$}^{\prime})=\boldsymbol{C}^{\prime}\text{vec}(\mbox{$\mathscr{X}$}^{\prime})$.
+Note that we omit the parentheses between the operator and the tensor for simplicity.
+We first want to show that for any $1\leq i_{k-1}\leq r_{k-1}$, and $1\leq i_{m}\leq p_{m}$ for $m=k,k+1,\dots,M$, and for all $\mbox{$\mathscr{X}$}\in\mathbb{R}^{p_{1}\times p_{2}\times\cdots\times p_{M}}$, the following holds:
+$(\mathcal{CT}\mbox{$\mathscr{X}$})_{i_{k-1},i_{k+1},i_{k},i_{k+2},\dots,i_{M}}=(\mathcal{C}\mbox{$\mathscr{X}$})_{i_{k-1},i_{k},i_{k+1},i_{k+2},\dots,i_{M}}$ (B.2)
+and thereby $\mathcal{C^{\prime}}$ can be constructed as $\mathcal{C}$, i.e., $\boldsymbol{G}_{m}^{\prime}=\boldsymbol{G}_{m}$ for $m=1,2,\dots,k-1$.
+We begin by first observing that $(\boldsymbol{I}_{p_{M}p_{M-1}\cdots p_{2}}\otimes\boldsymbol{G}_{1}^{\top})\operatorname*{vec}(\mathcal{T}\mbox{$\mathscr{X}$})=\operatorname*{vec}(\boldsymbol{G}_{1}^{\top}[\mathcal{T}\mbox{$\mathscr{X}$}]_{1})$.
+Let $\mbox{$\mathscr{X}$}^{(1)}\in\mathbb{R}^{r_{1}\times p_{2}\times\cdots\times p_{M}},\mbox{$\mathscr{X}$}^{\prime(1)}\in\mathbb{R}^{r_{1}\times\cdots p_{k-1}\times p_{k+1}\times p_{k}\times p_{k+2}\cdots\times p_{M}}$ be the tensors such that $\operatorname*{vec}(\mbox{$\mathscr{X}$}^{(1)})=(\boldsymbol{I}_{p_{M}p_{M-1}\cdots p_{2}}\otimes\boldsymbol{G}_{1}^{\top})\operatorname*{vec}(\mbox{$\mathscr{X}$})$, and $\operatorname*{vec}(\mbox{$\mathscr{X}$}^{\prime(1)})=(\boldsymbol{I}_{p_{M}p_{M-1}\cdots p_{2}}\otimes\boldsymbol{G}_{1}^{\top})\operatorname*{vec}(\mathcal{T}\mbox{$\mathscr{X}$})$.
+Then, we have
+$(\mbox{$\mathscr{X}$}^{\prime(1)})_{i_{1},\dots,i_{k-1},i_{k+1},i_{k},i_{k+2},\dots,i_{M}}$ $=\sum_{j_{1}=1}^{p_{1}}(\boldsymbol{G}_{1}^{\top})_{i_{1},j_{1}}(\mathcal{T}\mbox{$\mathscr{X}$})_{j_{1},i_{2},\dots,i_{k-1},i_{k+1},i_{k},i_{k+2},\dots,i_{M}}$
+$(\mbox{$\mathscr{X}$}^{(1)})_{i_{1},i_{2},\dots,i_{M}}$ $=\sum_{j_{1}=1}^{p_{1}}(\boldsymbol{G}_{1}^{\top})_{i_{1},j_{1}}\mbox{$\mathscr{X}$}_{j_{1},i_{2},\dots,i_{M}}.$
+But by the definition of $\mathcal{T}$, $(\mathcal{T}\mbox{$\mathscr{X}$})_{i_{1},i_{2},\dots,i_{k-1},i_{k+1},i_{k},i_{k+2},\dots,i_{M}}=\mbox{$\mathscr{X}$}_{i_{1},i_{2},\dots,i_{M}}$ for all applicable $i_{m}$, and thus we have
+$(\mbox{$\mathscr{X}$}^{\prime(1)})_{i_{1},\dots,i_{k-1},i_{k+1},i_{k},i_{k+2},\dots,i_{M}}$ $=(\mbox{$\mathscr{X}$}^{(1)})_{i_{1},i_{2},\dots,i_{M}}.$
+Now if we combine the first two modes and let $\mbox{$\mathscr{X}$}^{(2)}\in\mathbb{R}^{r_{1}p_{2}\times p_{3}\cdots\times p_{M}}$ to be the tensor such that $\operatorname*{vec}(\mbox{$\mathscr{X}$}^{(2)})=\operatorname*{vec}(\mbox{$\mathscr{X}$}^{(1)})$ and apply the same reasoning, we can show (B.2).
+The intuition of the first step is that, despite the change of order of certain modes, the operator $\mathcal{C}$ changes the values of the modes that are not involved in the swap in the same way.
+Next, we want to show that there exist orthogonal matrices $\boldsymbol{H}_{k+1}\in\mathbb{R}^{p_{k+1}r_{k-1}\times s_{k}}$ and $\boldsymbol{H}_{k}\in\mathbb{R}^{s_{k}p_{k}\times r_{k+1}}$ such that the following holds:
+$(\boldsymbol{I}_{p_{M}p_{M-1}\dots p_{k+2}}\otimes\boldsymbol{H}_{k}^{\top})(\boldsymbol{I}_{p_{M}p_{M-1}\dots p_{k+2}p_{k}}\otimes\boldsymbol{H}_{k+1}^{\top})\boldsymbol{T}^{\prime}=(\boldsymbol{I}_{p_{M}p_{M-1}\dots p_{k+2}}\otimes\boldsymbol{G}_{k+1}^{\top})(\boldsymbol{I}_{p_{M}p_{M-1}\dots p_{k+1}}\otimes\boldsymbol{G}_{k}^{\top}),$ (B.3)
+where $\boldsymbol{T}^{\prime}$ is a square matrix corresponding to the operator $\mathcal{T}^{\prime}:=\mathcal{T}(\cdot;(k-1,k+1,k,k+2,\dots,M))$, and $s_{k}:=r^{\prime}_{k-1}<p_{k+1}r_{k-1}$ is the surrogate of $r_{k}$ to be determined later.
+Re-define $\mbox{$\mathscr{X}$}\in\mathbb{R}^{r_{k-1}\times p_{k}\times p_{k+1}\times p_{res}}$ as the tensor sharing the same vectorization as the output of $\mathcal{C}$, where $p_{res}=p_{k+2}\cdots p_{M}$, and $\mbox{$\mathscr{X}$}^{\prime}:=\mathcal{T}^{\prime}\mbox{$\mathscr{X}$}\in\mathbb{R}^{r_{k-1}\times p_{k+1}\times p_{k}\times p_{res}}$.
+Denote
+$\boldsymbol{m}:=\operatorname*{vec}(\boldsymbol{G}_{k+1}^{\top}\text{Reshape}(\boldsymbol{G}_{k}^{\top}[\mbox{$\mathscr{X}$}]_{2};r_{k}p_{k+1},p_{res}))\in\mathbb{R}^{r_{k+1}p_{res}}$
+and
+$\boldsymbol{m}^{\prime}:=\operatorname*{vec}(\boldsymbol{H}_{k}^{\top}\text{Reshape}(\boldsymbol{H}_{k+1}^{\top}[\mbox{$\mathscr{X}$}^{\prime}]_{2};s_{k}p_{k},p_{res}))\in\mathbb{R}^{r_{k+1}p_{res}}.$
+In this step, we take the $l$-th element of the vector $\boldsymbol{m}$ and $\boldsymbol{m}^{\prime}$, denoted by $\boldsymbol{m}_{l}$ and $\boldsymbol{m}^{\prime}_{l}$, and suppose $l=a\cdot r_{k+1}+b$, where $a=\left\lfloor\frac{l-1}{r_{k+1}}\right\rfloor$ and $b\in\{1,\dots,r_{k+1}\}$ are the quotient and positive remainder of the division of $l$ by $r_{k+1}$, respectively.
+It can be shown that
+$\boldsymbol{m}_{l}$ $=\boldsymbol{G}_{k+1,b}^{\top}\operatorname*{vec}\left(\sum\nolimits_{i=1}^{r_{k-1}}\boldsymbol{G}_{k,i}^{\top}\mbox{$\mathscr{X}$}_{i,:,:,a}\right)=\operatorname*{tr}\left(\widetilde{\boldsymbol{G}}_{k+1,b}\sum\nolimits_{i=1}^{r_{k-1}}\boldsymbol{G}_{k,i}^{\top}\mbox{$\mathscr{X}$}_{i,:,:,a}\right)$ (B.4)
+$\boldsymbol{m}^{\prime}_{l}$ $=\boldsymbol{H}_{k,b}^{\top}\operatorname*{vec}\left(\sum\nolimits_{i=1}^{r_{k-1}}\boldsymbol{H}_{k+1,i}^{\top}\mbox{$\mathscr{X}$}^{\prime}_{i,:,:,a}\right)=\operatorname*{tr}\left(\widetilde{\boldsymbol{H}}_{k,b}\sum\nolimits_{i=1}^{r_{k-1}}\boldsymbol{H}_{k+1,i}^{\top}\mbox{$\mathscr{X}$}_{i,:,:,a}^{\top}\right)$
+$\hskip 175.83803pt=\operatorname*{tr}\left(\sum\nolimits_{i=1}^{r_{k-1}}\boldsymbol{H}_{k+1,i}\widetilde{\boldsymbol{H}}_{k,b}^{\top}\mbox{$\mathscr{X}$}_{i,:,:,a}\right)$ (B.5)
+where $\widetilde{\boldsymbol{G}}_{k+1,b}\in\mathbb{R}^{p_{k+1}\times r_{k}}$ and $\widetilde{\boldsymbol{H}}_{k,b}\in\mathbb{R}^{p_{k}\times s_{k}}$ can be vectorized into $\boldsymbol{G}_{k+1,b}$ and $\boldsymbol{H}_{k,b}$, which stand for the $b$-th column of $\boldsymbol{G}_{k+1}$ and $\boldsymbol{H}_{k}$, respectively; and $\boldsymbol{G}_{k,i}\in\mathbb{R}^{p_{k}\times r_{k}}$ and $\boldsymbol{H}_{k+1,i}\in\mathbb{R}^{p_{k+1}\times s_{k}}$ for $i\in\{1,2,\dots,r_{k-1}\}$ are formed by the rows of $\boldsymbol{G}_{k}$ and $\boldsymbol{G}_{k+1}$ such that (B.4) and (B.5) hold.
+We claim that
+$\langle\widetilde{\boldsymbol{G}}_{k+1,b},\widetilde{\boldsymbol{G}}_{k+1,c}\rangle=\langle\widetilde{\boldsymbol{H}}_{k,b},\widetilde{\boldsymbol{H}}_{k,c}\rangle=\delta_{bc}=\begin{cases}1,&b=c\\ 0,&b\neq c\end{cases}$ (B.6)
+holds for all $b,c\in\{1,\dots,r_{k+1}\}$, which is equivalent to the orthogonality of $\boldsymbol{G}_{k+1}$ and $\boldsymbol{H}_{k}$.
+The former part is assumed, and we later prove the latter.
+Thus, we need to show that to any orthogonal matrices $\boldsymbol{G}_{k+1}$ and $\boldsymbol{G}_{k}$ there correspond an orthogonal matrix $\boldsymbol{H}_{k+1}$ and a set of matrices $\widetilde{\boldsymbol{H}}_{k,b}$ satisfying (B.6) such that $\boldsymbol{m}_{l}=\boldsymbol{m}^{\prime}_{l}$, i.e.,
+$\widetilde{\boldsymbol{G}}_{k+1,b}\boldsymbol{G}_{k,i}^{\top}=\boldsymbol{H}_{k+1,i}\widetilde{\boldsymbol{H}}_{k,b}^{\top}\in\mathbb{R}^{p_{k+1}\times p_{k}}$ (B.7)
+holds for all $(b,i)\in\{1,\dots,r_{k+1}\}\times\{1,\dots,r_{k-1}\}$, where $\boldsymbol{G}_{k,i}\in\mathbb{R}^{p_{k}\times r_{k}}$ and $\boldsymbol{H}_{k+1,i}\in\mathbb{R}^{p_{k+1}\times s_{k}}$ are defined as above.
+Let $\boldsymbol{B}_{b}=\big[\boldsymbol{G}_{k,1}\widetilde{\boldsymbol{G}}_{k+1,b}^{\top};\dots;\boldsymbol{G}_{k,r_{k-1}}\widetilde{\boldsymbol{G}}_{k+1,b}^{\top}\big]^{\top}\in\mathbb{R}^{p_{k+1}r_{k-1}\times p_{k}}$ be the LHS of (B.7) stacked up, and set $V=\text{span}\left\{\text{col}(\boldsymbol{B}_{b}):b=1,\dots,r_{k+1}\right\}\subset\mathbb{R}^{p_{k+1}r_{k-1}}$.
+Set
+$s_{k}=\dim V.$ (B.8)
+By the construction of $\boldsymbol{B}_{b}$ we know that
+$\boldsymbol{B}_{b}=\left[\begin{array}[]{c}\widetilde{\boldsymbol{G}}_{k+1,b}\boldsymbol{G}_{k,1}^{\top}\\ \widetilde{\boldsymbol{G}}_{k+1,b}\boldsymbol{G}_{k,2}^{\top}\\ \vdots\\ \widetilde{\boldsymbol{G}}_{k+1,b}\boldsymbol{G}_{k,r_{k-1}}^{\top}\end{array}\right]=\left(\boldsymbol{I}_{r_{k-1}}\otimes\widetilde{\boldsymbol{G}}_{k+1,b}\right)\left[\begin{array}[]{c}\boldsymbol{G}_{k,1}^{\top}\\ \boldsymbol{G}_{k,2}^{\top}\\ \vdots\\ \boldsymbol{G}_{k,r_{k-1}}^{\top}\end{array}\right]\vcentcolon=(\boldsymbol{I}\otimes\widetilde{\boldsymbol{G}}_{k+1,b})\boldsymbol{B}_{\natural},$ (B.9)
+where $\boldsymbol{B}_{\natural}\in\mathbb{R}^{r_{k}r_{k-1}\times p_{k}}$ is not guaranteed to be orthogonal.
+Therefore, $\operatorname*{rank}(\boldsymbol{B}_{b})\leq r_{k}r_{k-1}$ and $s_{k}\leq r_{k-1}\min\{r_{k}r_{k+1},p_{k+1}\}$.
+Construct the matrix $\boldsymbol{H}_{k+1}=\left[\boldsymbol{H}_{k+1,1}^{\top};\dots;\boldsymbol{H}_{k+1,r_{k-1}}^{\top}\right]^{\top}\in\mathbb{O}^{p_{k+1}r_{k-1}\times s_{k}}$ so that $\boldsymbol{H}_{k+1}\boldsymbol{H}_{k+1}^{\top}=\mathcal{P}_{V}$, i.e., columns of $\boldsymbol{H}_{k+1}$ constitute an orthonormal basis of $V$, by, for example, Gram-Schmidt orthogonalization, and $\boldsymbol{H}_{k+1,i}\in\mathbb{R}^{p_{k+1}\times s_{k}}$ by chopping $\boldsymbol{H}_{k+1}$ accordingly.
+Consequently, set
+$\widetilde{\boldsymbol{H}}_{k,b}=\boldsymbol{B}_{b}^{\top}\boldsymbol{H}_{k+1}\in\mathbb{R}^{p_{k}\times s_{k}}.$ (B.10)
+It remains to verify the orthogonality condition (B.6) and the central equality (B.7) before we can conclude the proof by setting $\boldsymbol{G}_{m}^{\prime}=\boldsymbol{G}_{m}$ for $m=k+2,\dots,M$ so that $\boldsymbol{A}^{\prime}=\boldsymbol{A}$ in (B.1).
+Note that $\boldsymbol{H}_{k+1}\widetilde{\boldsymbol{H}}_{k,b}^{\top}=\boldsymbol{H}_{k+1}\boldsymbol{H}_{k+1}^{\top}\boldsymbol{B}_{b}=\boldsymbol{B}_{b}$, and taking the $i$-th block (component) out from both sides, we have
+$\boldsymbol{H}_{k+1,i}\widetilde{\boldsymbol{H}}_{k,b}^{\top}=\left[\boldsymbol{B}_{b}\right]_{\text{block i}}=\widetilde{\boldsymbol{G}}_{k+1,b}\boldsymbol{G}_{k,i}^{\top},$ (B.11)
+whence (B.7) holds.
+Similarly, $\operatorname*{tr}(\widetilde{\boldsymbol{H}}_{k,b}\widetilde{\boldsymbol{H}}_{k,c}^{\top})=\operatorname*{tr}(\boldsymbol{B}_{b}^{\top}\boldsymbol{H}_{k+1}\boldsymbol{H}_{k+1}^{\top}\boldsymbol{B}_{c})=\operatorname*{tr}(\boldsymbol{B}_{b}^{\top}\boldsymbol{B}_{c})$, but
+$\operatorname*{tr}(\boldsymbol{B}_{b}^{\top}\boldsymbol{B}_{c})$ $=\operatorname*{tr}\left(\left(\boldsymbol{I}_{r_{k-1}}\otimes\widetilde{\boldsymbol{G}}_{k+1,b}\right)\boldsymbol{B}_{\natural}\boldsymbol{B}_{\natural}^{\top}\left(\boldsymbol{I}_{r_{k-1}}\otimes\widetilde{\boldsymbol{G}}_{k+1,c}\right)^{\top}\right)$
+$=\operatorname*{tr}\left((\boldsymbol{I}_{r_{k-1}}\otimes\widetilde{\boldsymbol{G}}_{k+1,b}^{\top}\widetilde{\boldsymbol{G}}_{k+1,c})\boldsymbol{B}_{\natural}\boldsymbol{B}_{\natural}^{\top}\right)$
+$=\sum_{i=1}^{r_{k-1}}\operatorname*{tr}\left(\widetilde{\boldsymbol{G}}_{k+1,b}^{\top}\widetilde{\boldsymbol{G}}_{k+1,c}\boldsymbol{G}_{k,i}^{\top}\boldsymbol{G}_{k,i}\right)$
+$=\operatorname*{tr}\left(\widetilde{\boldsymbol{G}}_{k+1,b}^{\top}\widetilde{\boldsymbol{G}}_{k+1,c}\sum_{i}\boldsymbol{G}_{k,i}^{\top}\boldsymbol{G}_{k,i}\right)$
+$=\operatorname*{tr}\left(\widetilde{\boldsymbol{G}}_{k+1,b}^{\top}\widetilde{\boldsymbol{G}}_{k+1,c}\boldsymbol{G}_{k}^{\top}\boldsymbol{G}_{k}\right)$
+$=\delta_{bc},$
+which confirms the orthogonality condition (B.6). ∎
+B.2 ALS update rules
+Recall that the model is given by
+$\boldsymbol{y}_{t}=\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top})\boldsymbol{x}_{t}+\boldsymbol{e}_{t},$
+and the update rule comprises two parts: updating the component matrices sequentially by converting to a regression problem (minimization), and imposing the orthogonality constraints on the updated matrices (orthogonalization).
+When updating $\boldsymbol{G}_{m,k}$ in $\boldsymbol{\Lambda}_{x}$, denote $\boldsymbol{y}_{t,k}=\boldsymbol{y}_{t}-\sum_{l=1}^{L}\boldsymbol{\Lambda}_{y}\sum_{\gamma\neq k}^{K_{x}}\boldsymbol{\Theta}_{l,\gamma}\boldsymbol{\Lambda}_{x,\gamma}^{\top}$ as the response variable with the contributions of other action orders deducted, where $\boldsymbol{\Lambda}_{x,\gamma}$ is the $\gamma$’s block column of $\boldsymbol{\Lambda}_{x}$ and $\boldsymbol{\Theta}_{l,\gamma}$ is the $\gamma$’s block column of $\boldsymbol{\Theta}_{l}$ corresponding to the action order $k$.
+The update for $\boldsymbol{G}_{m,k}$ is then converted to the following OLS problem of linear regression:
+$\boldsymbol{y}_{t,k}=\bigg[\sum_{l=1}^{L}\boldsymbol{B}_{l,m,k}(\boldsymbol{Y}_{t-l,m,k}^{\top}\otimes\boldsymbol{I}_{r_{m}})\bigg]\operatorname*{vec}(\boldsymbol{G}_{m,k}^{\top})+\boldsymbol{e}_{t},$
+where $\boldsymbol{B}_{l,m,k}=\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}_{l,k}\boldsymbol{G}_{M,k}^{\top}(\boldsymbol{I}_{p_{\alpha^{k}_{(M)}}}\otimes\boldsymbol{G}_{M-1,k}^{\top})\cdots(\boldsymbol{I}_{p_{\alpha^{k}_{(m+2)}}\cdots p_{\alpha^{k}_{(M)}}}\otimes\boldsymbol{G}_{m+1,k}^{\top})$ accounts for the matrices before $\boldsymbol{G}_{m,k}$, and $\boldsymbol{Y}_{t-l,m,k}=\text{Reshape}((\boldsymbol{I}_{p_{\alpha^{k}_{(m)}}\cdots p_{\alpha^{k}_{(M)}}}\otimes\boldsymbol{G}_{m-1,k}^{\top})\cdots(\boldsymbol{I}_{p_{\alpha^{k}_{(2)}}\cdots p_{\alpha^{k}_{(M)}}}\otimes\boldsymbol{G}_{1,k}^{\top})\boldsymbol{T}_{x}(\alpha_{k})\boldsymbol{y}_{t-l};r_{m-1}p_{\alpha^{k}_{(m)}},p^{(m)})$, and $p^{(m)}:=p_{\alpha^{k}_{(m+1)}}\cdots p_{\alpha^{k}_{(M)}}$.
+To update $\boldsymbol{H}_{n,k^{\prime}}$ in $\boldsymbol{\Lambda}_{y}$, we define $\boldsymbol{g}_{t}=\boldsymbol{\Theta}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top})\boldsymbol{x}_{t}$, and thus the model becomes
 […]
-An identical procedure is then applied to $\mathcal{P}_{y}$, ensuring both $\boldsymbol{\Lambda}_{x}$ and $\boldsymbol{\Lambda}_{y}$ are represented in their most parsimonious forms.
-The final output of the rank reduction step is the optimal pair $(\mathcal{P}_{y},\mathcal{P}_{x})$ corresponding to the most efficient, yet equivalent, representation of the model structure.
-We summarize the overall hyperparameter selection procedure into Algorithm 2.
-5 Simulation studies
-This section conducts three simulation studies to evaluate the finite-sample performance of our OLS estimator proposed in Algorithm 1, investigate the effects of action order misspecification, as discussed in Proposition 1, and compare the performance of different feature extraction methods for hierarchical data under the framework of supervised factor modeling.
-In the first experiment, the data generating process follows (2.12) with $L=2$ and $N=3$ and equal dimension sizes, i.e., $\mbox{$\mathscr{A}$}_{\mathrm{AR}}]_{3}=([\mbox{$\mathscr{A}$}_{1}]_{3},[\mbox{$\mathscr{A}$}_{2}]_{3})$ and $\mbox{$\mathscr{Y}$}_{t}\in\mathbb{R}^{q\times q\times q}$.
-For simplicity, let intrinsic ranks $r_{n,1}=r_{n,2}=r$, $s_{n,1}=s_{n,2}=s$, where $r=s+1$.
-We consider three types of sub-Gaussian noise: (i) i.i.d.
-uniform noise, (ii) i.i.d.
-standard normal noise, and (iii) correlated gaussian noise.
-The detailed data generation process is relegated to Supplementary Material.
-From Theorem 2, the estimation error converges at a rate of $\sqrt{d_{\mathrm{AR}}/T}$, where $d_{\mathrm{AR}}=2(K_{x}r^{2}+K_{y}s^{2})Nq+2LK_{x}K_{y}rs$.
-To numerically verify this convergence rate, we consider three settings: (a) Fix $(r,T)=(3,2500)$ and vary $q\in\{10,11,12,13,14\}$; (b) Fix $(q,T)=(10,2500)$ and vary $r\in\{3,4,5,6,7\}$; (c) Fix $(q,r)=(10,3)$ and vary $T\in\{833,1000,1250,1670,2500\}$.
-In all cases, we apply Algorithm 1 to estimate $\mbox{$\mathscr{A}$}_{\mathrm{AR}}$.
-Figure 2 presents the estimation error $\|\mbox{$\mathscr{\widehat{A}}$}_{\mathrm{LR}}-\mbox{$\mathscr{A}$}^{*}_{\mathrm{LR}}\|_{\mathrm{F}}$ for these settings, averaged over 200 replications.
-The observed linearity in Figure 2 confirms the theoretical convergence rate in Theorem 2.
-It verifies that the estimation error depends on $q$ linearly, and $r$ quadratically, thereby demonstrating the efficiency of our proposed method.
 […]
-In the second experiment, we investigate the impact of action order misspecification, as discussed in Proposition 1.
-We set $M=3$, then the action set is $\mathcal{P}(3)$, which consists of all permutations of $\{1,2,3\}$ with cardinality $|\mathcal{P}(3)|=6$.
-The data are generated by $\boldsymbol{y}_{t}=\mathcal{M}(\mathcal{T}(\mbox{$\mathscr{X}$}_{t};\alpha_{1});\{\boldsymbol{G}_{m,1}\}_{m=1}^{3})$, where the true action order is $\alpha_{1}=123$.
-For simplicity, let $r_{1,k}=r_{2,k}=r_{3,k}=r_{k}$ for $k=1,\ldots,6$, and $\boldsymbol{G}_{m,1}$’s are generated similarly as in the first experiment.
-We consider the i.i.d.
-normal setting for $\mbox{$\mathscr{X}$}_{t}\in\mathbb{R}^{r_{k}\times r_{k}\times r_{k}}$ with $\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})\sim N(\boldsymbol{0},\boldsymbol{C})$, where $\boldsymbol{C}$ is a random matrix.
-By Proposition 1, for each $\alpha_{k}\in\mathcal{P}(3)$, there exists a corresponding set of $\{\boldsymbol{G}_{m,k}\}_{m=1}^{3}$ such that $\mathcal{M}(\mathcal{T}(\mbox{$\mathscr{X}$}_{t};\alpha_{1});\{\boldsymbol{G}_{m,1}\}_{m=1}^{3})=\mathcal{M}(\mathcal{T}(\mbox{$\mathscr{X}$}_{t};\alpha_{k});\{\boldsymbol{G}_{m,k}\}_{m=1}^{3})$, with the cost of potentially larger intrinsic rank $r_{k}$.
-We apply Algorithm 1 to estimate $\boldsymbol{G}_{m,k}$’s for each $k$, and compute the mean squared error (MSE) of the predictions.
-Figure 3(a) presents the MSEs for different values of $r_{k}$’s averaged over 50 replications.
-For both settings of $\mbox{$\mathscr{X}$}_{t}$, we observe that when $r_{k}$ is small, the MSEs for the misspecified action orders are significantly larger than that of the true action order $\alpha_{1}$.
-However, as $r_{k}$ increases, MSEs for all action orders converge to the same level as that of the true action order $\alpha_{1}$.
-This aligns with our theoretical findings in Proposition 1.
-Figure 3: (a) Average MSEs under different action orders with varying intrinsic rank $r_{k}$.
-(b) Average rolling forecast MSEs for HOFM, HFM, and our proposed method under supervised factor modeling framework.
-In the last experiment, we compare our proposed method with two hierarchical feature extraction methods: HFM (schmid1957development) and HOFM (yung1999relationship), under the framework of SFM.
-The data generating process follows (2.12) with the coefficient tensor $\mbox{$\mathscr{A}$}_{\mathrm{AR}}$ constructed by both HOFM and our proposed method.
-Details of the data generation and training procedure are relegated to Supplementary Material.
-To compare the one-step ahead forecast performance of HOFM, HFM, and our proposed method, we apply Algorithm 1 to estimate the coefficient tensor $\mbox{$\mathscr{A}$}_{\mathrm{AR}}$ for our model, and use the alternating least squares algorithm to estimate the coefficient tensor for HFM and HOFM.
-Figure 3(b) presents the average rolling forecast MSEs over 50 replications.
-We observe that our method consistently outperforms both HFM and HOFM, regardless of whether the data are generated by HOFM or our method.
-This demonstrating the efficiency of our model in capturing hierarchical structures with shared components within each mode in loading matrices for feature extractions.
-6 Real data analysis
-In this section, we demonstrate the application of our proposed model to the International Personality Item Pool (IPIP) NEO-120 personality inventory (johnson2014measuring), using data collected from June 1, 2015 to April 30, 2025.
-The dataset is available in the Hugging Face Datasets repository: https://huggingface.co/datasets/ecorbari/IPIP120-SCORES.
-It includes scores for the five major personality traits (Neuroticism, Extraversion, Conscientiousness, Agreeableness, and Openness to Experience), assessed by 24 questions.
-Participants are categorized by gender (two groups) and age (six groups: 10-19, 20-29, 30-39, 40-49, 50-59, and 60+), resulting in 12 demographic groups.
-To analyze trends in personality traits over time, we compute monthly average scores for each question (recorded on a 1-5 Likert scale) within each demographic group, forming a tensor time series $\{\mbox{$\mathscr{Y}$}_{t}\in\mathbb{R}^{12\times 5\times 24}\}_{t=1}^{T}$ with $T=115$ observations.
-There is one missing entry in the dataset, which we impute using spline interpolation.
-For preprocessing, the data are first detrended by first-order differencing, and then standardized to have zero means.
-The tensor autoregression model (2.12) is employed with each coefficient tensor $\mbox{$\mathscr{A}$}_{l}\in\mathbb{R}^{12\times 5\times 24\times 12\times 5\times 24}$.
-Observations are divided into a training set with the first 103 months and a testing set with the last 12 months.
-To enable a feasible estimation, we consider our proposed model, SFM based on CP and TT decompositions, and HOFM under SFM frameworks for feature extraction with the lag order fixed at $L=1$ by jointly consider the PACFs of all entries in data.
-We omit the comparison with Tucker decomposition and HFM based on SFM frameworks due to thier large number of parameters in the coefficient matrix, which makes them infeasible for this data analysis.
-For our proposed model, from prior knowledge of the data, we set the candidate action order sets as $\mathcal{P}_{x}^{(0)}=\mathcal{P}_{y}^{(0)}=\{(3,2,1),(1,3,2)\}$, i.e., the features are extracted by the order of questions, personalities, and demographic groups, or the order of demographic groups, questions, and personalities.
-We then apply Algorithm 2 to get the selected intrinsic ranks $\boldsymbol{r}_{1}=(1,1,2)$, $\boldsymbol{r}_{2}=(9,1,2)$, $\boldsymbol{s}_{1}=(3,1,2)$, and $\boldsymbol{s}_{2}=(3,2,2)$.
-For the two tensor decomposition based SFMs, we adopt the ALS methods in zhou2013tensor for CP-based tensor autoregression, and wang2022high for Tucker-based tensor autoregression.
-For TT-based SFM, we follow si2024efficient and qin2025computational to perform Riemannian gradient descent for estimation.
-The intrinsic ranks for these methods are selected by their respective methods.
-We next evaluate the rolling forecast performance on the testing set of these four methods, i.e., fit the model to historical data up to month $t$ to forecast the data at month $t+1$ for $t=103,\ldots,114$.
-Specifically, we use mean squared forecast errors (MSFE) and mean absolute forecast errors (MAFE) to assess the one-step ahead forecast accuracy.
-Table 2 summarizes the results, where our proposed method outperforms all competing methods in both MSFE and MAFE metrics.
-This result supports the existence of hierarchical structures in the IPIP NEO-120 data, which can be effectively captured by our proposed model.
-Moreover, the superior performance of our method over TT decomposition based SFM highlights the advantage of our model by considering multiple action orders for general hierarchical data.
-And the better performance compared to HOFM based SFM demonstrates the benefit of improved efficiency in feature extraction by sharing components within each mode in loading matrices.
-In conclusion, our proposed model provides an efficient and interpretable approach for modeling tensor time series data with hierarchical structures.
-Table 2: Mean squared forecast errors (MSFE) and mean absolute forecast errors (MAFE) of our model and three competing methods on the testing set of the IPIP NEO-120 data.
-The smallest numbers in each row are in bold.
-Ours CP $T$ HOFM
-MSFE 138.78 142.28 145.73 142.15
-MSAE 279.28 280.63 283.43 280.58
-Finally, we zoom in on the estimated coefficient tensor from our proposed model to interpret the hierarchical structures in the IPIP NEO-120 data and the importance of different action orders.
-We discuss the component matrices involved in feature extraction for responses under action order $(1,3,2)$ as an example.
-Figure 4(a) presents the heatmap of the component matrix $\boldsymbol{H}_{1,1}^{\top}$, which correspond to the first step of feature extraction, summarizing the information across the 24 questions for each personality trait.
-In addition, Figure 4(b) displays $\boldsymbol{H}_{2,1}^{\top}$ that represents the combined effects of personality traits and questions by a single extracted feature.
-Lastly, Figure 4(c) illustrates $\boldsymbol{H}_{3,1}^{\top}$, capturing the interactions between demographic groups and the combined features from questions and personality traits.
-The response tensor $\mbox{$\mathscr{Y}$}_{t}$ is then summarized into two factors as its low-dimensional representation.
-Component matrices with the second action order $(1,3,2)$ for responses and those for predictors with both action orders can be found in Supplementary Material, which have similar interpretations as those for predictors.
-Finally, Figure 4(d) presents the heatmap of $\boldsymbol{\Theta}$, the estimated coefficient matrix that captures the relationships between the extracted factors from predictors and responses by two action orders.
-We observe that features extracted by action order $(3,2,1)$ have a more significant impact on extracted features from responses under both action orders compared to those extracted by action order $(1,3,2)$, as indicated by the larger absolute values in the corresponding blocks of $\boldsymbol{\Theta}$.
-Figure 4: Heatmaps of (a) $\boldsymbol{H}_{1,1}^{\top}$, (b) $\boldsymbol{H}_{2,1}^{\top}$, (c) $\boldsymbol{H}_{3,1}^{\top}$, and (d) $\boldsymbol{\Theta}$.
-7 Conclusion and discussion
-In this paper, we introduce a novel autoregressive model under SFM framework for tensor time series data that effectively captures hierarchical structures through the incorporation of multiple action orders in the construction of the coefficient tensor.
-We employ a structure of shared components within each mode in loading matrices for efficient feature extraction.
-An alternating least squares algorithm is developed for parameter estimation, and a boosting-based method is proposed for hyperparameter selection, which aims to reveal the underlying hierarchical structures present in the data.
-Theoretical analyses demonstrate the consistency of our estimator under sub-Gaussian noise with misspecification of action orders.
-Simulation studies and real data analysis illustrate the superior performance of our method compared to existing approaches.
-For future research, we may extend the proposed model to accommodate more complex time series structures, such as moving average (MA) or ARMA models.
-We can also develop robust estimation techniques for handling heavy-tailed noise distributions, so that our methods can be deployed to a wider range of real-world applications.
-Additionally, investigating statistical inference procedures for the selected action orders and intrinsic ranks would provide valuable insights into the interpretability and reliability of the model in practical settings.
+$\boldsymbol{y}_{t,k^{\prime}}=\boldsymbol{H}_{n,k^{\prime}}^{\mathrm{fore}}(\boldsymbol{Y}_{t,n,k^{\prime}}^{\top}\otimes\boldsymbol{I}_{q_{\beta^{k^{\prime}}_{(n)}}s_{n-1,k^{\prime}}})\operatorname*{vec}(\boldsymbol{H}_{n,k^{\prime}})+\boldsymbol{e}_{t},$
+[…]
+For the coefficient matrix $\boldsymbol{\Theta}$, it is a bilinear regression problem equivalent to solving
+$\boldsymbol{y}_{t}=\big[(\boldsymbol{x}_{t}^{\top}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}))\otimes\boldsymbol{\Lambda}_{y}\big]\operatorname*{vec}(\boldsymbol{\Theta})+\boldsymbol{e}_{t}.$
+Each of these OLS problems can then be solved efficiently with closed forms.
+To satisfy the orthogonality assumptions required by the model, an additional orthogonalization step is therefore necessary.
+Specifically, for any set of invertible matrices $\boldsymbol{O}_{m,k}\in\mathbb{R}^{r_{m,k}\times r_{m,k}}$, where $1\leq m\leq M$, the following identity holds with proof provided in Section B.3:
+$\begin{split}\boldsymbol{G}_{M,k}^{\top}(\boldsymbol{I}_{p_{\alpha^{k}_{(M)}}}\otimes\boldsymbol{G}_{M-1,k}^{\top})&\cdots(\boldsymbol{I}_{p_{\alpha^{k}_{(2)}}\cdots p_{\alpha^{k}_{(M)}}}\otimes\boldsymbol{G}_{1,k}^{\top})=\\ &(\boldsymbol{O}_{M,k}^{\top})^{-1}\widetilde{\boldsymbol{G}}_{M,k}^{\top}(\boldsymbol{I}_{p_{\alpha^{k}_{(M)}}}\otimes\widetilde{\boldsymbol{G}}_{M-1,k}^{\top})\cdots(\boldsymbol{I}_{p_{\alpha^{k}_{(2)}}\cdots p_{\alpha^{k}_{(M)}}}\otimes\widetilde{\boldsymbol{G}}_{1,k}^{\top})\end{split}$ (B.12)
+where the transformed components $\widetilde{\boldsymbol{G}}_{m,k}$ are defined recursively as $\widetilde{\boldsymbol{G}}_{1,k}=\boldsymbol{G}_{1,k}\boldsymbol{O}_{1,k}$, $\widetilde{\boldsymbol{G}}_{2,k}=(\boldsymbol{I}_{p_{\alpha^{k}_{(2)}}}\otimes\boldsymbol{O}_{1,k}^{-1})\boldsymbol{G}_{2,k}\boldsymbol{O}_{2,k}$, $\ldots$, $\widetilde{\boldsymbol{G}}_{M,k}=(\boldsymbol{I}_{p_{\alpha^{k}_{(M)}}}\otimes\boldsymbol{O}_{M-1,k}^{-1})\boldsymbol{G}_{M,k}\boldsymbol{O}_{M,k}$.
+By appropriately selecting the matrices $\boldsymbol{O}_{m,k}$ for $1\leq m\leq M$, we can enforce orthogonality across all components $\{\boldsymbol{G}_{m,k}\}_{m=1}^{M}$.
+We apply this orthogonalization procedure to all block columns of $\boldsymbol{\Lambda}_{x}$ and $\boldsymbol{\Lambda}_{y}$, and denote the whole process by the operator $\mathrm{SSVD}(\cdot)$, detailed in Section B.4, which is applied after the updates of all $\{\{\boldsymbol{H}_{n,k^{\prime}}\}_{n=1}^{N}\}_{k^{\prime}=1}^{K_{y}}$ and $\{\{\boldsymbol{G}_{m,k}\}_{m=1}^{M}\}_{k=1}^{K_{x}}$ have been completed.
+B.3 Proof of (B.12)
+We begin by verifying the identity for the last two terms on the left-hand side of (B.12).
+Let $\boldsymbol{O}_{1,k}\in\mathbb{R}^{r_{1,k}\times r_{1,k}}$ be an arbitrary invertible matrix.
+Consider the following product:
+$(\boldsymbol{I}_{p_{\alpha^{k}_{(3)}}\cdots p_{\alpha^{k}_{(M)}}}\otimes\boldsymbol{G}_{2,k}^{\top})(\boldsymbol{I}_{p_{\alpha^{k}_{(2)}}p_{\alpha^{k}_{(3)}}\cdots p_{\alpha^{k}_{(M)}}}\otimes\boldsymbol{G}_{1,k}^{\top})$
+$=$ $\penalty 10000\ \boldsymbol{I}_{p_{\alpha^{k}_{(3)}}\cdots p_{\alpha^{k}_{(M)}}}\otimes[\boldsymbol{G}_{2,k}^{\top}(\boldsymbol{I}_{p_{\alpha^{k}_{(2)}}}\otimes\boldsymbol{G}_{1,k}^{\top})]$
+$=$ $\penalty 10000\ \boldsymbol{I}_{p_{\alpha^{k}_{(3)}}\cdots p_{\alpha^{k}_{(M)}}}\otimes[\boldsymbol{G}_{2,k}^{\top}(\boldsymbol{I}_{p_{\alpha^{k}_{(2)}}}\otimes((\boldsymbol{O}_{1,k}^{-1})^{\top}\boldsymbol{O}_{1,k}^{\top}\boldsymbol{G}_{1,k}^{\top}))]$
+$=$ $\penalty 10000\ \boldsymbol{I}_{p_{\alpha^{k}_{(3)}}\cdots p_{\alpha^{k}_{(M)}}}\otimes[\boldsymbol{G}_{2,k}^{\top}(\boldsymbol{I}_{p_{\alpha^{k}_{(2)}}}\otimes\boldsymbol{O}_{1,k}^{-1})^{\top}(\boldsymbol{I}_{p_{\alpha^{k}_{(2)}}}\otimes(\boldsymbol{O}_{1,k}^{\top}\boldsymbol{G}_{1,k}^{\top}))]$
+$=$ $\penalty 10000\ (\boldsymbol{I}_{p_{\alpha^{k}_{(3)}}\cdots p_{\alpha^{k}_{(M)}}}\otimes[(\boldsymbol{I}_{p_{\alpha^{k}_{(2)}}}\otimes\boldsymbol{O}_{1,k}^{-1})\boldsymbol{G}_{2,k}]^{\top})(\boldsymbol{I}_{p_{\alpha^{k}_{(2)}}p_{\alpha^{k}_{(3)}}\cdots p_{\alpha^{k}_{(M)}}}\otimes(\boldsymbol{G}_{1,k}\boldsymbol{O}_{1,k})^{\top}).$
+This computation shows that the matrix $\boldsymbol{O}_{1,k}$ can be absorbed into $\boldsymbol{G}_{1,k}$ and $\boldsymbol{G}_{2,k}$ through the following change of variables:
+$\widetilde{\boldsymbol{G}}_{1,k}=\boldsymbol{G}_{1,k}\boldsymbol{O}_{1,k},\qquad\bar{\boldsymbol{G}}_{2,k}=(\boldsymbol{I}_{p_{\alpha^{k}_{(2)}}}\otimes\boldsymbol{O}_{1,k}^{-1})\boldsymbol{G}_{2,k},$ (B.13)
+where $\bar{\boldsymbol{G}}_{2,k}$ is an intermediate form of the second component. In the subsequent steps, we apply a similar change of variables recursively:
+$\widetilde{\boldsymbol{G}}_{m,k}=\bar{\boldsymbol{G}}_{m,k}\boldsymbol{O}_{m,k},\qquad\bar{\boldsymbol{G}}_{m+1,k}=(\boldsymbol{I}_{p_{\alpha^{k}_{(m+1)}}}\otimes\boldsymbol{O}_{m,k}^{-1})\boldsymbol{G}_{m+1,k},$ (B.14)
+for $2\leq m\leq M-1$, and finally define $\widetilde{\boldsymbol{G}}_{M,k}=\bar{\boldsymbol{G}}_{M,k}\boldsymbol{O}_{M,k}$.
+By successively applying these transformations, each $\boldsymbol{G}_{m,k}$ is mapped to $\widetilde{\boldsymbol{G}}_{m,k}$ through a sequence of invertible matrix multiplications involving Kronecker products with identity matrices.
+This construction guarantees that the full product on the left-hand side of (B.12) is equal to the transformed product on the right-hand side, up to a scaling by the inverse transpose of the final matrix $\boldsymbol{O}_{M,k}^{\top}$.
+This completes the proof.
+B.4 Algorithm for SSVD
+In this subsection, we propose an algorithm to orthogonalize all the components in $\boldsymbol{\Lambda}_{x}$ and $\boldsymbol{\Lambda}_{y}$ required in Section 4.1 of the manuscript.
+We apply (B.12) to orthogonalize the components $\{\{\boldsymbol{G}_{m,k}\}_{m=1}^{M}\}_{k=1}^{K_{x}}$, and the specific update rule of each orthogonal component follows (B.13) or (B.14).
+The matrices concerned, i.e., $\widetilde{\boldsymbol{G}}_{m,k}$’s and $\boldsymbol{O}_{m,k}$’s, are obtained via SVD.
+After obtaining the orthogonalized components, we turn to the product $\boldsymbol{\Theta}_{l}\boldsymbol{\Lambda}_{x}^{\top}$ ($l=1,2,\ldots,L$), which can be written as
+$\begin{bmatrix}\boldsymbol{\Theta}_{l,1}&\cdots&\boldsymbol{\Theta}_{l,K_{x}}\end{bmatrix}\begin{bmatrix}(\boldsymbol{O}_{M,1}^{\top})^{-1}\boldsymbol{\Lambda}_{x,1}^{\top}\\ \vdots\\ (\boldsymbol{O}_{M,K_{x}}^{\top})^{-1}\boldsymbol{\Lambda}_{x,K_{x}}^{\top}\end{bmatrix}=\begin{bmatrix}\boldsymbol{\Theta}_{l,1}(\boldsymbol{O}_{M,1}^{\top})^{-1}&\cdots&\boldsymbol{\Theta}_{l,K_{x}}(\boldsymbol{O}_{M,K_{x}}^{\top})^{-1}\end{bmatrix}\begin{bmatrix}\boldsymbol{\Lambda}_{x,1}^{\top}\\ \vdots\\ \boldsymbol{\Lambda}_{x,K_{x}}^{\top}\end{bmatrix}$.
+(B.15)
+This shows that the effect of the scaling matrices $\{(\boldsymbol{O}_{M,k}^{\top})^{-1}\}_{k=1}^{K_{x}}$ can be absorbed into the coefficient matrices $\boldsymbol{\Theta}_{l,k}$ for each lag $l$ and action $k$.
+Therefore, we can enforce orthogonality on all components $\{\boldsymbol{G}_{m,k}\}_{m=1}^{M}$ of $\boldsymbol{\Lambda}_{x}$ without affecting the model structure.
+A similar procedure can be applied to the components $\{\{\boldsymbol{H}_{n,k^{\prime}}\}_{n=1}^{N}\}_{k^{\prime}=1}^{K_{y}}$ of $\boldsymbol{\Lambda}_{y}$ to achieve orthogonality, and by viewing the updated matrix $\boldsymbol{\Theta}$ as a whole and transposing $\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}$, (B.15) can be used analogously to absorb the scaling matrices $\{(\boldsymbol{O}_{N,k^{\prime}}^{\top})^{-1}\}_{k^{\prime}=1}^{K_{y}}$ into $\boldsymbol{\Theta}$.
+We summarize this sequential orthogonalization process into the following algorithm, denoted by $\text{SSVD}(\cdot)$.
+Algorithm S.1 SSVD procedure for orthogonalizing core components
+1:Input: $\{\{\boldsymbol{G}_{m,k}\}_{m=1}^{M}\}_{k=1}^{K_{x}},\{\{\boldsymbol{H}_{n,k^{\prime}}\}_{n=1}^{N}\}_{k^{\prime}=1}^{K_{y}}$ and $\boldsymbol{\Theta}=(\boldsymbol{\Theta}_{1},\dots,\boldsymbol{\Theta}_{L})$.
+2:for $k=1$ to $K_{x}$ do
+3:  for $m=1$ to $M$ do
+4:   Set $\bar{\boldsymbol{G}}_{m,k}=(\boldsymbol{I}_{q_{\alpha^{k}_{(m)}}}\otimes\boldsymbol{O}_{G,m-1,k}^{-1})\boldsymbol{G}_{m,k}$ if $m>1$; else $\bar{\boldsymbol{G}}_{m,k}=\boldsymbol{G}_{m,k}$.
+5:   Conduct SVD on $\bar{\boldsymbol{G}}_{m,k}$ as $\bar{\boldsymbol{G}}_{m,k}=\boldsymbol{U}_{G,m,k}\boldsymbol{\Sigma}_{G,m,k}\boldsymbol{V}^{\top}_{G,m,k}$.
+6:   Set $\widetilde{\boldsymbol{G}}_{m,k}=\boldsymbol{U}_{G,m,k}$ and $\boldsymbol{O}_{G,m,k}=(\boldsymbol{\Sigma}_{G,m,k}\boldsymbol{V}^{\top}_{G,m,k})^{-1}$.
+7:  end for
+8:  Update $\boldsymbol{\Theta}^{(1)}_{l,k}=\boldsymbol{\Theta}_{l,k}(\boldsymbol{O}_{G,M,k}^{\top})^{-1}$ for each $l\in\{1,2,\dots,L\}$ with the form of (B.15).
+9:end for
+10:for $k^{\prime}=1$ to $K_{y}$ do
+11:  Repeat line 37 for $\{\{\boldsymbol{H}_{n,k^{\prime}}\}_{n=1}^{N}\}_{k^{\prime}=1}^{K_{y}}$ to get $\{\{\widetilde{\boldsymbol{H}}_{n,k^{\prime}}\}_{n=1}^{N}\}_{k^{\prime}=1}^{K_{y}}$ and $\boldsymbol{O}_{H,N,k^{\prime}}$.
+12:  Update $(\boldsymbol{\Theta}^{(2)})^{\top}]_{k^{\prime}}=[(\boldsymbol{\Theta}^{(1)})^{\top}]_{k^{\prime}}(\boldsymbol{O}_{H,N,k^{\prime}}^{\top})^{-1}$ similar to line 8.
+13:end for
+14:Return $\{\{\widetilde{\boldsymbol{G}}_{m,k}\}_{m=1}^{M}\}_{k=1}^{K_{x}}$, $\{\{\widetilde{\boldsymbol{H}}_{n,k^{\prime}}\}_{n=1}^{N}\}_{k^{\prime}=1}^{K_{y}}$, and $\boldsymbol{\Theta}^{(2)}$.
+Remark S.1.
+In the unlikely event that any $\bar{\boldsymbol{G}}_{m,k}$ or $\bar{\boldsymbol{H}}_{n,k^{\prime}}$ is rank-deficient (take $\bar{\boldsymbol{G}}_{m,k}$ for example and omit $k$ for convenience), we pad zero columns to the corresponding $\boldsymbol{U}_{m}$ to ensure that the output component $\widetilde{\boldsymbol{G}}_{m}$ has the same dimensions as the input components $\boldsymbol{G}_{m}$, and we add rows to $\boldsymbol{O}_{G,m}$ to make it square and invertible.
+Line 6 can be so modified if necessary, and the successive procedure of Section B.3 still applies.
+To remove the zero columns in $\widetilde{\boldsymbol{G}}_{m}$, we notice that $(\boldsymbol{I}_{p_{\alpha_{(m+2)}}\cdots p_{\alpha_{(M)}}}\otimes\widetilde{\boldsymbol{G}}_{m+1}^{\top})(\boldsymbol{I}_{p_{\alpha_{(m+1)}}p_{\alpha_{(m+2)}}\cdots p_{\alpha_{(M)}}}\otimes\widetilde{\boldsymbol{G}}_{m}^{\top})=\boldsymbol{I}_{p_{\alpha_{(m+2)}}\cdots p_{\alpha_{(M)}}}\otimes[\widetilde{\boldsymbol{G}}_{m+1}^{\top}(\boldsymbol{I}_{p_{\alpha_{(m+1)}}}\otimes\widetilde{\boldsymbol{G}}_{m}^{\top})]$, and we remove the zero rows in $\boldsymbol{I}_{p_{\alpha_{(m+1)}}}\otimes\widetilde{\boldsymbol{G}}_{m}^{\top}$ and the corresponding columns in $\widetilde{\boldsymbol{G}}_{m+1}^{\top}$.
+Finally, we re-run Algorithm S.1 to restore the orthogonality of all components.
+B.5 Proof of Proposition 2
+We aim to show that there exist $\boldsymbol{r}_{k,3}=\boldsymbol{r}_{k,1}+\boldsymbol{r}_{k,2}$ (in terms of elementwise ordering) and a matrix $\boldsymbol{O}$ such that $\widetilde{\boldsymbol{\Lambda}}_{x}(\alpha_{k},\boldsymbol{r}_{k,3})=\boldsymbol{O}\left[\widetilde{\boldsymbol{\Lambda}}_{x}(\alpha_{k},\boldsymbol{r}_{k,1}),\widetilde{\boldsymbol{\Lambda}}_{x}(\alpha_{k},\boldsymbol{r}_{k,2})\right]$, that is, absent orthogonality constraints, identical action orders can be merged with additive $\boldsymbol{r}_{k}$’s.
+The proof is constructive.
+We assume without loss of generality that $\alpha_{k}=(1,2,\cdots,M)$ and omit the subscript $k$.
+Let $\boldsymbol{G}_{m,1}\in\mathbb{O}^{p_{m}r_{1,m-1}\times r_{1,m}}$ and $\boldsymbol{G}_{m,2}\in\mathbb{O}^{p_{m}r_{2,m-1}\times r_{2,m}}$ be the respective core matrices corresponding to the reversed construction of $\widetilde{\boldsymbol{\Lambda}}_{x}(\alpha,\boldsymbol{r}_{1})$ and $\widetilde{\boldsymbol{\Lambda}}_{x}(\alpha,\boldsymbol{r}_{2})$, i.e., the left hand side of (B.17).
+We construct the new core matrices $\boldsymbol{G}_{m}^{\top}=\left[\begin{smallmatrix}\boldsymbol{G}_{m,1}^{\top}&\\ &\boldsymbol{G}_{m,2}^{\top}\\ \end{smallmatrix}\right]\in\mathbb{R}^{(r_{1,m}+r_{2,m})\times p_{m}(r_{1,m-1}+r_{2,m-1})}$ for $m=2,\ldots,M$, and $\boldsymbol{G}_{1}^{\top}=\left[\begin{smallmatrix}\boldsymbol{G}_{1,1}^{\top}\\ \boldsymbol{G}_{1,2}^{\top}\\ \end{smallmatrix}\right]\in\mathbb{R}^{(r_{1,1}+r_{2,1})\times p_{1}}$, and the preliminary row block
+$\widetilde{\boldsymbol{\Lambda}}_{x}^{\top}(\alpha,\boldsymbol{r}_{1}+\boldsymbol{r}_{2})=\begin{bmatrix}\boldsymbol{G}_{M,1}^{\top}&\\ &\boldsymbol{G}_{M,2}^{\top}\\ \end{bmatrix}$ $\left(\begin{bmatrix}\boldsymbol{G}_{M-1,1}^{\top}&\\ &\boldsymbol{G}_{M-1,2}^{\top}\\ \end{bmatrix}\otimes\boldsymbol{I}_{p_{M}}\right)\cdots$
+$\left(\begin{bmatrix}\boldsymbol{G}_{2,1}^{\top}&\\ &\boldsymbol{G}_{2,2}^{\top}\\ \end{bmatrix}\otimes\boldsymbol{I}_{p_{3}\cdots p_{M}}\right)\left(\begin{bmatrix}\boldsymbol{G}_{1,1}^{\top}\\ \boldsymbol{G}_{1,2}^{\top}\\ \end{bmatrix}\otimes\boldsymbol{I}_{p_{2}\cdots p_{M}}\right)\boldsymbol{T}(\alpha).$
+The equivalence can be verified by expanding the above expression by block matrix multiplication rules.
+Then, we convert the reversed form back to the standard form, i.e., the right hand side of (B.17), according to Section B.6.
+The matrix $\boldsymbol{O}$ shall be chosen according to Section B.4 to ensure orthogonality of the new core matrices.
+B.6 Proof of Remark 2
+Remark 2 states that the proposed model generalizes the Tensor Train (TT) decomposition when $K_{y}=K_{x}=1$ with other appropriate specifications.
+Here, we provide a detailed proof of this claim by demonstrating how the TT decomposition can be represented within our model framework.
+We begin by recalling the representation of TT decomposition.
+For a $(N+M)$-th order tensor $\mbox{$\mathscr{A}$}\in\mathbb{R}^{q_{1}\times\cdots\times q_{N}\times p_{1}\times\cdots\times p_{M}}$, its entry can be expressed in the TT format as
+$\mbox{$\mathscr{A}$}_{j_{1}\ldots j_{N}i_{1}\ldots i_{M}}=\boldsymbol{H}_{1}(j_{1})\cdots\boldsymbol{H}_{N}(j_{N})\boldsymbol{G}_{1}(i_{1})\cdots\boldsymbol{G}_{M}(i_{M}),$
+[…]
+According to Lemma 3.1 in Zhou et al.
+(2022), Proposition 1 in Si et al.
+(2024), and Qin and Zhu (2025), the $N$-th sequential matricization of $\mathscr{A}$ with TT ranks $(s_{1},\ldots,s_{N},r_{2},\ldots,r_{M})$, admits the following decomposition:
+[…]
+[…]
+To connect this decomposition with our model, consider the case where $K_{y}=K_{x}=1$, and $\mathcal{P}_{y}(N,K_{y})=\{(1,2,\ldots,N)\}$, $\mathcal{P}_{x}(M,K_{x})=\{(M,M-1,\ldots,1)\}$. We claim that
+[…]
+Consequently, $\mbox{$\mathscr{A}$}\in\mathcal{S}_{\mathrm{LR}}(\{(r_{M},\ldots,r_{1})\},\{(s_{1},\ldots,s_{N})\},\mathcal{P}_{x}(M,K_{x}),\mathcal{P}_{y}(N,K_{y}))$, which corresponds to a special case of our proposed regression model, where $\boldsymbol{\Theta}_{\mathrm{TT}}$ is diagonal.
+It remains to verify (B.16).
+We only need to show $\left([\mbox{$\mathscr{G}$}_{1}]_{1}([\mbox{$\mathscr{G}$}_{2}]_{1}\otimes\boldsymbol{I}_{p_{1}})\cdots(\boldsymbol{G}_{M}^{\top}\otimes\boldsymbol{I}_{p_{1}\cdots p_{M-1}})\right)^{\top}\in\mathcal{R}_{x}(\{(r_{M},\ldots,r_{1})\},\mathcal{P}_{x}(M,K_{x}))$, and by changing the notations and indexing to conform with our settings, the target of proof becomes that, for all $\{\boldsymbol{G}_{m}\in\mathbb{O}^{p_{m}r_{m-1}\times r_{m}}\}_{m=1}^{M}$ ($r_{0}\vcentcolon=1$), there exist $\{\boldsymbol{H}_{m}\in\mathbb{O}^{p_{m}r_{m-1}\times r_{m}}\}_{m=1}^{M}$ and a permutation matrix $\boldsymbol{T}_{M}$ such that for any $\boldsymbol{x}_{M}\in\mathbb{R}^{p_{1}p_{2}\cdots p_{M}}$,
+$\begin{split}&\boldsymbol{G}_{M}^{\top}(\boldsymbol{G}_{M-1}^{\top}\otimes\boldsymbol{I}_{p_{M}})\cdots(\boldsymbol{G}_{2}^{\top}\otimes\boldsymbol{I}_{p_{3}\cdots p_{M}})(\boldsymbol{G}_{1}^{\top}\otimes\boldsymbol{I}_{p_{2}\cdots p_{M}})\boldsymbol{x}_{M}\\ =\penalty 10000\ &\boldsymbol{H}_{M}^{\top}(\boldsymbol{I}_{p_{M}}\otimes\boldsymbol{H}_{M-1}^{\top})\cdots(\boldsymbol{I}_{p_{3}\cdots p_{M}}\otimes\boldsymbol{H}_{2}^{\top})(\boldsymbol{I}_{p_{2}\cdots p_{M}}\otimes\boldsymbol{H}_{1}^{\top})\boldsymbol{T}_{M}\boldsymbol{x}_{M}.\end{split}$ (B.17)
+[…]
+$\boldsymbol{H}_{m}^{\top}(i,(j-1)r_{m-1}+k)=\boldsymbol{G}_{m}^{\top}(i,j+(k-1)p_{m}).$ (B.18)
+Specifically, $\boldsymbol{H}_{1}=\boldsymbol{G}_{1}$.
+Note that this construction preserves orthogonality, i.e., $\boldsymbol{H}_{m}\in\mathbb{O}^{p_{m}r_{m-1}\times r_{m}}$ for all $m$, because only the ordering of the rows of $\boldsymbol{G}_{m}$ is changed to form $\boldsymbol{H}_{m}$.
+Let $\mbox{$\mathscr{X}$}_{m}\in\mathbb{R}^{p_{m}\times p_{m-1}\times\cdots\times p_{1}}$ ($m\leq M$) satisfy $\operatorname*{vec}(\mbox{$\mathscr{X}$}_{M})=\boldsymbol{x}_{M}$, and $\widetilde{\mbox{$\mathscr{X}$}}_{m}\in\mathbb{R}^{p_{1}\times p_{2}\times\cdots\times p_{m}}$ be the tensor satisfying $\widetilde{\mbox{$\mathscr{X}$}}_{m}(i_{1},i_{2},\ldots,i_{m})=\mbox{$\mathscr{X}$}_{m}(i_{m},i_{m-1},\ldots,i_{1})$ for all $i_{1},\dots,i_{m}$, and $\boldsymbol{T}_{m}$ be the permutation matrix such that $\operatorname*{vec}(\widetilde{\mbox{$\mathscr{X}$}}_{m})=\boldsymbol{T}_{m}\boldsymbol{x}_{m}$.
+Some properties of $\boldsymbol{T}_{m}$ follow this construction: If we take the $j$-th row vector of $\mbox{$\mathscr{X}$}_{m+1}]_{1}\in\mathbb{R}^{p_{m+1}\times p_{m}\cdots p_{1}}$ as $\mbox{$\mathscr{X}$}_{m+1}]_{1}^{\top}(:,j)$ and the $j$-th column of $\widetilde{\mbox{$\mathscr{X}$}}_{m+1}]_{m}\in\mathbb{R}^{p_{1}p_{2}\cdots p_{m}\times p_{m+1}}$ as $\widetilde{\mbox{$\mathscr{X}$}}_{m+1}]_{m}(:,j)$ for $j\leq p_{m+1}$, then we have
+$\mbox{$\mathscr{X}$}_{m+1}]_{1}^{\top}(:,j)=\boldsymbol{T}_{m}^{-1}[\widetilde{\mbox{$\mathscr{X}$}}_{m+1}]_{m}(:,j)$ (B.19)
+as both sides equal $\operatorname*{vec}(\mbox{$\mathscr{X}$}_{m+1}(j,:,:,\ldots,:))$.
+We now corroborate (B.17) by induction.
+When $M=1$, we have $\widetilde{\mbox{$\mathscr{X}$}}_{1}=\mbox{$\mathscr{X}$}_{1}$, and hence $\boldsymbol{T}_{1}=\boldsymbol{I}_{p_{1}}$, implying $\boldsymbol{G}_{1}^{\top}=\boldsymbol{H}_{1}^{\top}\boldsymbol{T}_{1}$.
+Assume (B.17) holds for $M=m$ with the above construction.
+We will show it also holds for $M=m+1$.
+Note that
+$\boldsymbol{G}_{m+1}^{\top}(\boldsymbol{G}_{m}^{\top}\otimes\boldsymbol{I}_{p_{m+1}})\cdots(\boldsymbol{G}_{2}^{\top}\otimes\boldsymbol{I}_{p_{3}\cdots p_{m+1}})(\boldsymbol{G}_{1}^{\top}\otimes\boldsymbol{I}_{p_{2}\cdots p_{m+1}})\boldsymbol{x}_{m+1}$
+[…]
+[…]
+[…]
+and similarly,
+$\boldsymbol{H}_{m+1}^{\top}(\boldsymbol{I}_{p_{m+1}}\otimes\boldsymbol{H}_{m}^{\top})\cdots(\boldsymbol{I}_{p_{3}\cdots p_{m+1}}\otimes\boldsymbol{H}_{2}^{\top})(\boldsymbol{I}_{p_{2}\cdots p_{m+1}}\otimes\boldsymbol{H}_{1}^{\top})\boldsymbol{T}_{m+1}\boldsymbol{x}_{m+1}$
+[…]
+[…]
+where the last equality follows from (B.19). But by the induction hypothesis, we have
+$\boldsymbol{G}_{m}^{\top}(\boldsymbol{G}_{m-1}^{\top}\otimes\boldsymbol{I}_{p_{m}})\cdots(\boldsymbol{G}_{1}^{\top}\otimes\boldsymbol{I}_{p_{2}\cdots p_{m}})=\boldsymbol{H}_{m}^{\top}(\boldsymbol{I}_{p_{m}}\otimes\boldsymbol{H}_{m-1}^{\top})\cdots(\boldsymbol{I}_{p_{m}\cdots p_{2}}\otimes\boldsymbol{H}_{1}^{\top})\boldsymbol{T}_{m}$
+and thereby we proceed to establish for any $\boldsymbol{M}\in\mathbb{R}^{p_{m+1}\times r_{m}}$,
+$\boldsymbol{G}_{m+1}^{\top}\operatorname*{vec}(\boldsymbol{M})=\boldsymbol{H}_{m+1}^{\top}\operatorname*{vec}(\boldsymbol{M}^{\top}).$
+Inspecting the $i$-th element of both sides for $i\leq r_{m+1}$, we have
+$\boldsymbol{G}_{m+1}^{\top}(i,:)\operatorname*{vec}(\boldsymbol{M})$ $=\sum_{j=1}^{p_{m+1}}\sum_{k=1}^{r_{m}}\boldsymbol{G}_{m+1}^{\top}(i,j+(k-1)p_{m+1})\boldsymbol{M}(j,k),$
+$\boldsymbol{H}_{m+1}^{\top}(i,:)\operatorname*{vec}(\boldsymbol{M}^{\top})$ $=\sum_{j=1}^{p_{m+1}}\sum_{k=1}^{r_{m}}\boldsymbol{H}_{m+1}^{\top}(i,(j-1)r_{m}+k)\boldsymbol{M}^{\top}(k,j).$
+By (B.18), these two quantities are equal, and hence the induction is complete.
+Appendix C Proofs of Two Theorems
+This section provides the detailed proofs of propositions and theorems in Section 2 and 3 of the manuscript.
+C.1 Proof of Theorem 1
+Proof.
+Denote sets
+$\mathcal{W}_{\mathrm{LR}}(\{\boldsymbol{r}_{k}\}_{k=1}^{K_{x}},\{\boldsymbol{s}_{k^{\prime}}\}_{k^{\prime}=1}^{K_{y}},\mathcal{P}_{x}(M,K_{x}),\mathcal{P}_{y}(N,K_{y}))=\{\boldsymbol{\Delta}\in\mathbb{R}^{\prod_{n=1}^{N}q_{n}\times\prod_{m=1}^{M}p_{m}}:$
+$\hskip 28.45274pt\boldsymbol{\Delta}=\boldsymbol{\Lambda}_{y,1}\boldsymbol{\Theta}_{1}\boldsymbol{\Lambda}_{x,1}^{\top}-\boldsymbol{\Lambda}_{y,2}\boldsymbol{\Theta}_{2}\boldsymbol{\Lambda}_{x,2}^{\top};\boldsymbol{\Lambda}_{y,1},\boldsymbol{\Lambda}_{y,2}\in\mathcal{R}_{y}(\{\boldsymbol{s}_{k^{\prime}}\}_{k^{\prime}=1}^{K_{y}},\mathcal{P}_{y}(N,K_{y})),$
+$\hskip 28.45274pt\boldsymbol{\Lambda}_{x,1},\boldsymbol{\Lambda}_{x,2}\in\mathcal{R}_{x}(\{\boldsymbol{r}_{k}\}_{k=1}^{K_{x}},\mathcal{P}_{x}(M,K_{x})),\boldsymbol{\Theta}_{1},\boldsymbol{\Theta}_{2}\in\mathbb{R}^{s_{\mathrm{tot}}\times r_{\mathrm{tot}}},\left\lVert\boldsymbol{\Theta}_{1}\right\rVert_{2}\leq g_{1},\left\lVert\boldsymbol{\Theta}_{2}\right\rVert_{2}\leq g_{1}\},$
+abbreviated to $\mathcal{W}_{\mathrm{LR}}$. Further denote
+$\mathcal{W}_{\mathrm{LR}}(\gamma)=\{\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{LR}}:\left\lVert\boldsymbol{\Delta}\right\rVert_{\mathrm{F}}\geq\gamma\}\quad\text{and}\quad\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma)=\{\boldsymbol{\Delta}^{\prime}=\frac{\boldsymbol{\Delta}}{\left\lVert\boldsymbol{\Delta}\right\rVert_{\mathrm{F}}}:\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{LR}}(\gamma)\},$
+[…]
+Recall the quadratic loss function
+$\mathcal{L}_{T}(\mbox{$\mathscr{A}$})=\frac{1}{T}\sum_{t=1}^{T}\left\lVert\operatorname*{vec}(\mbox{$\mathscr{Y}$}_{t})-[\mbox{$\mathscr{A}$}]_{N}\operatorname*{vec}(\mbox{$\mathscr{X}$}_{t})\right\rVert_{2}^{2}=\frac{1}{T}\sum_{t=1}^{T}\left\lVert\boldsymbol{y}_{t}-[\mbox{$\mathscr{A}$}]_{N}\boldsymbol{x}_{t}\right\rVert_{2}^{2}.$
+By the optimality of $\mbox{$\mathscr{\widehat{A}}$}_{\mathrm{LR}}$, we have
+$\frac{1}{T}\sum_{t=1}^{T}\left\lVert\boldsymbol{y}_{t}-[\mbox{$\mathscr{\widehat{A}}$}_{\mathrm{LR}}]_{N}\boldsymbol{x}_{t}\right\rVert_{2}^{2}\leq\frac{1}{T}\sum_{t=1}^{T}\left\lVert\boldsymbol{y}_{t}-[\mbox{$\mathscr{A}$}_{\mathrm{LR}}^{\ast}]_{N}\boldsymbol{x}_{t}\right\rVert_{2}^{2}$
+$\Rightarrow\penalty 10000\$$\frac{1}{T}\sum_{t=1}^{T}\left\lVert\boldsymbol{\Delta}\boldsymbol{x}_{t}\right\rVert_{2}^{2}\leq\frac{2}{T}\sum_{t=1}^{T}\langle\boldsymbol{e}_{t},\boldsymbol{\Delta}\boldsymbol{x}_{t}\rangle$
+where $\boldsymbol{\Delta}=[\mbox{$\mathscr{\widehat{A}}$}_{\mathrm{LR}}]_{N}-[\mbox{$\mathscr{A}$}_{\mathrm{LR}}^{\ast}]_{N}\in\mathcal{W}_{\mathrm{LR}}$.
+Multiplying both sides by the non-negative indicator function $\mathbbm{1}_{\{\left\lVert\boldsymbol{\Delta}\right\rVert_{\mathrm{F}}\geq\gamma\}}$ for some $\gamma\in(0,1)$ to be specified later, we have
+$\frac{1}{T}\sum_{t=1}^{T}\left\lVert\boldsymbol{\Delta}\boldsymbol{x}_{t}\right\rVert_{2}^{2}\mathbbm{1}_{\{\left\lVert\boldsymbol{\Delta}\right\rVert_{\mathrm{F}}\geq\gamma\}}\leq\frac{2}{T}\sum_{t=1}^{T}\langle\boldsymbol{e}_{t},\boldsymbol{\Delta}\boldsymbol{x}_{t}\rangle\mathbbm{1}_{\{\left\lVert\boldsymbol{\Delta}\right\rVert_{\mathrm{F}}\geq\gamma\}}.$ (C.1)
+Denoting $\boldsymbol{\Delta}_{\geq\gamma}=\boldsymbol{\Delta}\mathbbm{1}_{\{\left\lVert\boldsymbol{\Delta}\right\rVert_{\mathrm{F}}\geq\gamma\}}\in\mathcal{W}_{\mathrm{LR}}(\gamma)\cup\{\boldsymbol{0}\}$, and also noticing that $\mathbbm{1}_{\{\left\lVert\boldsymbol{\Delta}\right\rVert_{\mathrm{F}}\geq\gamma\}}\in\{0,1\}$, we can rewrite (C.1) as
+$\frac{1}{T}\sum_{t=1}^{T}\left\lVert\boldsymbol{\Delta}_{\geq\gamma}\boldsymbol{x}_{t}\right\rVert_{2}^{2}\leq\frac{2}{T}\sum_{t=1}^{T}\langle\boldsymbol{e}_{t},\boldsymbol{\Delta}_{\geq\gamma}\boldsymbol{x}_{t}\rangle,$ (C.2)
+which yields the deterministic bound
+$\left\lVert\boldsymbol{\Delta}_{\geq\gamma}\right\rVert_{\mathrm{F}}\inf_{\boldsymbol{\Delta}^{\prime}\in\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma)}\sum_{t=1}^{T}\left\lVert\boldsymbol{\Delta}^{\prime}\boldsymbol{x}_{t}\right\rVert_{2}^{2}\leq 2\sup_{\boldsymbol{\Delta}^{\prime}\in\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma)}\sum_{t=1}^{T}\langle\boldsymbol{e}_{t},\boldsymbol{\Delta}^{\prime}\boldsymbol{x}_{t}\rangle.$ (C.3)
+Combining Lemmas 2 and 3, we can show when $T\gtrsim\max(1,{\sigma^{4}}/{c_{x}^{2}})(d_{\mathrm{LR}}d_{\mathrm{LR}}^{\prime}(\gamma))$, with probability at least $1-2\exp\left\{-c_{1}d_{\mathrm{LR}}d_{\mathrm{LR}}^{\prime}(\gamma)\right\}$, the following event $\mathcal{E}_{\mathrm{LR}}(\gamma)$ holds:
+$\frac{c_{x}}{2}\left\lVert\boldsymbol{\Delta}_{\geq\gamma}\right\rVert_{\mathrm{F}}\leq 2C\kappa\sigma\sqrt{\frac{d_{\mathrm{LR}}d_{\mathrm{LR}}^{\prime}(\gamma)}{T}}$
+where
+$d_{\mathrm{LR}}=2K_{x}r_{\max}^{2}\sum_{m=1}^{M}p_{m}+2K_{y}s_{\max}^{2}\sum_{n=1}^{N}q_{n}+2K_{x}K_{y}r_{\max}s_{\max},$ (C.4)
+$d_{\mathrm{LR}}^{\prime}(\gamma)=\log\left(\frac{(M\vee N)\sqrt{[(K_{y}s_{\max})\wedge(K_{x}r_{\max})]K_{x}K_{y}(r_{\max}\vee s_{\max})}}{\gamma}\right),$
+$d_{\mathrm{LR}}^{\prime}=d_{\mathrm{LR}}^{\prime}(1)=\log\left((M\vee N)\sqrt{[(K_{y}s_{\max})\wedge(K_{x}r_{\max})]K_{x}K_{y}(r_{\max}\vee s_{\max})}\right),$
+and $C$, $c_{1}$, $c_{2}$, $c_{3}$ are some positive constants.
+Note that $\sigma^{2}\geq\left\lVert\boldsymbol{x}_{t}\right\rVert_{\psi_{2}}^{2}=\sup_{\left\lVert\boldsymbol{u}\right\rVert=1}\left\lVert\langle\boldsymbol{u},\boldsymbol{x}_{t}\rangle\right\rVert_{\psi_{2}}^{2}\gtrsim\sup_{\left\lVert\boldsymbol{u}\right\rVert=1}\left\lVert\langle\boldsymbol{u},\boldsymbol{x}_{t}\rangle\right\rVert_{L_{2}}^{2}\geq c_{x}$, and therefore $\max(1,{\sigma^{4}}/{c_{x}^{2}})\asymp{\sigma^{4}}/{c_{x}^{2}}$.
+Under the event $\mathcal{E}_{\mathrm{LR}}(\gamma)$, it holds that
+$\left\lVert\boldsymbol{\Delta}\right\rVert_{\mathrm{F}}=\left\lVert\boldsymbol{\Delta}_{\geq\gamma}+\boldsymbol{\Delta}{\mathbbm{1}_{\{\left\lVert\boldsymbol{\Delta}\right\rVert_{\mathrm{F}}<\gamma\}}}\right\rVert_{\mathrm{F}}\leq\gamma+\left\lVert\boldsymbol{\Delta}_{\geq\gamma}\right\rVert_{\mathrm{F}}\lesssim\gamma+\frac{\kappa\sigma}{c_{x}}\sqrt{\frac{d_{\mathrm{LR}}d_{\mathrm{LR}}^{\prime}(\gamma)}{T}}.$
+Take $\gamma=\exp(d^{\prime}_{\mathrm{LR}})/T\ll\sqrt{\frac{d_{\mathrm{LR}}}{T}}$.
+With the convexity of $\log$ function, the sample size requirement can be relaxed to $T\gtrsim\frac{\sigma^{4}}{c_{x}^{2}}d_{\mathrm{LR}}\log(\frac{\sigma^{4}}{c_{x}^{2}}d_{\mathrm{LR}})$, and with probability at least $1-2\exp\{-c_{1}d_{\mathrm{LR}}\log T\}$,
+$\left\lVert\boldsymbol{\Delta}\right\rVert_{\mathrm{F}}\lesssim\frac{\kappa\sigma}{c_{x}}\sqrt{\frac{d_{\mathrm{LR}}\log T}{T}}.$
+C.2 Proof of Theorem 2
+Proof.
+Denote sets
+$\mathcal{W}_{\mathrm{AR}}(\{\boldsymbol{r}_{k}\}_{k=1}^{K_{x}},\{\boldsymbol{s}_{k^{\prime}}\}_{k^{\prime}=1}^{K_{y}},\mathcal{P}_{x}(N,K_{x}),\mathcal{P}_{y}(N,K_{y}),L)=\{\boldsymbol{\Delta}\in\mathbb{R}^{\prod_{n=1}^{N}q_{n}\times\prod_{n=1}^{N}q_{n}}:$
+$\hskip 28.45274pt\boldsymbol{\Delta}=\boldsymbol{\Lambda}_{y,1}\boldsymbol{\Theta}_{1}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x,1}^{\top})-\boldsymbol{\Lambda}_{y,2}\boldsymbol{\Theta}_{2}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x,2}^{\top}),\boldsymbol{\Lambda}_{y,1},\boldsymbol{\Lambda}_{y,2}\in\mathcal{R}_{y}(\{\boldsymbol{s}_{k^{\prime}}\}_{k^{\prime}=1}^{K_{y}},\mathcal{P}_{y}(N,K_{y})),$
+$\hskip 28.45274pt\boldsymbol{\Lambda}_{x,1},\boldsymbol{\Lambda}_{x,2}\in\mathcal{R}_{x}(\{\boldsymbol{r}_{k}\}_{k=1}^{K_{x}},\mathcal{P}_{x}(N,K_{x})),\boldsymbol{\Theta}_{1},\boldsymbol{\Theta}_{2}\in\mathbb{R}^{s_{\mathrm{tot}}\times r_{\mathrm{tot}}L},\left\lVert\boldsymbol{\Theta}_{1}\right\rVert_{2}\leq g_{2},\left\lVert\boldsymbol{\Theta}_{2}\right\rVert_{2}\leq g_{2}\},$
+abbreviated to $\mathcal{W}_{\mathrm{AR}}$. Further denote
+$\mathcal{W}_{\mathrm{AR}}(\gamma)=\{\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{AR}}:\left\lVert\boldsymbol{\Delta}\right\rVert_{\mathrm{F}}\geq\gamma\}\quad\text{and}\quad\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma)=\{\boldsymbol{\Delta}^{\prime}=\frac{\boldsymbol{\Delta}}{\left\lVert\boldsymbol{\Delta}\right\rVert_{\mathrm{F}}}:\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{AR}}(\gamma)\},$
+[…]
+By the optimality of estimator $\mbox{$\mathscr{\widehat{A}}$}_{\mathrm{AR}}$, similar to the proof of Theorem 1, we could obtain
+$\frac{1}{T}\sum_{t=1}^{T}\left\lVert\boldsymbol{\Delta}\boldsymbol{x}_{t}\right\rVert_{2}^{2}\leq\frac{2}{T}\sum_{t=1}^{T}\langle\boldsymbol{e}_{t},\boldsymbol{\Delta}\boldsymbol{x}_{t}\rangle,$
+which implies that deterministically,
+$\left\lVert\boldsymbol{\Delta}_{\geq\gamma}\right\rVert_{\mathrm{F}}\inf_{\boldsymbol{\Delta}^{\prime}\in\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma)}\sum_{t=1}^{T}\left\lVert\boldsymbol{\Delta}^{\prime}\boldsymbol{x}_{t}\right\rVert_{2}^{2}\leq 2\sup_{\boldsymbol{\Delta}^{\prime}\in\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma)}\sum_{t=1}^{T}\langle\boldsymbol{e}_{t},\boldsymbol{\Delta}^{\prime}\boldsymbol{x}_{t}\rangle.$ (C.5)
+where $\boldsymbol{\Delta}=[\mbox{$\mathscr{\widehat{A}}$}_{\mathrm{AR}}]_{N}-[\mbox{$\mathscr{A}$}_{\mathrm{AR}}^{\ast}]_{N}\in\mathcal{W}_{\mathrm{AR}}$, and $\boldsymbol{\Delta}_{\geq\gamma}=\boldsymbol{\Delta}\mathbbm{1}_{\{\left\lVert\boldsymbol{\Delta}\right\rVert_{\mathrm{F}}\geq\gamma\}}\in\mathcal{W}_{\mathrm{AR}}(\gamma)\cup\{\boldsymbol{0}\}$.
+Applying Lemmas 4 and 5 to (C.5), we can show when $T\gtrsim\max(\kappa^{4}\kappa_{U,B}^{2}/\kappa_{L,A}^{2},1)d_{\mathrm{AR}}d_{\mathrm{AR}}^{\prime}(\gamma)$, with probability at least $1-c_{1}\exp\left\{-c_{2}d_{\mathrm{AR}}d_{\mathrm{AR}}^{\prime}(\gamma)\right\}$, the following event $\mathcal{E}_{\mathrm{AR}}(\gamma)$ holds:
+$\frac{\kappa_{L,A}}{16}\left\lVert\boldsymbol{\Delta}_{\geq\gamma}\right\rVert_{\mathrm{F}}\leq 2C\kappa^{2}\kappa_{U,B}^{\prime}\sqrt{\frac{d_{\mathrm{AR}}d_{\mathrm{AR}}^{\prime}(\gamma)}{T}}$
+where
+$d_{\mathrm{AR}}=2(K_{x}r_{\max}^{2}+K_{y}s_{\max}^{2})\sum_{n=1}^{N}q_{n}+2LK_{x}K_{y}r_{\max}s_{\max},$ (C.6)
+$d_{\mathrm{AR}}^{\prime}(\gamma)=\log\left(\frac{N\sqrt{[(K_{y}s_{\max})\wedge(LK_{x}r_{\max})]LK_{x}K_{y}(r_{\max}\vee s_{\max})}}{\gamma}\right),$
+$d_{\mathrm{AR}}^{\prime}=d_{\mathrm{AR}}^{\prime}(1)=\log\left(N\sqrt{[(K_{y}s_{\max})\wedge(LK_{x}r_{\max})]LK_{x}K_{y}(r_{\max}\vee s_{\max})}\right),$
+[…]
+and $C$, $c_{1}$, $c_{2}$ are some positive constants. By Lemma 6, $\max(\kappa^{4}\kappa_{U,B}^{2}/\kappa_{L,A}^{2},1)\asymp\kappa^{4}\kappa_{U,B}^{2}/\kappa_{L,A}^{2}$.
+Under the event $\mathcal{E}_{\mathrm{AR}}(\gamma)$, it holds that
+$\left\lVert\boldsymbol{\Delta}\right\rVert_{\mathrm{F}}=\left\lVert\boldsymbol{\Delta}_{\geq\gamma}+\boldsymbol{\Delta}{\mathbbm{1}_{\{\left\lVert\boldsymbol{\Delta}\right\rVert_{\mathrm{F}}<\gamma\}}}\right\rVert_{\mathrm{F}}\leq\gamma+\left\lVert\boldsymbol{\Delta}_{\geq\gamma}\right\rVert_{\mathrm{F}}\lesssim\gamma+\frac{\kappa^{2}\kappa_{U,B}^{\prime}}{\kappa_{L,A}}\sqrt{\frac{d_{\mathrm{AR}}d_{\mathrm{AR}}^{\prime}(\gamma)}{T}}.$
+Take $\gamma=\exp(d^{\prime}_{\mathrm{AR}})/T$.
+Whenever $T\gtrsim\frac{\kappa^{4}\kappa_{U,B}^{2}}{\kappa_{L,A}^{2}}d_{\mathrm{AR}}\log(\frac{\kappa^{4}\kappa_{U,B}^{2}}{\kappa_{L,A}^{2}}d_{\mathrm{AR}})$, with probability at least $1-c_{1}\exp\left\{-c_{2}d_{\mathrm{AR}}\log T\right\}$,
+$\left\lVert\boldsymbol{\Delta}\right\rVert_{\mathrm{F}}\lesssim\frac{\kappa^{2}\kappa_{U,B}^{\prime}}{\kappa_{L,A}}\sqrt{\frac{d_{\mathrm{AR}}\log T}{T}}.$
+Appendix D Auxiliary Lemmas
+This section provides the auxiliary lemmas that are used in establishing the theorems in Section C and their proofs.
+We first introduce the concepts of Gaussian width and Talagrand’s $\gamma_{\alpha}$ functional, which will be frequently utilized in the upcoming lemmas.
+Definition 1 (Gaussian width, Definition 7.5.1 of Vershynin (2018)).
+The Gaussian width of a subset $T\subset\mathbb{R}^{n}$ is defined as
+$\omega(T)\vcentcolon=\mathbb{E}\sup_{\boldsymbol{x}\in T}\langle\boldsymbol{g},\boldsymbol{x}\rangle\quad\text{where}\quad\boldsymbol{g}\sim N(\boldsymbol{0},\boldsymbol{I}_{n}).$
+Definition 2 (Talagrand’s $\gamma_{\alpha}$ functional, Dirksen (2015)).
+Let $(T,d)$ be a metric space and $\alpha>0$.
+A sequence of subsets $\left(T_{k}\right)_{k=0}^{\infty}$ of TT is called an admissible sequence if the cardinalities of $T_{k}$ satisfy $|T_{0}|=1$ and $|T_{k}|\leq 2^{2^{k}}$ for all $k\geq 1$.
+The $\gamma_{\alpha}$ functional of TT is defined as
+$\gamma_{\alpha}(T,d)=\inf_{(T_{k})}\sup_{t\in T}\sum_{k=0}^{\infty}2^{k/\alpha}d(t,T_{k})$
+where the infimum is with respect to all admissible sequences.
+Remark S.2.
+This definition is consistent with Definition 8.5.1 of Vershynin (2018) when $\alpha=2$, motivated by the covering number argument.
+Some of the references in this paper define the $\gamma_{\alpha}$ functional as in Talagrand (2005), based on the notion of ultrametricity.
+These two definitions are equivalent up to a constant factor depending only on $\alpha$ according to Talagrand (2001).
+Lemma 1 (Covering number).
+This lemma derives the covering number for a series of parameter space that are involved in the subsequent proof.
+  • (a)
+    The cardinality of the $\epsilon$-covering net of $\mathcal{R}_{x}(\{\boldsymbol{r}_{k}\}_{k=1}^{K_{x}},\mathcal{P}_{x}(M,K_{x}))$, denoted as $\mathcal{N}(\mathcal{R}_{x},\left\lVert\cdot\right\rVert_{\mathrm{F}},\epsilon)$, satisfies
+    $\mathcal{N}(\mathcal{R}_{x},\left\lVert\cdot\right\rVert_{\mathrm{F}},\epsilon)\leq\left(1+\frac{2M\sqrt{K_{x}r_{\max}}}{\epsilon}\right)^{K_{x}r_{\max}^{2}\sum_{m=1}^{M}p_{m}}.$
+  • (b)
+    The cardinality of the $\epsilon$-covering net of $\mathcal{R}_{y}(\{\boldsymbol{s}_{k^{\prime}}\}_{k^{\prime}=1}^{K_{y}},\mathcal{P}_{y}(N,K_{y}))$, denoted as $\mathcal{N}(\mathcal{R}_{y},\left\lVert\cdot\right\rVert_{\mathrm{F}},\epsilon)$, satisfies
+    $\mathcal{N}(\mathcal{R}_{y},\left\lVert\cdot\right\rVert_{\mathrm{F}},\epsilon)\leq\left(1+\frac{2N\sqrt{K_{y}s_{\max}}}{\epsilon}\right)^{K_{y}s_{\max}^{2}\sum_{n=1}^{N}q_{n}}.$
+  • (c)
+    The cardinality of the $\epsilon$-covering net of $\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma)$, denoted as $\mathcal{N}(\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma),\left\lVert\cdot\right\rVert_{\mathrm{F}},\epsilon)$, with $\gamma\in(0,1)$, satisfies
+    $\mathcal{N}(\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma),\left\lVert\cdot\right\rVert_{\mathrm{F}},\epsilon)\leq\left(1+\frac{24g_{2}^{\prime}N\sqrt{LK_{x}K_{y}(r_{\max}\vee s_{\max})(s_{\mathrm{tot}}\wedge(Lr_{\mathrm{tot}}))}}{\gamma\epsilon}\right)^{d_{\mathrm{AR}}},$
+    where $g_{2}^{\prime}=\sqrt{s_{\mathrm{tot}}\wedge(Lr_{\mathrm{tot}})}g_{2}$, and $d_{\mathrm{AR}}=2(K_{x}r_{\max}^{2}+K_{y}s_{\max}^{2})\sum_{n=1}^{N}q_{n}+2LK_{x}K_{y}r_{\max}s_{\max}$.
+  • (d)
+    The cardinality of the $\epsilon$-covering net of $\mathcal{W}_{\mathrm{LR}}^{\prime}$, denoted as $\mathcal{N}(\mathcal{W}_{\mathrm{LR}}^{\prime},\left\lVert\cdot\right\rVert_{\mathrm{F}},\epsilon)$, satisfies
+    $\mathcal{N}(\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma),\left\lVert\cdot\right\rVert_{\mathrm{F}},\epsilon)\leq\left(1+\frac{24g_{1}^{\prime}(M\vee N)\sqrt{K_{x}K_{y}(r_{\max}\vee s_{\max})(s_{\mathrm{tot}}\wedge r_{\mathrm{tot}})}}{\gamma\epsilon}\right)^{d_{\mathrm{LR}}},$
+    where $g_{1}^{\prime}=\sqrt{s_{\mathrm{tot}}\wedge r_{\mathrm{tot}}}g_{1}$, and $d_{\mathrm{LR}}=2K_{x}r_{\max}^{2}\sum_{m=1}^{M}p_{m}+2K_{y}s_{\max}^{2}\sum_{n=1}^{N}q_{n}+2K_{x}K_{y}r_{\max}s_{\max}$.
+Proof.
+(a) Denote the $\epsilon/(M\sqrt{r_{\mathrm{tot}}})$-covering net of $\mathbb{O}^{p\times r}$ by $\bar{\mathbb{O}}^{p\times r}$.
+By Lemma 7 in Zhang and Xia (2018), the cardinality of $\bar{\mathbb{O}}^{p\times r}$ satisfies $|\bar{\mathbb{O}}^{p\times r}|\leq((1+2M\sqrt{r_{\mathrm{tot}}})/\epsilon)^{pr}$.
+Then for any $\boldsymbol{G}\in\mathbb{O}^{p\times r}$, there exists a $\bar{\boldsymbol{G}}\in\bar{\mathbb{O}}^{p\times r}$ such that $\left\lVert\boldsymbol{G}-\bar{\boldsymbol{G}}\right\rVert_{\mathrm{F}}\leq\epsilon/(M\sqrt{r_{\mathrm{tot}}})$.
+Set the $\epsilon$-covering net of $\mathcal{R}_{x}(\{\boldsymbol{r}_{k}\}_{k=1}^{K_{x}},\mathcal{P}_{x}(M,K_{x}))$, with abbreviation $\mathcal{R}_{x}$, to be
+$\bar{\mathcal{R}}_{x}(\epsilon)=\{$ $\bar{\boldsymbol{\Lambda}}_{x}\in\mathbb{R}^{\prod_{m=1}^{M}p_{m}\times r_{\mathrm{tot}}}:$
+$\bar{\boldsymbol{\Lambda}}_{x}^{\top}=\begin{bmatrix}\bar{\boldsymbol{G}}_{M,1}^{\top}(\boldsymbol{I}_{p_{\alpha^{1}_{(M)}}}\otimes\bar{\boldsymbol{G}}_{M-1,1}^{\top})\cdots(\boldsymbol{I}_{p_{\alpha^{1}_{(2)}}\cdots p_{\alpha^{1}_{(M)}}}\otimes\bar{\boldsymbol{G}}_{1,1}^{\top})\boldsymbol{T}_{x}(\alpha_{1})\\ \bar{\boldsymbol{G}}_{M,2}^{\top}(\boldsymbol{I}_{p_{\alpha^{2}_{(M)}}}\otimes\bar{\boldsymbol{G}}_{M-1,2}^{\top})\cdots(\boldsymbol{I}_{p_{\alpha^{2}_{(2)}}\cdots p_{\alpha^{2}_{(M)}}}\otimes\bar{\boldsymbol{G}}_{1,2}^{\top})\boldsymbol{T}_{x}(\alpha_{2})\\ \cdots\\ \bar{\boldsymbol{G}}_{M,K_{x}}^{\top}(\boldsymbol{I}_{p_{\alpha^{K_{x}}_{(M)}}}\otimes\bar{\boldsymbol{G}}_{M-1,K_{x}}^{\top})\cdots(\boldsymbol{I}_{p_{\alpha^{K_{x}}_{(2)}}\cdots p_{\alpha^{K_{x}}_{(M)}}}\otimes\bar{\boldsymbol{G}}_{1,K_{x}}^{\top})\boldsymbol{T}_{x}(\alpha_{K_{x}})\end{bmatrix}$,
+$\bar{\boldsymbol{G}}_{m,k}\in\bar{\mathbb{O}}^{r_{m-1,k}p_{\alpha^{k}_{(m)}}\times r_{m,k}},\text{ for }1\leq m\leq M,1\leq k\leq K_{x}\}.$
+It remains to show that for any $\boldsymbol{\Lambda}_{x}\in\mathcal{R}_{x}$, there exists a $\bar{\boldsymbol{\Lambda}}_{x}\in\bar{\mathcal{R}}_{x}(\epsilon)$ such that $\left\lVert\boldsymbol{\Lambda}_{x}-\bar{\boldsymbol{\Lambda}}_{x}\right\rVert_{\mathrm{F}}\leq\epsilon$.
+Let $\phi(\boldsymbol{G}_{1,k},\ldots,\boldsymbol{G}_{M,k})=\boldsymbol{G}_{M,k}^{\top}(\boldsymbol{I}_{p_{\alpha^{k}_{(M)}}}\otimes\boldsymbol{G}_{M-1,k}^{\top})\cdots(\boldsymbol{I}_{p_{\alpha^{k}_{(2)}}\cdots p_{\alpha^{k}_{(M)}}}\otimes\boldsymbol{G}_{1,k}^{\top})\boldsymbol{T}_{x}(\alpha_{k})$.
+Note that
+$\left\lVert\boldsymbol{\Lambda}_{x}-\bar{\boldsymbol{\Lambda}}_{x}\right\rVert_{\mathrm{F}}^{2}=$ $\sum_{k=1}^{K_{x}}\left\lVert\phi(\boldsymbol{G}_{1,k},\ldots,\boldsymbol{G}_{M,k})-\phi(\bar{\boldsymbol{G}}_{1,k},\ldots,\bar{\boldsymbol{G}}_{M,k})\right\rVert_{\mathrm{F}}^{2}$
+$\leq$ $\sum_{k=1}^{K_{x}}(\sum_{m=1}^{M}\|\phi(\bar{\boldsymbol{G}}_{1,k},\ldots,\bar{\boldsymbol{G}}_{m-1,k},\boldsymbol{G}_{m,k},\boldsymbol{G}_{m+1,k},\ldots,\boldsymbol{G}_{M,k})$
+$\hskip 42.67912pt-\phi(\bar{\boldsymbol{G}}_{1,k},\ldots,\bar{\boldsymbol{G}}_{m-1,k},\bar{\boldsymbol{G}}_{m,k},\boldsymbol{G}_{m+1,k},\ldots,\boldsymbol{G}_{M,k})\|_{\mathrm{F}})^{2}$
+$=$ $\sum_{k=1}^{K_{x}}(\sum_{m=1}^{M}\|\phi(\bar{\boldsymbol{G}}_{1,k},\ldots,\bar{\boldsymbol{G}}_{m-1,k},\boldsymbol{G}_{m,k}-\bar{\boldsymbol{G}}_{m,k},\boldsymbol{G}_{m+1,k},\ldots,\boldsymbol{G}_{M,k})\|_{\mathrm{F}})^{2}.$
+In the following, we will bound each term of the last equality in the right hand side. For $m=1,2,\ldots,M-1$,
+$\|\phi(\bar{\boldsymbol{G}}_{1,k},\ldots,\bar{\boldsymbol{G}}_{m-1,k},\boldsymbol{G}_{m,k}-\bar{\boldsymbol{G}}_{m,k},\boldsymbol{G}_{m+1,k},\ldots,\boldsymbol{G}_{M,k})\|_{\mathrm{F}}$
+[…]
+[…]
+$\cdots\|\boldsymbol{I}_{p_{\alpha^{k}_{(2)}}\cdots p_{\alpha^{k}_{(M)}}}\otimes\bar{\boldsymbol{G}}_{1,k}^{\top}\|_{2}\|\boldsymbol{T}_{x}(\alpha_{k})\|_{2}$
+[…]
+where the first inequality follows from $\|\boldsymbol{A}\boldsymbol{B}\|_{\mathrm{F}}\leq\|\boldsymbol{A}\|_{\mathrm{F}}\|\boldsymbol{B}\|_{2}$ for any matrices $\boldsymbol{A}\in\mathbb{R}^{p\times r}$ and $\boldsymbol{B}\in\mathbb{R}^{r\times q}$.
+The second inequality uses $\|\boldsymbol{I}_{q}\otimes\boldsymbol{A}\|_{2}=\|\boldsymbol{A}\|_{2}\leq 1$ for any orthogonal matrix $\boldsymbol{A}$.
+When $m=M$, we can bound the term in a similar way,
+$\|\phi(\bar{\boldsymbol{G}}_{1,k},\ldots,\bar{\boldsymbol{G}}_{M-1,k},\boldsymbol{G}_{M,k}-\bar{\boldsymbol{G}}_{M,k})\|_{\mathrm{F}}$
+$=\penalty 10000\$$\|(\boldsymbol{G}_{M,k}^{\top}-\bar{\boldsymbol{G}}_{M,k}^{\top})(\boldsymbol{I}_{p_{\alpha^{k}_{(M)}}}\otimes\bar{\boldsymbol{G}}_{M-1,k}^{\top})\cdots(\boldsymbol{I}_{p_{\alpha^{k}_{(2)}}\cdots p_{\alpha^{k}_{(M)}}}\otimes\bar{\boldsymbol{G}}_{1,k}^{\top})\boldsymbol{T}_{x}(\alpha_{k})\|_{\mathrm{F}}$
+[…]
+[…]
+Thus,
+$\left\lVert\boldsymbol{\Lambda}_{x}-\bar{\boldsymbol{\Lambda}}_{x}\right\rVert_{\mathrm{F}}^{2}\leq$ $\sum_{k=1}^{K_{x}}(\sum_{m=1}^{M}\sqrt{r_{M,k}}\|\boldsymbol{G}_{m,k}-\bar{\boldsymbol{G}}_{m,k}\|_{\mathrm{F}})^{2}\leq\sum_{k=1}^{K_{x}}(\sum_{m=1}^{M}\epsilon\sqrt{r_{M,k}}/(M\sqrt{r_{\mathrm{tot}}}))^{2}$
+$\leq$ $\sum_{k=1}^{K_{x}}(\epsilon\sqrt{r_{M,k}}/\sqrt{r_{\mathrm{tot}}})^{2}\leq\epsilon^{2}.$
+So the cardinality of $\bar{\mathcal{R}}_{x}(\epsilon)$ is
+$\mathcal{N}(\mathcal{R}_{x},\left\lVert\cdot\right\rVert_{\mathrm{F}},\epsilon)\leq$ $\prod_{k=1}^{K_{x}}\prod_{m=1}^{M}\left(\frac{1+2M\sqrt{r_{\mathrm{tot}}}}{\epsilon}\right)^{r_{m-1,k}p_{\alpha_{k}(m)}r_{m,k}}$
+$\leq$ $\left(1+\frac{2M\sqrt{K_{x}r_{\max}}}{\epsilon}\right)^{K_{x}r_{\max}^{2}\sum_{m=1}^{M}p_{m}}.$
+(b) The proof of this lemma is identical to the proof of part (a).
+(c) Denote the radius-$\nu$ ball by $\mathcal{B}_{2}^{d}(\nu)=\{\boldsymbol{b}\in\mathbb{R}^{d}:\lVert\boldsymbol{b}\rVert_{2}\leq\nu\}$.
+By Assumption 6, $\left\lVert\boldsymbol{\Theta}_{\mathrm{AR}}\right\rVert_{\mathrm{F}}\leq\sqrt{s_{\mathrm{tot}}\wedge(Lr_{\mathrm{tot}})}g_{2}\vcentcolon=g_{2}^{\prime}$.
+Set the $\epsilon/(3\sqrt{K_{x}K_{y}})$-covering net of $\mathcal{B}_{2}^{s_{\mathrm{tot}}r_{\mathrm{tot}}L}(g_{2}^{\prime})$ by $\bar{\mathcal{B}}^{s_{\mathrm{tot}}r_{\mathrm{tot}}L}$ with cardinality $|\bar{\mathcal{B}}^{s_{\mathrm{tot}}r_{\mathrm{tot}}L}|\leq(1+6g_{2}^{\prime}\sqrt{K_{x}K_{y}}/\epsilon)^{s_{\mathrm{tot}}r_{\mathrm{tot}}L}$.
+We first construct the $\epsilon$-covering net of $\mathcal{S}_{\mathrm{AR}}(\{\boldsymbol{r}_{k}\}_{k=1}^{K_{x}},\{\boldsymbol{s}_{k^{\prime}}\}_{k^{\prime}=1}^{K_{y}},\mathcal{P}_{x}(M,K_{x}),\mathcal{P}_{y}(N,K_{y}),L)$, with abbreviation $\mathcal{S}_{\mathrm{AR}}$, to be
+$\bar{\mathcal{S}}_{\mathrm{AR}}(\epsilon)=\{$ $\bar{\mbox{$\mathscr{A}$}}\in\mathbb{R}^{q_{1}\times\cdots\times q_{N}\times p_{1}\times\cdots\times p_{M}\times L}:[\bar{\mbox{$\mathscr{A}$}}]_{N}=\bar{\boldsymbol{\Lambda}}_{y}\bar{\boldsymbol{\Theta}}_{\mathrm{AR}}(\boldsymbol{I}_{L}\otimes\bar{\boldsymbol{\Lambda}}_{x}^{\top}),$
+$\bar{\boldsymbol{\Lambda}}_{y}\in\bar{\mathcal{R}}_{y}(\epsilon/(3g_{2}\sqrt{K_{x}})),\bar{\boldsymbol{\Lambda}}_{x}\in\bar{\mathcal{R}}_{x}(\epsilon/(3g_{2}\sqrt{LK_{y}})),$
+$\bar{\boldsymbol{\Theta}}_{\mathrm{AR}}\in\mathbb{R}^{s_{\mathrm{tot}}\times r_{\mathrm{tot}}L}\text{ and }\operatorname*{vec}(\bar{\boldsymbol{\Theta}}_{\mathrm{AR}})\in\bar{\mathcal{B}}^{s_{\mathrm{tot}}r_{\mathrm{tot}}L}\}.$
+It remains to show that for any $\mbox{$\mathscr{A}$}\in\mathcal{S}_{\mathrm{AR}}$, there exists a $\bar{\mbox{$\mathscr{A}$}}\in\bar{\mathcal{S}}_{\mathrm{AR}}(\epsilon)$ such that $\left\lVert\mbox{$\mathscr{A}$}-\bar{\mbox{$\mathscr{A}$}}\right\rVert_{\mathrm{F}}\leq\epsilon$.
+Note that
+$\left\lVert\mbox{$\mathscr{A}$}-\bar{\mbox{$\mathscr{A}$}}\right\rVert_{\mathrm{F}}$ $=\left\lVert\boldsymbol{\Lambda}_{y}\boldsymbol{\Theta}_{\mathrm{AR}}(\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top})-\bar{\boldsymbol{\Lambda}}_{y}\bar{\boldsymbol{\Theta}}_{\mathrm{AR}}(\boldsymbol{I}_{L}\otimes\bar{\boldsymbol{\Lambda}}_{x}^{\top})\right\rVert_{\mathrm{F}}$
+$\leq\left\lVert\boldsymbol{\Lambda}_{y}-\bar{\boldsymbol{\Lambda}}_{y}\right\rVert_{\mathrm{F}}\left\lVert\boldsymbol{\Theta}_{\mathrm{AR}}\right\rVert_{2}\left\lVert\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top}\right\rVert_{2}+\left\lVert\bar{\boldsymbol{\Lambda}}_{y}\right\rVert_{2}\left\lVert\boldsymbol{\Theta}_{\mathrm{AR}}-\bar{\boldsymbol{\Theta}}_{\mathrm{AR}}\right\rVert_{\mathrm{F}}\left\lVert\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top}\right\rVert_{2}$
+$+\left\lVert\bar{\boldsymbol{\Lambda}}_{y}\right\rVert_{2}\left\lVert\bar{\boldsymbol{\Theta}}_{\mathrm{AR}}\right\rVert_{2}\left\lVert\boldsymbol{I}_{L}\otimes(\boldsymbol{\Lambda}_{x}^{\top}-\bar{\boldsymbol{\Lambda}}_{x}^{\top})\right\rVert_{\mathrm{F}}$
+$\leq g_{2}\sqrt{K_{x}}\left\lVert\boldsymbol{\Lambda}_{y}-\bar{\boldsymbol{\Lambda}}_{y}\right\rVert_{\mathrm{F}}+\sqrt{K_{x}K_{y}}\left\lVert\boldsymbol{\Theta}_{\mathrm{AR}}-\bar{\boldsymbol{\Theta}}_{\mathrm{AR}}\right\rVert_{\mathrm{F}}+g_{2}\sqrt{LK_{y}}\left\lVert\boldsymbol{\Lambda}_{x}-\bar{\boldsymbol{\Lambda}}_{x}\right\rVert_{\mathrm{F}}$
+$\leq\epsilon,$
+where the second inequality uses $\left\lVert\boldsymbol{\Theta}_{\mathrm{AR}}\right\rVert_{2}\leq g_{2}$ and $\left\lVert\bar{\boldsymbol{\Theta}}_{\mathrm{AR}}\right\rVert_{2}\leq g_{2}$ by assumption 6.
+And we have $\left\lVert\bar{\boldsymbol{\Lambda}}_{y}\right\rVert_{2}\leq\sqrt{K_{y}}$ and $\left\lVert\boldsymbol{I}_{L}\otimes\boldsymbol{\Lambda}_{x}^{\top}\right\rVert_{2}=\left\lVert\boldsymbol{\Lambda}_{x}^{\top}\right\rVert_{2}\leq\sqrt{K_{x}}$ given that $\bar{\boldsymbol{\Lambda}}_{y}$ and $\boldsymbol{\Lambda}_{x}$ are composed of some orthogonal matrices.
+By Lemma 1(a) and (b), the cardinality of $\bar{\mathcal{S}}_{\mathrm{AR}}(\epsilon)$ is
+$\mathcal{N}(\mathcal{S}_{\mathrm{AR}},\left\lVert\cdot\right\rVert_{\mathrm{F}},\epsilon)\leq$ $\left(1+\frac{6g_{2}N\sqrt{LK_{x}K_{y}r_{\max}}}{\epsilon}\right)^{K_{x}r_{\max}^{2}\sum_{n=1}^{N}q_{n}}\left(1+\frac{6g_{2}^{\prime}\sqrt{K_{x}K_{y}}}{\epsilon}\right)^{LK_{x}K_{y}r_{\max}s_{\max}}$
+$\left(1+\frac{6g_{2}N\sqrt{K_{x}K_{y}s_{\max}}}{\epsilon}\right)^{K_{y}s_{\max}^{2}\sum_{n=1}^{N}q_{n}}$
+$\leq$ $\left(1+\frac{6g_{2}^{\prime}N\sqrt{LK_{x}K_{y}(r_{\max}\vee s_{\max})}}{\epsilon}\right)^{(K_{x}r_{\max}^{2}+K_{y}s_{\max}^{2})\sum_{n=1}^{N}q_{n}+LK_{x}K_{y}r_{\max}s_{\max}}.$
+Now we are ready to construct the $\epsilon$-covering net of $\mathcal{W}_{\mathrm{AR}}$.
+Note that for any $\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{AR}}$, we have $\boldsymbol{\Delta}=[\mbox{$\mathscr{A}$}_{1}]_{N}-[\mbox{$\mathscr{A}$}_{2}]_{N}$ with $\mbox{$\mathscr{A}$}_{1},\mbox{$\mathscr{A}$}_{2}\in\mathcal{S}_{\mathrm{AR}}$.
+Then the covering net can be constructed as
+$\bar{\mathcal{W}}_{\mathrm{AR}}=\{$ $\bar{\boldsymbol{\Delta}}\in\mathbb{R}^{\prod_{n=1}^{N}q_{n}\times L\prod_{m=1}^{M}p_{m}}:\bar{\boldsymbol{\Delta}}=[\bar{\mbox{$\mathscr{A}$}}_{1}]_{N}-[\bar{\mbox{$\mathscr{A}$}}_{2}]_{N},\bar{\mbox{$\mathscr{A}$}}_{1},\bar{\mbox{$\mathscr{A}$}}_{2}\in\bar{\mathcal{S}}_{\mathrm{AR}}(\epsilon/2)\}.$
+It can be easily verified that $\left\lVert\boldsymbol{\Delta}-\bar{\boldsymbol{\Delta}}\right\rVert_{\mathrm{F}}\leq\left\lVert[\mbox{$\mathscr{A}$}_{1}]_{N}-[\bar{\mbox{$\mathscr{A}$}}_{1}]_{N}\right\rVert_{\mathrm{F}}+\left\lVert[\mbox{$\mathscr{A}$}_{2}]_{N}-[\bar{\mbox{$\mathscr{A}$}}_{2}]_{N}\right\rVert_{\mathrm{F}}\leq\epsilon$.
+And the cardinality of $\bar{\mathcal{W}}_{\mathrm{AR}}$ is given by
+$\penalty 10000\ \mathcal{N}(\mathcal{W}_{\mathrm{AR}},\left\lVert\cdot\right\rVert_{\mathrm{F}},\epsilon)$
+$\leq$ $\penalty 10000\ \mathcal{N}\left(\mathcal{S}_{\mathrm{AR}},\left\lVert\cdot\right\rVert_{\mathrm{F}},\frac{\epsilon}{2}\right)^{2}$
+$\leq$ $\left(1+\frac{12g_{2}^{\prime}(M\vee N)\sqrt{LK_{x}K_{y}(r_{\max}\vee s_{\max})}}{\epsilon}\right)^{2(K_{x}r_{\max}^{2}+K_{y}s_{\max}^{2})\sum_{n=1}^{N}q_{n}+2LK_{x}K_{y}r_{\max}s_{\max}}.$ (D.1)
+Note that $\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma)=\{\boldsymbol{\Delta}/\left\lVert\boldsymbol{\Delta}\right\rVert_{\mathrm{F}}:\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{AR}},\left\lVert\boldsymbol{\Delta}\right\rVert_{\mathrm{F}}>\gamma\}\subset\mathcal{W}_{\mathrm{AR}}/\gamma\vcentcolon=\{\boldsymbol{\Delta}/\gamma:\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{AR}}\}$.
+Each element in $\mathcal{W}_{\mathrm{AR}}/\gamma$ is a scaled version of an element in $\mathcal{W}_{\mathrm{AR}}$ by a factor $1/\gamma$, which can be multiplied to $\boldsymbol{\Theta}_{\mathrm{AR}}$ and $\bar{\boldsymbol{\Theta}}_{\mathrm{AR}}$ with their upper bound $g_{2}$ replaced by $g_{2}/\gamma$ and other components unchanged.
+Thus, the covering number of $\mathcal{W}_{\mathrm{AR}}/\gamma$ can be obtained by replacing $g_{2}^{\prime}$ with $g_{2}^{\prime}/\gamma$ in (D.1).
+Denote $\mathcal{D}(\cdot,\left\lVert\cdot\right\rVert,\epsilon)$ as the $\epsilon$-packing number as in Definition 5.4 of Wainwright (2019).
+By the relationship between covering number and packing number (Wainwright (2019), Lemma 5.5), we have $\mathcal{N}(\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma),\left\lVert\cdot\right\rVert_{\mathrm{F}},\epsilon)\leq\mathcal{D}(\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma),\left\lVert\cdot\right\rVert_{\mathrm{F}},\epsilon)\leq\mathcal{D}(\mathcal{W}_{\mathrm{AR}}/\gamma,\left\lVert\cdot\right\rVert_{\mathrm{F}},\epsilon)\leq\mathcal{N}(\mathcal{W}_{\mathrm{AR}}/\gamma,\left\lVert\cdot\right\rVert_{\mathrm{F}},\epsilon/2)$, yielding the desired result.
+(d) This is similar to part (c) with $L=1$ and $g_{1}^{\prime}\vcentcolon=\sqrt{s_{\mathrm{tot}}\wedge r_{\mathrm{tot}}}g_{1}$. ∎
+Lemma 2 (Truncated Regression RSC).
+Suppose Assumptions 2, 3, 4 hold with $T\gtrsim\frac{\sigma^{4}}{c_{x}^{2}}(d_{\mathrm{LR}}d_{\mathrm{LR}}^{\prime}(\gamma))$ for some $\gamma\in(0,1)$. Then
+$\frac{1}{2}c_{x}\left\lVert\boldsymbol{\Delta}\right\rVert_{\mathrm{F}}^{2}\leq\frac{1}{T}\sum_{t=1}^{T}\left\lVert\boldsymbol{\Delta}\boldsymbol{x}_{t}\right\rVert_{2}^{2}\leq 2C_{x}\left\lVert\boldsymbol{\Delta}\right\rVert_{\mathrm{F}}^{2}\quad\text{for all }\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{LR}}(\gamma)$
+with probability at least $1-\exp\left\{-c_{1}d_{\mathrm{LR}}d_{\mathrm{LR}}^{\prime}(\gamma)\right\}$, where $c_{1}>0$ is a constant, and $d_{\mathrm{LR}}$, $d_{\mathrm{LR}}^{\prime}(\gamma)$ are defined in (C.4).
+Proof.
+This lemma establishes the probabilistic uniform convexity condition for the empirical loss function confined on the set $\mathcal{W}_{\mathrm{LR}}(\gamma)$.
+By the definition of $\mathcal{W}_{\mathrm{LR}}(\gamma)$ and $\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma)$, it suffices to show that $\sup_{\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma)}\sum_{t=1}^{T}\left\lVert\boldsymbol{\Delta}\boldsymbol{x}_{t}\right\rVert_{2}^{2}/T\leq 2C_{x}$ and $\inf_{\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma)}\sum_{t=1}^{T}\left\lVert\boldsymbol{\Delta}\boldsymbol{x}_{t}\right\rVert_{2}^{2}/T\geq c_{x}/2$ hold with high probability.
+Truncating $\mathcal{B}_{\mathrm{F}}^{Q\times\prod_{m=1}^{M}p_{m}}(\gamma)\vcentcolon=\{\boldsymbol{\Delta}\in\mathbb{R}^{Q\times\prod_{m=1}^{M}p_{m}}:\lVert\boldsymbol{\Delta}\rVert_{\mathrm{F}}\leq\gamma\}$ out from the restricted strong convexity (RSC) condition avoids the anomalous behavior of the normalized process when $\boldsymbol{\Delta}$ approaches the origin.
+We first bound the $\psi_{1}$-norm of $\left\lVert\boldsymbol{\Delta}_{0}\boldsymbol{x}_{t}\right\rVert_{2}^{2}$ for a fixed $\boldsymbol{\Delta}_{0}\in\mathbb{R}^{Q\times\prod_{m=1}^{M}p_{m}}$ with unit Frobenius norm.
+Note that $\left\lVert\langle\boldsymbol{\Delta}_{0}^{j},\boldsymbol{x}_{t}\rangle\right\rVert_{\psi_{2}}^{2}\leq\sigma^{2}\left\lVert\boldsymbol{\Delta}_{0}^{j}\right\rVert_{2}^{2}$, where $\boldsymbol{\Delta}_{0}^{j}$ is the $j$-th row vector of $\boldsymbol{\Delta}_{0}$.
+Hence,
+$\left\lVert\left\lVert\boldsymbol{\Delta}_{0}\boldsymbol{x}_{t}\right\rVert_{2}^{2}\right\rVert_{\psi_{1}}$ $=\left\lVert\sum_{j=1}^{Q}\langle\boldsymbol{\Delta}_{0}^{j},\boldsymbol{x}_{t}\rangle^{2}\right\rVert_{\psi_{1}}\leq\sum_{j=1}^{Q}\left\lVert\langle\boldsymbol{\Delta}_{0}^{j},\boldsymbol{x}_{t}\rangle^{2}\right\rVert_{\psi_{1}}=\sum_{j=1}^{Q}\left\lVert\langle\boldsymbol{\Delta}_{0}^{j},\boldsymbol{x}_{t}\rangle\right\rVert_{\psi_{2}}^{2}$
+$\leq\sigma^{2}\sum_{j=1}^{Q}\left\lVert\boldsymbol{\Delta}_{0}^{j}\right\rVert_{\mathrm{F}}^{2}=\sigma^{2}\left\lVert\boldsymbol{\Delta}_{0}\right\rVert_{\mathrm{F}}^{2}=\sigma^{2},$
+which implies that $\left\lVert\left\lVert\boldsymbol{\Delta}_{0}\boldsymbol{x}_{t}\right\rVert_{2}\right\rVert_{\psi_{2}}\leq\sigma$.
+To proceed from the pointwise sub-Gaussian property to a uniform concentration over the set $\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma)$, we utilize the generic chaining technique in Theorem 10 from Banerjee and et al.
+(2015).
+Specifically, we consider the following class of functions indexed by $\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma)$:
+[…]
+Let $(\Omega,\mu)$ be the probability space where $\boldsymbol{x}_{t}$’s are defined. For any $f_{\boldsymbol{\Delta}}\in F$,
+$\left\lVert f_{\boldsymbol{\Delta}}\right\rVert_{L_{2}(\mu)}^{2}=\frac{1}{\operatorname*{tr}(\boldsymbol{\Delta}\boldsymbol{\Sigma}_{x}\boldsymbol{\Delta}^{\top})}\mathbb{E}[\operatorname*{tr}(\boldsymbol{x}_{t}^{\top}\boldsymbol{\Delta}^{\top}\boldsymbol{\Delta}\boldsymbol{x}_{t})]=1,$
+so that $F$ is a subset of the unit sphere of $L_{2}(\mu)$, i.e., $F\subseteq S_{L_{2}(\mu)}$.
+Meanwhile, $\sup_{f_{\boldsymbol{\Delta}}\in F}\left\lVert f_{\boldsymbol{\Delta}}\right\rVert_{\psi_{2}}=\sup_{\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma)}\left\lVert\frac{\left\lVert\boldsymbol{\Delta}\boldsymbol{x}_{t}\right\rVert_{2}}{\sqrt{\operatorname*{tr}(\boldsymbol{\Delta}\boldsymbol{\Sigma}_{x}\boldsymbol{\Delta}^{\top})}}\right\rVert_{\psi_{2}}\leq\sigma/\sqrt{c_{x}}$.
+As a result,
+$\gamma_{2}(F,\left\lVert\cdot\right\rVert_{\psi_{2}})\leq\sigma\gamma_{2}(F,\left\lVert\cdot\right\rVert_{L_{2}(\mu)})/\sqrt{c_{x}}\leq c_{1}\sigma\omega(\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma))/\sqrt{c_{x}}.$
+In the context of Theorem 10 from Banerjee and et al.
+(2015), let $\sqrt{T}\geq 2c_{1}c_{2}\sigma^{2}\omega(\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma))/c_{x}$ and $\theta=1/2$, then the condition $\theta\sqrt{T}\geq c_{2}\sigma\gamma_{2}(F,\left\lVert\cdot\right\rVert_{\psi_{2}})/\sqrt{c_{x}}$ is satisfied.
+Therefore, by Theorem 10 from Banerjee and et al.
+(2015), we have
+$\sup_{\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma)}\left|\frac{1}{T}\frac{1}{\operatorname*{tr}(\boldsymbol{\Delta}\boldsymbol{\Sigma}_{x}\boldsymbol{\Delta}^{\top})}\sum_{t=1}^{T}\left\lVert\boldsymbol{\Delta}\boldsymbol{x}_{t}\right\rVert_{2}^{2}-1\right|\leq\frac{1}{2}$
+with probability at least $1-\exp\left\{-\eta c_{x}^{2}T/\sigma^{4}\right\}$, where $\eta$ is some positive constant.
+But by Assumption 2 and Rayleigh’s quotient, $\operatorname*{tr}(\boldsymbol{\Delta}\boldsymbol{\Sigma}_{x}\boldsymbol{\Delta}^{\top})=\sum_{j=1}^{Q}\boldsymbol{\Delta}^{j}{}^{\top}\boldsymbol{\Sigma}_{x}\boldsymbol{\Delta}^{j}\in[c_{x},C_{x}]$ for any $\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma)$.
+Putting together the pieces, we have when $T\gtrsim\sigma^{4}\omega^{2}(\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma))/c_{x}^{2}$,
+$\frac{1}{2}c_{x}\leq\inf_{\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma)}\frac{1}{T}\sum_{t=1}^{T}\left\lVert\boldsymbol{\Delta}\boldsymbol{x}_{t}\right\rVert_{2}^{2}\leq\sup_{\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma)}\frac{1}{T}\sum_{t=1}^{T}\left\lVert\boldsymbol{\Delta}\boldsymbol{x}_{t}\right\rVert_{2}^{2}\leq\frac{3}{2}C_{x}$
+with probability at least $1-\exp\left\{-\eta c_{x}^{2}T/\sigma^{4}\right\}$.
+Finally, we study the Gaussian width $\omega(\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma))$. By Dudley’s inequality and Lemma 1(d), we have
+$\omega(\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma))$ $\leq C_{1}\int_{0}^{2}\sqrt{\log\mathcal{N}\left(\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma),\left\lVert\cdot\right\rVert_{2},\epsilon\right)}\mathrm{d}\epsilon$
+$\leq C_{2}\int_{0}^{2}\left[d_{\mathrm{LR}}\log\left(\frac{24(M\vee N)\sqrt{[(K_{y}s_{\max})\wedge(K_{x}r_{\max})]K_{x}K_{y}(r_{\max}\vee s_{\max})}}{\gamma\epsilon}\right)\right]^{1/2}\mathrm{d}\epsilon$
+$\leq C_{3}d_{\mathrm{LR}}^{1/2}\left[\log\left(\frac{(M\vee N)\sqrt{[(K_{y}s_{\max})\wedge(K_{x}r_{\max})]K_{x}K_{y}(r_{\max}\vee s_{\max})}}{\gamma}\right)\right]^{1/2}$
+$=C_{3}\sqrt{d_{\mathrm{LR}}d_{\mathrm{LR}}^{\prime}(\gamma)},$ (D.2)
+where $d_{\mathrm{LR}},d_{\mathrm{LR}}^{\prime}(\gamma)$ are defined in (C.4).
+The third inequality follows from Mill’s ratio after integration by parts.
+Then, when $T\gtrsim(\sigma^{4}/c_{x}^{2})(d_{\mathrm{LR}}d_{\mathrm{LR}}^{\prime}(\gamma))$, with probability at least $1-\exp\left\{-c_{1}d_{\mathrm{LR}}d_{\mathrm{LR}}^{\prime}(\gamma)\right\}$,
+$\frac{c_{x}}{2}\left\lVert\boldsymbol{\Delta}\right\rVert_{\mathrm{F}}^{2}\leq\frac{1}{T}\sum_{t=1}^{T}\left\lVert\boldsymbol{\Delta}\boldsymbol{x}_{t}\right\rVert_{2}^{2}\leq 2C_{x}\left\lVert\boldsymbol{\Delta}\right\rVert_{\mathrm{F}}^{2}$
+holds simultaneously for all $\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{LR}}(\gamma)$. ∎
+Lemma 3 (Truncated regression deviation bound).
+Suppose that Assumptions 2, 3, 4 hold. Then, when $T\gtrsim d_{\mathrm{LR}}d_{\mathrm{LR}}^{\prime}(\gamma)$,
+$\sup_{\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma)}\frac{1}{T}\sum_{t=1}^{T}\langle\boldsymbol{e}_{t},\boldsymbol{\Delta}\boldsymbol{x}_{t}\rangle\leq C\kappa\sigma\sqrt{d_{\mathrm{LR}}d_{\mathrm{LR}}^{\prime}(\gamma)/T}$
+with probability at least $1-\exp\left\{-cd_{\mathrm{LR}}d_{\mathrm{LR}}^{\prime}(\gamma)\right\}$, where $d_{\mathrm{LR}},d_{\mathrm{LR}}^{\prime}(\gamma)$ are defined in (C.4), and $C,c$ are some positive constants.
+Proof.
+The key approach to proving this lemma is to utilize Lemma 7 for the random process $\boldsymbol{W}(\boldsymbol{\Delta}):=\frac{1}{T}\sum_{t=1}^{T}\langle\boldsymbol{e}_{t},\boldsymbol{\Delta}\boldsymbol{x}_{t}\rangle$.
+To apply Lemma 7, we first need to verify the mixed tail condition.
+For any $\boldsymbol{\Delta}_{a},\boldsymbol{\Delta}_{b}\in\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma)\cup\{\boldsymbol{0}\}$ and $0<\lambda\leq c/\left(\kappa\sigma\left\lVert(\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b})/T\right\rVert_{\mathrm{op}}\right)$, we have
+$\mathbb{P}\left\{\boldsymbol{W}(\boldsymbol{\Delta}_{a})-\boldsymbol{W}(\boldsymbol{\Delta}_{b})\geq u\right\}$ $\leq\exp(-\lambda u)\mathbb{E}\left[\exp\left(\lambda\sum_{t=1}^{T}\langle\boldsymbol{e}_{t},(\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b})\boldsymbol{x}_{t}/T\rangle\right)\right]$
+$=\exp(-\lambda u)\prod_{t=1}^{T}\mathbb{E}_{x}\left\{\mathbb{E}_{e}\left[\exp\left(\lambda\langle\boldsymbol{e}_{t},(\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b})\boldsymbol{x}_{t}/T\rangle\right)|\boldsymbol{x}_{t}\right]\right\}$
+$\leq\exp(-\lambda u)\prod_{t=1}^{T}\mathbb{E}\left[\exp\left(C_{1}\lambda^{2}\kappa^{2}\left\lVert(\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b})\boldsymbol{x}_{t}/T\right\rVert_{2}^{2}\right)\right]$
+$\leq\exp\left(C_{2}\lambda^{2}T\kappa^{2}\sigma^{2}\left\lVert(\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b})/T\right\rVert_{\mathrm{F}}^{2}-\lambda u\right),$
+where the first inequality is a direct conclusion from Markov’s inequality and the definition of $\boldsymbol{W}(\boldsymbol{\Delta})$, the first equality is due to the independence condition in Assumptions 2 and 3, and the last inequality is from Vershynin (2018), Exercise 6.2.6.
+We have also used the fact that, conditional on $\boldsymbol{x}_{t}$, $\langle\boldsymbol{e}_{t},(\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b})\boldsymbol{x}_{t}/T\rangle$ is a sub-Gaussian random variable with parameter $\kappa^{2}\left\lVert(\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b})\boldsymbol{x}_{t}/T\right\rVert_{2}^{2}$.
+By choosing $\lambda$ according to the sub-exponential tail bound, we get
+$\mathbb{P}\left\{\boldsymbol{W}(\boldsymbol{\Delta}_{a})-\boldsymbol{W}(\boldsymbol{\Delta}_{b})\geq u\right\}\leq\exp\left\{-\min\left(\frac{u^{2}}{C_{3}^{2}\kappa^{2}\sigma^{2}\left\lVert\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b}\right\rVert_{\mathrm{F}}^{2}/T},\frac{u}{C_{3}\kappa\sigma\left\lVert\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b}\right\rVert_{\mathrm{F}}/T}\right)\right\}.$
+The other direction of the tail bound can be easily verified with the same approach.
+By the resulting fixed-point mixed tail for $\boldsymbol{W}(\boldsymbol{\Delta})$, we let $d_{1}(\boldsymbol{\Delta}_{a},\boldsymbol{\Delta}_{b})=C_{3}\kappa\sigma\left\lVert\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b}\right\rVert_{\mathrm{F}}/T$, and $d_{2}(\boldsymbol{\Delta}_{a},\boldsymbol{\Delta}_{b})=C_{3}\kappa\sigma\left\lVert\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b}\right\rVert_{\mathrm{F}}/\sqrt{T}$.
+By Theorem 8.6.1 of Vershynin (2018) and Eq.
+(46) and Lemma 2.7 of Melnyk and Banerjee (2016), we have
+$\gamma_{2}(\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma),d_{2})\leq\left(C_{3}\kappa\sigma/\sqrt{T}\right)\gamma_{2}(\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma),\left\lVert\cdot\right\rVert_{\mathrm{F}})\leq\left(C_{5}\kappa\sigma/\sqrt{T}\right)\omega(\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma)),$
+$\gamma_{1}(\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma),d_{1})\leq\left(C_{3}\kappa\sigma/T\right)\gamma_{2}^{2}(\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma),\left\lVert\cdot\right\rVert_{\mathrm{F}})\leq\left(C_{6}\kappa\sigma/T\right)\omega^{2}(\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma)),$
+$\mathrm{Diam}_{2}(\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma))=2C_{3}\kappa\sigma/\sqrt{T},\quad\mathrm{Diam}_{1}(\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma))=2C_{3}\kappa\sigma/T.$
+Note that adding the origin $\boldsymbol{0}$ to the set $\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma)$ does not change its diameter or its Gaussian width, as $\boldsymbol{0}$ is already in the convex hull of $\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma)$.
+Let $u=\omega(\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma))$ given by (D.2).
+Substituting all above terms into Lemma 7, we establish that when $T\gtrsim d_{\mathrm{LR}}d_{\mathrm{LR}}^{\prime}(\gamma)$,
+$\sup_{\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{LR}}^{\prime}(\gamma)}\boldsymbol{W}(\boldsymbol{\Delta})\leq C\kappa\sigma\sqrt{d_{\mathrm{LR}}d_{\mathrm{LR}}^{\prime}(\gamma)/T}$
+with probability at least $1-\exp\{-cd_{\mathrm{LR}}d_{\mathrm{LR}}^{\prime}(\gamma)\}$. ∎
+Lemma 4 (Truncated Autoregression RSC).
+Suppose Assumptions 5, 1, 6 hold. When $T\gtrsim(\kappa^{4}\kappa_{U,B}^{2}/\kappa_{L,A}^{2})d_{\mathrm{AR}}d_{\mathrm{AR}}^{\prime}(\gamma)$, we have
+$\frac{\kappa_{L,A}}{16}\|\boldsymbol{\Delta}\|_{\mathrm{F}}^{2}\leq\frac{1}{T}\sum_{t=1}^{T}\left\lVert\boldsymbol{\Delta}\boldsymbol{x}_{t}\right\rVert_{2}^{2}\leq 4\kappa_{U,A}\|\boldsymbol{\Delta}\|_{\mathrm{F}}^{2}\quad\text{for all }\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{AR}}(\gamma)$
+with probability at least $1-2\exp\left\{-cd_{\mathrm{AR}}d_{\mathrm{AR}}^{\prime}(\gamma)\right\}$, where $c>0$ is a constant, $d_{\mathrm{AR}}$, $d_{\mathrm{LR}}^{\prime}(\gamma)$ are defined in (C.6), and $\kappa_{U,B}=C_{e}/\mu_{\mathrm{min}}(\boldsymbol{B}^{\ast})$, $\kappa_{L,A}=c_{e}/\mu_{\mathrm{max}}(\mbox{$\mathscr{A}$}^{\ast})$, $\kappa_{U,A}=C_{e}/\mu_{\mathrm{min}}(\mbox{$\mathscr{A}$}^{\ast})$.
+Proof.
+We define the semi-norm $\left\lVert\boldsymbol{\Delta}\right\rVert_{T}=\sqrt{\frac{1}{T}\sum_{t=1}^{T}\left\lVert\boldsymbol{\Delta}\boldsymbol{x}_{t}\right\rVert_{2}^{2}}$.
+It is sufficient to show that $\sup_{\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma)}\left\lVert\boldsymbol{\Delta}\right\rVert_{T}^{2}\leq 4\kappa_{U,A}$ and $\inf_{\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma)}\left\lVert\boldsymbol{\Delta}\right\rVert_{T}^{2}\geq\kappa_{L,A}/16$ hold with high probability.
+For any coefficient tensor $\mbox{$\mathscr{A}$}\in\mathcal{S}_{\mathrm{AR}}$, denote $\boldsymbol{A}=[\mbox{$\mathscr{A}$}]_{N}$.
+Then we have $\boldsymbol{A}=(\boldsymbol{A}_{1},\dots,\boldsymbol{A}_{L})$ where $\boldsymbol{A}_{l}=[\mbox{$\mathscr{A}$}_{l}]_{N}\in\mathbb{R}^{Q\times Q}$ and $Q=\prod_{n=1}^{N}q_{n}$.
+We first rewrite the VAR($L$) representation $\boldsymbol{y}_{t}=\sum_{l=1}^{L}\boldsymbol{A}_{l}\boldsymbol{y}_{t-l}+\boldsymbol{e}_{t}$ into a VAR($1$) form, $\boldsymbol{x}_{t}=\boldsymbol{B}\boldsymbol{x}_{t-1}+\boldsymbol{\zeta}_{t}$.
+From Luetkepohl (2005) we know $\{\boldsymbol{x}_{t}\}$ is also stationary.
+Also as a centered process, by Wold representation theorem, $\{\boldsymbol{x}_{t}\}$ admits of a VMA($\infty$) representation, given by $\boldsymbol{x}_{t}=\sum_{j=0}^{\infty}\boldsymbol{B}^{j}\boldsymbol{\zeta}_{t-j}$ or $\boldsymbol{z}=\boldsymbol{P}\boldsymbol{e}$, where
+$\boldsymbol{B}=\begin{bmatrix}\boldsymbol{A}_{1}&\boldsymbol{A}_{2}&\cdots&\boldsymbol{A}_{L-1}&\boldsymbol{A}_{L}\\ \boldsymbol{I}_{Q}&\boldsymbol{O}&\cdots&\boldsymbol{O}&\boldsymbol{O}\\ \boldsymbol{O}&\boldsymbol{I}_{Q}&\cdots&\boldsymbol{O}&\boldsymbol{O}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ \boldsymbol{O}&\boldsymbol{O}&\cdots&\boldsymbol{I}_{Q}&\boldsymbol{O}\end{bmatrix},\quad\boldsymbol{P}=\begin{bmatrix}\boldsymbol{I}_{QL}&\boldsymbol{B}&\boldsymbol{B}^{2}&\cdots&\boldsymbol{B}^{T-1}&\cdots\\ \boldsymbol{O}&\boldsymbol{I}_{QL}&\boldsymbol{B}&\cdots&\boldsymbol{B}^{T-2}&\cdots\\ \boldsymbol{O}&\boldsymbol{O}&\boldsymbol{I}_{QL}&\cdots&\boldsymbol{B}^{T-3}&\cdots\\ \vdots&\vdots&\vdots&\ddots&\vdots&\cdots\\ \boldsymbol{O}&\boldsymbol{O}&\boldsymbol{O}&\cdots&\boldsymbol{I}_{QL}&\cdots\end{bmatrix}$, (D.3)
+$\boldsymbol{\zeta}_{t}=(\boldsymbol{e}_{t-1}^{\top},\boldsymbol{0}^{\top},\dots,\boldsymbol{0}^{\top})^{\top}\in\mathbb{R}^{QL}$, $\boldsymbol{e}=(\boldsymbol{\zeta}_{{T}}^{\top},\boldsymbol{\zeta}_{{T}-1}^{\top},\dots,\boldsymbol{\zeta}_{1}^{\top},\dots)^{\top}$ and $\boldsymbol{z}=(\boldsymbol{x}_{{T}}^{\top},\boldsymbol{x}_{{T}-1}^{\top},\dots,$ $\boldsymbol{x}_{1}^{\top})^{\top}\in\mathbb{R}^{TQL}$.
+By Assumption 5, we have $\boldsymbol{e}_{t}=\boldsymbol{\Sigma}_{e}^{1/2}\boldsymbol{\xi}_{t}$ and thus $\boldsymbol{e}=(\boldsymbol{I}_{\infty}\otimes\boldsymbol{\Sigma}_{e}^{1/2})\bar{\boldsymbol{\xi}}$, where $\bar{\boldsymbol{\xi}}=(\bar{\boldsymbol{\xi}}_{T}^{\top},\bar{\boldsymbol{\xi}}_{T-1}^{\top},\dots,\bar{\boldsymbol{\xi}}_{1}^{\top},\dots)^{\top}$ and $\bar{\boldsymbol{\xi}}_{t}=(\boldsymbol{\xi}_{t-1}^{\top},\boldsymbol{0}^{\top},\dots,\boldsymbol{0}^{\top})^{\top}\in\mathbb{R}^{QL}$.
+Similar to the proof of Lemma 3, we consider a random process $\boldsymbol{W}(\boldsymbol{\Delta})=\left\lVert\boldsymbol{\Delta}\right\rVert_{T}^{2}-\mathbb{E}\left\lVert\boldsymbol{\Delta}\right\rVert_{T}^{2}$ and apply the conclusion in Lemma 7.
+We first verify the mixed tail condition for $\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma)$.
+For any $\boldsymbol{\Delta}_{a},\boldsymbol{\Delta}_{b}\in\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma)\subset\mathcal{W}_{\mathrm{AR}}$, we have
+$\left|\boldsymbol{W}(\boldsymbol{\Delta}_{a})-\boldsymbol{W}(\boldsymbol{\Delta}_{b})\right|=\left|\left(\left\lVert\boldsymbol{\Delta}_{a}\right\rVert_{T}^{2}-\left\lVert\boldsymbol{\Delta}_{b}\right\rVert_{T}^{2}\right)-\mathbb{E}\left(\left\lVert\boldsymbol{\Delta}_{a}\right\rVert_{T}^{2}-\left\lVert\boldsymbol{\Delta}_{b}\right\rVert_{T}^{2}\right)\right|.$
+Note that
+$\left\lVert\boldsymbol{\Delta}_{a}\right\rVert_{T}^{2}$ $=\frac{1}{T}\sum_{t=1}^{T}\left\lVert\boldsymbol{\Delta}_{a}\boldsymbol{x}_{t}\right\rVert_{2}^{2}=\frac{1}{T}\boldsymbol{z}^{\top}\left(\boldsymbol{I}_{T}\otimes(\boldsymbol{\Delta}_{a}^{\top}\boldsymbol{\Delta}_{a})\right)\boldsymbol{z}$
+$=\frac{1}{T}\bar{\boldsymbol{\xi}}^{\top}(\boldsymbol{I}\otimes\boldsymbol{\Sigma}_{e}^{1/2})^{\top}\boldsymbol{P}^{\top}\left(\boldsymbol{I}_{T}\otimes(\boldsymbol{\Delta}_{a}^{\top}\boldsymbol{\Delta}_{a})\right)\boldsymbol{P}(\boldsymbol{I}\otimes\boldsymbol{\Sigma}_{e}^{1/2})\bar{\boldsymbol{\xi}}$
+$=\bar{\boldsymbol{\xi}}^{\top}\left[\frac{1}{T}(\boldsymbol{I}\otimes\boldsymbol{\Sigma}_{e}^{1/2})^{\top}\boldsymbol{P}^{\top}\left(\boldsymbol{I}_{T}\otimes(\boldsymbol{\Delta}_{a}^{\top}\boldsymbol{\Delta}_{a})\right)\boldsymbol{P}(\boldsymbol{I}\otimes\boldsymbol{\Sigma}_{e}^{1/2})\right]\bar{\boldsymbol{\xi}}.$
+Similar transformation can be applied to $\left\lVert\boldsymbol{\Delta}_{b}\right\rVert_{T}^{2}$. So we have
+$\left\lVert\boldsymbol{\Delta}_{a}\right\rVert_{T}^{2}-\left\lVert\boldsymbol{\Delta}_{b}\right\rVert_{T}^{2}=\bar{\boldsymbol{\xi}}^{\top}\left[\frac{1}{T}(\boldsymbol{I}\otimes\boldsymbol{\Sigma}_{e}^{1/2})^{\top}\boldsymbol{P}^{\top}\left(\boldsymbol{I}_{T}\otimes(\boldsymbol{\Delta}_{a}^{\top}\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b}^{\top}\boldsymbol{\Delta}_{b})\right)\boldsymbol{P}(\boldsymbol{I}\otimes\boldsymbol{\Sigma}_{e}^{1/2})\right]\bar{\boldsymbol{\xi}}.$
+Denote $\boldsymbol{M}=\boldsymbol{\Delta}_{a}^{\top}\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b}^{\top}\boldsymbol{\Delta}_{b}$.
+Given that $\boldsymbol{\Delta}_{a},\boldsymbol{\Delta}_{b}\in\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma)$ and hence that $\left\lVert\boldsymbol{\Delta}_{a}\right\rVert_{\mathrm{op}}\leq\left\lVert\boldsymbol{\Delta}_{a}\right\rVert_{\mathrm{F}}=1$, $\left\lVert\boldsymbol{\Delta}_{b}\right\rVert_{\mathrm{op}}\leq\left\lVert\boldsymbol{\Delta}_{b}\right\rVert_{\mathrm{F}}=1$, we have $\left\lVert\boldsymbol{M}\right\rVert_{\mathrm{F}}=\left\lVert(\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b})^{\top}\boldsymbol{\Delta}_{a}+\boldsymbol{\Delta}_{b}^{\top}(\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b})\right\rVert_{\mathrm{F}}\leq 2\left\lVert\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b}\right\rVert_{\mathrm{F}}$.
+Further denote $\boldsymbol{Q}=(\boldsymbol{I}\otimes\boldsymbol{\Sigma}_{e}^{1/2})^{\top}\boldsymbol{P}^{\top}(\boldsymbol{I}_{T}\otimes\boldsymbol{M})\boldsymbol{P}(\boldsymbol{I}\otimes\boldsymbol{\Sigma}_{e}^{1/2})/T$.
+Then, we can upper bound the operator norm and Frobenius norm of $\boldsymbol{Q}$ by
+$\left\lVert\boldsymbol{Q}\right\rVert_{{\mathrm{op}}}$ $=\left\lVert\frac{1}{T}(\boldsymbol{I}\otimes\boldsymbol{\Sigma}_{e}^{1/2})^{\top}\boldsymbol{P}^{\top}\left(\boldsymbol{I}_{T}\otimes\boldsymbol{M}\right)\boldsymbol{P}(\boldsymbol{I}\otimes\boldsymbol{\Sigma}_{e}^{1/2})\right\rVert_{{\mathrm{op}}}$
+$\leq\frac{1}{{T}}\left\lVert\boldsymbol{I}\otimes\boldsymbol{\Sigma}_{e}^{1/2}\right\rVert_{{\mathrm{op}}}^{2}\left\lVert\boldsymbol{P}\right\rVert_{{\mathrm{op}}}^{2}\left\lVert\boldsymbol{I}_{T}\otimes\boldsymbol{M}\right\rVert_{{\mathrm{op}}}$
+$\leq\frac{1}{{T}}C_{e}\lambda_{\mathrm{max}}(\boldsymbol{P}\boldsymbol{P}^{\top})\left\lVert\boldsymbol{M}\right\rVert_{\mathrm{F}}\leq\frac{2}{{T}}C_{e}\lambda_{\mathrm{max}}(\boldsymbol{P}\boldsymbol{P}^{\top})\left\lVert\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b}\right\rVert_{\mathrm{F}}.$
+$\left\lVert\boldsymbol{Q}\right\rVert_{\mathrm{F}}^{2}$ $=\left\lVert\frac{1}{T}(\boldsymbol{I}\otimes\boldsymbol{\Sigma}_{e}^{1/2})^{\top}\boldsymbol{P}^{\top}\left(\boldsymbol{I}_{T}\otimes\boldsymbol{M}\right)\boldsymbol{P}(\boldsymbol{I}\otimes\boldsymbol{\Sigma}_{e}^{1/2})\right\rVert_{\mathrm{F}}^{2}$
+$\leq\frac{1}{T^{2}}\left\lVert\boldsymbol{I}\otimes\boldsymbol{\Sigma}_{e}^{1/2}\right\rVert_{{\mathrm{op}}}^{4}\left\lVert\boldsymbol{P}\right\rVert_{{\mathrm{op}}}^{4}\left\lVert\boldsymbol{I}_{T}\otimes\boldsymbol{M}\right\rVert_{\mathrm{F}}^{2}$
+$=\frac{1}{T^{2}}C_{e}^{2}\lambda_{\mathrm{max}}(\boldsymbol{P}\boldsymbol{P}^{\top})^{2}\left\lVert\boldsymbol{I}_{T}\right\rVert_{\mathrm{F}}^{2}\left\lVert\boldsymbol{M}\right\rVert_{\mathrm{F}}^{2}\leq\frac{4}{{T}}C_{e}^{2}\lambda_{\mathrm{max}}(\boldsymbol{P}\boldsymbol{P}^{\top})^{2}\left\lVert\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b}\right\rVert_{\mathrm{F}}^{2}.$
+With the upper bound of $\left\lVert\boldsymbol{Q}\right\rVert_{{\mathrm{op}}}$ and $\left\lVert\boldsymbol{Q}\right\rVert_{\mathrm{F}}^{2}$, the tail probability for the difference process holds for any $u>0$ as follows:
+$\mathbb{P}\left\{\left|\boldsymbol{W}(\boldsymbol{\Delta}_{a})-\boldsymbol{W}(\boldsymbol{\Delta}_{b})\right|\geq u\right\}$
+$=\penalty 10000\$$\mathbb{P}\left\{\left|\bar{\boldsymbol{\xi}}^{\top}\boldsymbol{Q}\bar{\boldsymbol{\xi}}-\mathbb{E}\left(\bar{\boldsymbol{\xi}}^{\top}\boldsymbol{Q}\bar{\boldsymbol{\xi}}\right)\right|\geq u\right\}$
+$=\penalty 10000\$$\mathbb{P}\left\{\left|\tilde{\boldsymbol{\xi}}^{\top}\tilde{\boldsymbol{Q}}\tilde{\boldsymbol{\xi}}-\mathbb{E}\left(\tilde{\boldsymbol{\xi}}^{\top}\tilde{\boldsymbol{Q}}\tilde{\boldsymbol{\xi}}\right)\right|\geq u\right\}$
+$\leq\penalty 10000\$$2\exp\left\{-c\min\left(\frac{u}{\kappa^{2}\|\tilde{\boldsymbol{Q}}\|_{\mathrm{op}}},\frac{u^{2}}{\kappa^{4}\|\tilde{\boldsymbol{Q}}\|_{\mathrm{F}}^{2}}\right)\right\}$
+$\leq\penalty 10000\$$2\exp\left\{-c\min\left(\frac{u}{2\kappa^{2}\kappa_{U,B}\left\lVert\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b}\right\rVert_{\mathrm{F}}/{T}},\frac{u^{2}}{(2\kappa^{2}\kappa_{U,B}\left\lVert\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b}\right\rVert_{\mathrm{F}}/\sqrt{{T}})^{2}}\right)\right\},$
+where $\kappa_{U,B}=C_{e}\lambda_{\mathrm{max}}(\boldsymbol{P}\boldsymbol{P}^{\top})$, $\tilde{\boldsymbol{\xi}}=(\boldsymbol{\xi}_{T}^{\top},\dots,\boldsymbol{\xi}_{1}^{\top})^{\top}$ contains all non-zero entries of $\boldsymbol{\xi}$, and $\tilde{\boldsymbol{Q}}$ is the corresponding sub-matrix of $\boldsymbol{Q}$.
+The first inequality is the Hanson-Wright inequality (Vershynin (2018), Theorem 6.2.1) as the entries of $\tilde{\boldsymbol{\xi}}$ are independent $\kappa^{2}$-sub-Gaussian random variables, and the second inequality is due to $\|\tilde{\boldsymbol{Q}}\|_{\mathrm{F}}\leq\left\lVert\boldsymbol{Q}\right\rVert_{\mathrm{F}}$ and $\|\tilde{\boldsymbol{Q}}\|_{\mathrm{op}}\leq\left\lVert\boldsymbol{Q}\right\rVert_{\mathrm{op}}$.
+Also, by the resulting mixed tail for $\boldsymbol{W}(\boldsymbol{\Delta})$, we define $d_{1}(\boldsymbol{\Delta}_{a},\boldsymbol{\Delta}_{b})=2c^{\prime}\kappa^{2}\kappa_{U,B}$ $\left\lVert\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b}\right\rVert_{\mathrm{F}}/T$, and $d_{2}(\boldsymbol{\Delta}_{a},\boldsymbol{\Delta}_{b})=2c^{\prime\prime}\kappa^{2}\kappa_{U,B}\left\lVert\boldsymbol{\Delta}_{i}-\boldsymbol{\Delta}_{j}\right\rVert_{\mathrm{F}}/\sqrt{T}$.
+By Theorem 8.6.1 of Vershynin (2018) and Eq.
+(46) and Lemma 2.7 of Melnyk and Banerjee (2016), we have
+$\begin{split}&\gamma_{2}(\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma),d_{2})=\left(2c^{\prime\prime}\kappa^{2}\kappa_{U,B}/\sqrt{T}\right)\gamma_{2}(\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma),\left\lVert\cdot\right\rVert_{\mathrm{F}})\leq\left(c_{2}\kappa^{2}\kappa_{U,B}/\sqrt{T}\right)\cdot\omega(\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma)),\\ &\gamma_{1}(\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma),d_{1})=\left(2c^{\prime}\kappa^{2}\kappa_{U,B}/{T}\right)\cdot\gamma_{1}(\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma),\left\lVert\cdot\right\rVert_{\mathrm{F}})\leq\left(c_{1}\kappa^{2}\kappa_{U,B}/T\right)\cdot\omega^{2}(\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma)),\\ &\mathrm{Diam}_{2}(\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma))=4c^{\prime\prime}\kappa^{2}\kappa_{U,B}/\sqrt{T},\quad\mathrm{Diam}_{1}(\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma))=4c^{\prime}\kappa^{2}\kappa_{U,B}/T.\end{split}$
+On the other hand, by Dudley’s inequality and Lemma 1(c), we have
+$\omega(\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma))$ $\leq C_{1}\int_{0}^{2}\sqrt{\log\mathcal{N}\left(\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma),\left\lVert\cdot\right\rVert_{\mathrm{F}},\epsilon\right)}\mathrm{d}\epsilon$
+$\leq C_{2}\int_{0}^{2}\left[d_{\mathrm{AR}}\log\left(\frac{24N\sqrt{[(K_{y}s_{\max})\wedge(LK_{x}r_{\max})]LK_{x}K_{y}(r_{\max}\vee s_{\max})}}{\gamma\epsilon}\right)\right]^{1/2}\mathrm{d}\epsilon$
+$\leq C_{3}d_{\mathrm{AR}}^{1/2}\left[\log\left(\frac{N\sqrt{[(K_{y}s_{\max})\wedge(LK_{x}r_{\max})]LK_{x}K_{y}(r_{\max}\vee s_{\max})}}{\gamma}\right)\right]^{1/2}$
+$=C_{3}\sqrt{d_{\mathrm{AR}}d_{\mathrm{AR}}^{\prime}(\gamma)}$ (D.4)
+where $d_{\mathrm{AR}},d_{\mathrm{AR}}^{\prime}(\gamma)$ are defined in (C.6).
+Let $u=\omega(\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma))$. Substituting all above terms into Lemma 7, we have
+$\mathbb{P}\left\{\sup_{\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma)}\left|\left\lVert\boldsymbol{\Delta}\right\rVert_{T}^{2}-\mathbb{E}\left\lVert\boldsymbol{\Delta}\right\rVert_{T}^{2}\right|=\sup_{\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma)}\left|\boldsymbol{W}(\boldsymbol{\Delta})\right|>c_{3}\kappa^{2}\kappa_{U,B}\left(\sqrt{\frac{d_{\mathrm{AR}}d_{\mathrm{AR}}^{\prime}(\gamma)}{T}}+\frac{d_{\mathrm{AR}}d_{\mathrm{AR}}^{\prime}(\gamma)}{T}\right)\right\}$
+$\hskip 330.0pt\leq 2\exp\left\{-cd_{\mathrm{AR}}d_{\mathrm{AR}}^{\prime}(\gamma)\right\}.$
+We now turn to the upper and lower bound for $\mathbb{E}\left\lVert\boldsymbol{\Delta}\right\rVert_{T}^{2}$. According to Basu and Michailidis (2015) (Proposition 2.3 and Equation (2.6)), we have
+$\lambda_{\mathrm{min}}\{\mathbb{E}(\boldsymbol{x}_{t}\boldsymbol{x}_{t}^{\top})\}\geq\lambda_{\mathrm{min}}(\boldsymbol{\Sigma}_{e})/\mu_{\mathrm{max}}(\mbox{$\mathscr{A}$}^{\ast})\quad\text{and}\quad\lambda_{\mathrm{max}}\{\mathbb{E}(\boldsymbol{x}_{t}\boldsymbol{x}_{t}^{\top})\}\leq\lambda_{\mathrm{max}}(\boldsymbol{\Sigma}_{e})/\mu_{\mathrm{min}}(\mbox{$\mathscr{A}$}^{\ast}).$
+We denote $\kappa_{L,A}=\lambda_{\mathrm{min}}(\boldsymbol{\Sigma}_{e})/\mu_{\mathrm{max}}(\mbox{$\mathscr{A}$}^{\ast})$ and $\kappa_{U,A}=\lambda_{\mathrm{max}}(\boldsymbol{\Sigma}_{e})/\mu_{\mathrm{min}}(\mbox{$\mathscr{A}$}^{\ast})$.
+Note that
+$\mathbb{E}\left\lVert\boldsymbol{\Delta}\right\rVert_{T}^{2}=$ $\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\mathrm{vec}(\boldsymbol{\Delta}^{\top})^{\top}(\boldsymbol{I}_{Q}\otimes\boldsymbol{x}_{t}\boldsymbol{x}_{t}^{\top})\mathrm{vec}(\boldsymbol{\Delta}^{\top})\right]=\mathrm{vec}(\boldsymbol{\Delta}^{\top})^{\top}(\boldsymbol{I}_{Q}\otimes\mathbb{E}[\boldsymbol{x}_{t}\boldsymbol{x}_{t}^{\top}])\mathrm{vec}(\boldsymbol{\Delta}^{\top}),$
+and by Rayleigh’s quotient, $\kappa_{L,A}\leq\mathbb{E}\left\lVert\boldsymbol{\Delta}\right\rVert_{T}^{2}\leq\kappa_{U,A}$ for all $\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma)$.
+Note that
+$\inf_{\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma)}\left\lVert\boldsymbol{\Delta}\right\rVert_{T}^{2}\geq\inf_{\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma)}\mathbb{E}\left\lVert\boldsymbol{\Delta}\right\rVert_{T}^{2}-\sup_{\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma)}\left|\left\lVert\boldsymbol{\Delta}\right\rVert_{T}^{2}-\mathbb{E}\left\lVert\boldsymbol{\Delta}\right\rVert_{T}^{2}\right|\geq\kappa_{L,A}-\sup_{\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{AR}}^{\prime}}\left|\left\lVert\boldsymbol{\Delta}\right\rVert_{T}^{2}-\mathbb{E}\left\lVert\boldsymbol{\Delta}\right\rVert_{T}^{2}\right|$
+and vice versa for the supremum. We get the uniform concentration inequality
+$\mathbb{P}\Biggl\{$ $\kappa_{L,A}-c_{3}\kappa^{2}\kappa_{U,B}\left(\sqrt{\frac{d_{\mathrm{AR}}d_{\mathrm{AR}}^{\prime}(\gamma)}{T}}+\frac{d_{\mathrm{AR}}d_{\mathrm{AR}}^{\prime}(\gamma)}{T}\right)\leq\inf_{\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma)}\left\lVert\boldsymbol{\Delta}\right\rVert_{T}^{2}\leq\sup_{\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma)}\left\lVert\boldsymbol{\Delta}\right\rVert_{T}^{2}$
+$\leq\kappa_{U,A}+c_{3}\kappa^{2}\kappa_{U,B}\left(\sqrt{\frac{d_{\mathrm{AR}}d_{\mathrm{AR}}^{\prime}(\gamma)}{T}}+\frac{d_{\mathrm{AR}}d_{\mathrm{AR}}^{\prime}(\gamma)}{T}\right)\Biggl\}\geq 1-2\exp\left\{-cd_{\mathrm{AR}}d_{\mathrm{AR}}^{\prime}(\gamma)\right\}.$
+By letting $T\gtrsim\max\{\kappa^{4}\kappa_{U,B}^{2}/\kappa_{L,A}^{2},\kappa^{2}\kappa_{U,B}/\kappa_{L,A}\}d_{\mathrm{AR}}d_{\mathrm{AR}}^{\prime}(\gamma)$, we have
+$\frac{\kappa_{L,A}}{16}\leq\inf_{\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma)}\left\lVert\boldsymbol{\Delta}\right\rVert_{T}^{2}\leq\sup_{\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma)}\left\lVert\boldsymbol{\Delta}\right\rVert_{T}^{2}\leq 4\kappa_{U,A}$
+with probability at least $1-2\exp\left\{-cd_{\mathrm{AR}}d_{\mathrm{AR}}^{\prime}(\gamma)\right\}$.
+[…]
+Lemma 5 (Autoregression deviation bound).
+Suppose that Assumptions 5, 1, 6 hold. Then, when $T\gtrsim d_{\mathrm{AR}}d_{\mathrm{AR}}^{\prime}(\gamma)$,
+$\sup_{\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma)}\frac{1}{T}\sum_{t=1}^{T}\langle\boldsymbol{e}_{t},\boldsymbol{\Delta}\boldsymbol{x}_{t}\rangle\leq C\kappa^{2}\kappa_{U,B}^{\prime}\sqrt{d_{\mathrm{AR}}d_{\mathrm{AR}}^{\prime}(\gamma)/T}$
+with probability at least $1-\exp\{-cd_{\mathrm{AR}}d_{\mathrm{AR}}^{\prime}(\gamma)\}$, where $d_{\mathrm{AR}}$, $d_{\mathrm{AR}}^{\prime}(\gamma)$, $\kappa_{U,B}^{\prime}$ are defined in (C.6), and $C,c$ are some positive constants.
+Proof.
+Similar to the proof of Lemma 3, we consider the random process $\boldsymbol{W}(\boldsymbol{\Delta})\vcentcolon=\frac{1}{T}\sum_{t=1}^{T}\langle\boldsymbol{e}_{t},\boldsymbol{\Delta}\boldsymbol{x}_{t}\rangle$ and apply the conclusion in Lemma 7.
+We first need to verify the mixed tail condition.
+For any $\boldsymbol{\Delta}_{a},\boldsymbol{\Delta}_{b}\in\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma)$ and $\lambda>0$, we have
+$\mathbb{P}\left\{\left|\boldsymbol{W}(\boldsymbol{\Delta}_{a})-\boldsymbol{W}(\boldsymbol{\Delta}_{b})\right|\geq u\right\}$ $\leq\exp(-\lambda u)\mathbb{E}\left[\exp\left(\lambda\sum_{t=1}^{T}\langle\boldsymbol{e}_{t},(\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b})\boldsymbol{x}_{t}/T\rangle\right)\right]$
+$=\exp(-\lambda u)\mathbb{E}\left[\exp\left(\lambda\sum_{t=1}^{T}\langle\boldsymbol{\xi}_{t},\boldsymbol{\Sigma}_{e}^{1/2}(\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b})\boldsymbol{x}_{t}/T\rangle\right)\right]$
+$\leq\exp(-\lambda u)\mathbb{E}\left[\exp\left(C^{\prime}\lambda^{2}\kappa^{2}\sum_{t=1}^{T}\|\boldsymbol{\Sigma}_{e}^{1/2}(\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b})\boldsymbol{x}_{t}/T\|_{2}^{2}\right)\right]$
+$\leq\exp(-\lambda u)\mathbb{E}\left[\exp\left(C^{\prime}\lambda^{2}\kappa^{2}C_{e}\sum_{t=1}^{T}\|(\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b})\boldsymbol{x}_{t}/T\|_{2}^{2}\right)\right],$
+where the first inequality is a direct conclusion from Markov’s inequality, the equality follows from Assumption 5, and the second inequality uses Lemma 8.
+Denote $\boldsymbol{M}=(\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b})/T$. By a similar argument to Lemma 4, we can rewrite
+$\sum_{t=1}^{T}\|\boldsymbol{M}\boldsymbol{x}_{t}\|_{2}^{2}=\bar{\boldsymbol{\xi}}^{\top}\left((\boldsymbol{I}\otimes\boldsymbol{\Sigma}_{e}^{1/2})^{\top}\boldsymbol{P}^{\top}(\boldsymbol{I}_{T}\otimes\boldsymbol{M})^{\top}(\boldsymbol{I}_{T}\otimes\boldsymbol{M})\boldsymbol{P}(\boldsymbol{I}\otimes\boldsymbol{\Sigma}_{e}^{1/2})\right)\bar{\boldsymbol{\xi}}.$
+Let $\boldsymbol{Q}_{1}=(\boldsymbol{I}_{T}\otimes\boldsymbol{M})\boldsymbol{P}(\boldsymbol{I}\otimes\boldsymbol{\Sigma}_{e}^{1/2})$, we can upper bound $\left\lVert\boldsymbol{Q}_{1}\right\rVert_{{\mathrm{op}}}$ and $\left\lVert\boldsymbol{Q}_{1}\right\rVert_{\mathrm{F}}$ by
+$\left\lVert\boldsymbol{Q}_{1}\right\rVert_{{\mathrm{op}}}$ $=\left\lVert(\boldsymbol{I}_{T}\otimes\boldsymbol{M})\boldsymbol{P}(\boldsymbol{I}\otimes\boldsymbol{\Sigma}_{e}^{1/2})\right\rVert_{{\mathrm{op}}}\leq\left\lVert\boldsymbol{I}_{T}\otimes\boldsymbol{M}\right\rVert_{{\mathrm{op}}}\left\lVert\boldsymbol{P}\right\rVert_{{\mathrm{op}}}\left\lVert\boldsymbol{I}\otimes\boldsymbol{\Sigma}_{e}^{1/2}\right\rVert_{{\mathrm{op}}}$
+$\leq\frac{1}{{T}}C_{e}^{1/2}\lambda_{\mathrm{max}}^{1/2}(\boldsymbol{P}\boldsymbol{P}^{\top})\left\lVert\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b}\right\rVert_{\mathrm{F}}.$
+$\left\lVert\boldsymbol{Q}_{1}\right\rVert_{\mathrm{F}}^{2}$ $=\left\lVert(\boldsymbol{I}_{T}\otimes\boldsymbol{M})\boldsymbol{P}(\boldsymbol{I}\otimes\boldsymbol{\Sigma}_{e}^{1/2})\right\rVert_{\mathrm{F}}^{2}\leq\left\lVert\boldsymbol{I}_{T}\otimes\boldsymbol{M}\right\rVert_{\mathrm{F}}^{2}\left\lVert\boldsymbol{P}\right\rVert_{{\mathrm{op}}}^{2}\left\lVert\boldsymbol{I}\otimes\boldsymbol{\Sigma}_{e}^{1/2}\right\rVert_{{\mathrm{op}}}^{2}$
+$=C_{e}\lambda_{\mathrm{max}}(\boldsymbol{P}\boldsymbol{P}^{\top})\left\lVert\boldsymbol{I}_{T}\right\rVert_{\mathrm{F}}^{2}\left\lVert\boldsymbol{M}\right\rVert_{\mathrm{F}}^{2}\leq\frac{1}{{T}}C_{e}\lambda_{\mathrm{max}}(\boldsymbol{P}\boldsymbol{P}^{\top})\left\lVert\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b}\right\rVert_{\mathrm{F}}^{2}.$
+Note that $\sum_{t=1}^{T}\|\boldsymbol{M}\boldsymbol{x}_{t}\|_{2}^{2}=\left\lVert\boldsymbol{Q}_{1}\bar{\boldsymbol{\xi}}\right\rVert_{2}^{2}$, and therefore for any $0<\lambda\leq c_{1}/(\kappa^{2}C_{e}^{1/2}\left\lVert\boldsymbol{Q}_{1}\right\rVert_{{\mathrm{op}}})$,
+$\mathbb{P}\left\{\left|\boldsymbol{W}(\boldsymbol{\Delta}_{a})-\boldsymbol{W}(\boldsymbol{\Delta}_{b})\right|\geq u\right\}$ $\leq\exp(-\lambda u)\mathbb{E}\left[\exp\left(C^{\prime}\lambda^{2}\kappa^{2}C_{e}\sum_{t=1}^{T}\|\boldsymbol{M}\boldsymbol{x}_{t}\|_{2}^{2}\right)\right]$
+$=\exp(-\lambda u)\mathbb{E}\left[\exp\left(C^{\prime}\lambda^{2}\kappa^{2}C_{e}\left\lVert\boldsymbol{Q}_{1}\bar{\boldsymbol{\xi}}\right\rVert_{2}^{2}\right)\right]$
+$\leq\exp\left(C\lambda^{2}\kappa^{4}C_{e}\left\lVert\boldsymbol{Q}_{1}\right\rVert_{\mathrm{F}}^{2}-\lambda u\right).$
+The last inequality is due to Vershynin (2018), Exercise 6.2.6 and ignoring all zero entries. By choosing $\lambda$ similar to a sub-exponential tail, we have
+$\mathbb{P}\left\{\left|\boldsymbol{W}(\boldsymbol{\Delta}_{a})-\boldsymbol{W}(\boldsymbol{\Delta}_{b})\right|\geq u\right\}$
+$\leq$ $\exp\left\{-c^{\prime}\min\left(\frac{u}{\kappa^{2}C_{e}^{1/2}\left\lVert\boldsymbol{Q}_{1}\right\rVert_{{\mathrm{op}}}},\frac{u^{2}}{\kappa^{4}C_{e}\left\lVert\boldsymbol{Q}_{1}\right\rVert_{\mathrm{F}}^{2}}\right)\right\}$
+$\leq$ $\exp\left\{-\min\left(\frac{u}{c^{\prime\prime}\kappa^{2}\kappa_{U,B}^{\prime}\left\lVert\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b}\right\rVert_{\mathrm{F}}/T},\frac{u^{2}}{\left(c^{\prime\prime}\kappa^{2}\kappa_{U,B}^{\prime}\left\lVert\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b}\right\rVert_{\mathrm{F}}/\sqrt{T}\right)^{2}}\right)\right\},$
+where $\kappa_{U,B}^{\prime}=C_{e}\lambda_{\mathrm{max}}^{1/2}(\boldsymbol{P}\boldsymbol{P}^{\top})$, and the second inequality follows from the upper bounds of $\left\lVert\boldsymbol{Q}_{1}\right\rVert_{{\mathrm{op}}}$ and $\left\lVert\boldsymbol{Q}_{1}\right\rVert_{\mathrm{F}}$.
+By the resulting mixed tail bound for $\boldsymbol{W}(\boldsymbol{\Delta})$, we let $d_{1}(\boldsymbol{\Delta}_{a},\boldsymbol{\Delta}_{b})=c^{\prime\prime}\kappa^{2}\kappa_{U,B}^{\prime}\left\lVert\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b}\right\rVert_{\mathrm{F}}/T$ and $d_{2}(\boldsymbol{\Delta}_{a},\boldsymbol{\Delta}_{b})=c^{\prime\prime}\kappa^{2}\kappa_{U,B}^{\prime}\left\lVert\boldsymbol{\Delta}_{a}-\boldsymbol{\Delta}_{b}\right\rVert_{\mathrm{F}}/\sqrt{T}$.
+By Theorem 8.6.1 of Vershynin (2018) and Eq.
+(46) and Lemma 2.7 of Melnyk and Banerjee (2016), we have
+$\gamma_{2}(\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma),d_{2})=\left(c^{\prime\prime}\kappa^{2}\kappa_{U,B}^{\prime}/\sqrt{T}\right)\cdot\gamma_{2}(\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma),\left\lVert\cdot\right\rVert_{\mathrm{F}})\leq\left(c_{2}\kappa^{2}\kappa_{U,B}^{\prime}/\sqrt{T}\right)\cdot\omega(\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma)),$
+$\gamma_{1}(\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma),d_{1})=\left(c^{\prime\prime}\kappa^{2}\kappa_{U,B}^{\prime}/T\right)\cdot\gamma_{1}(\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma),\left\lVert\cdot\right\rVert_{\mathrm{F}})\leq\left(c_{1}\kappa^{2}\kappa_{U,B}^{\prime}/T\right)\cdot\omega^{2}(\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma)),$
+$\mathrm{Diam}^{\prime}_{2}({\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma)})=2c^{\prime\prime}\kappa^{2}\kappa_{U,B}^{\prime}/\sqrt{T},\quad\mathrm{Diam}_{1}^{\prime}(\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma))=2c^{\prime\prime}\kappa^{2}\kappa_{U,B}^{\prime}/T.$
+By (D), we have $\omega(\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma))\lesssim d_{\mathrm{AR}}^{1/2}d_{\mathrm{AR}}^{\prime 1/2}$, with $d_{\mathrm{AR}}=2(K_{x}r_{\max}^{2}+K_{y}s_{\max}^{2})\sum_{n=1}^{N}q_{n}+2LK_{x}K_{y}r_{\max}s_{\max}$ and $d_{\mathrm{AR}}^{\prime}=\log\left(N\sqrt{[(K_{y}s_{\max})\wedge(LK_{x}r_{\max})]LK_{x}K_{y}(r_{\max}\vee s_{\max})}\right)$.
+Let $u=\omega(\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma))$. Substituting all above terms into Lemma 7, we have when $T\gtrsim d_{\mathrm{AR}}d_{\mathrm{AR}}^{\prime}(\gamma)$,
+$\sup_{\boldsymbol{\Delta}\in\mathcal{W}_{\mathrm{AR}}^{\prime}(\gamma)}\boldsymbol{W}(\boldsymbol{\Delta})\leq C\kappa^{2}\kappa_{U,B}^{\prime}\sqrt{d_{\mathrm{AR}}d_{\mathrm{AR}}^{\prime}(\gamma)/T}$
+[…]
+Lemma 6 (Spectral properties of VAR($L$) process).
+Suppose that Assumptions 5 and 1 hold. With the notations of Lemma 4, we have
+  • (a)
+[…]
+  • (b)
+[…]
+Proof.
+(a) Given an arbitrary vector $\boldsymbol{e}=(\boldsymbol{e}_{T}^{\top},\boldsymbol{e}_{T-1}^{\top},\dots,\boldsymbol{e}_{1}^{\top},\dots)^{\top}$ where $\boldsymbol{e}_{t}\in\mathbb{R}^{QL}$, the $n$-th block of $\boldsymbol{P}^{\top}\boldsymbol{e}$ is $\sum_{j=-\infty}^{n}(\boldsymbol{B}^{\top})^{n-j}\boldsymbol{e}_{j}$ for $n=0,1,\dots,T$.
+Then, we observe that $\boldsymbol{P}^{\top}\boldsymbol{e}$ is part of $\{(\boldsymbol{B}^{\top})^{k}\}\ast\{\boldsymbol{e}_{k}\}=\{\sum_{j=-\infty}^{k}(\boldsymbol{B}^{\top})^{k-j}\boldsymbol{e}_{j}\}_{k=0}^{\infty}$, where the asterisk denotes the convolution operator.
+The Fourier transform of $\{(\boldsymbol{B}^{\top})^{k}\}\ast\{\boldsymbol{e}_{k}\}$ is given by $\mathcal{F}(\{(\boldsymbol{B}^{\top})^{k}\}\ast\{\boldsymbol{e}_{k}\})(\theta)=(\sum_{k=0}^{\infty}(\boldsymbol{B}^{\top})^{k}e^{-ik\theta})\mathcal{F}(\{\boldsymbol{e}_{k}\})(\theta)=(\boldsymbol{I}_{Q}-\boldsymbol{B}^{\top}e^{-i\theta})^{-1}\mathcal{F}(\{\boldsymbol{e}_{k}\})(\theta)$.
+Hence, by Parseval’s theorem, we have
+$\left\lVert\boldsymbol{P}^{\top}\boldsymbol{e}\right\rVert_{2}^{2}$ $\leq\left\lVert\{(\boldsymbol{B}^{\top})^{k}\}\ast\{\boldsymbol{e}_{k}\}\right\rVert_{2}^{2}$
+$\leq\frac{1}{2\pi}\int_{0}^{2\pi}\left\lVert\mathcal{F}(\{(\boldsymbol{B}^{\top})^{k}\})(\theta)\right\rVert_{{\mathrm{op}}}^{2}\left\lVert\mathcal{F}(\{\boldsymbol{e}_{k}\})(\theta)\right\rVert_{2}^{2}\mathrm{d}\theta$
+$=\frac{1}{2\pi}\int_{0}^{2\pi}\lambda_{\max}\left((\boldsymbol{I}_{QL}-\boldsymbol{B}^{\top}e^{-i\theta})^{-1}(\boldsymbol{I}_{QL}-\boldsymbol{B}e^{i\theta})^{-1}\right)\left\lVert\mathcal{F}(\{\boldsymbol{e}_{k}\})(\theta)\right\rVert_{2}^{2}\mathrm{d}\theta$
+$=\frac{1}{2\pi}\int_{0}^{2\pi}(1/\lambda_{\min}\left((\boldsymbol{I}_{QL}-\boldsymbol{B}^{\top}e^{-i\theta})(\boldsymbol{I}_{QL}-\boldsymbol{B}e^{i\theta})\right))\left\lVert\mathcal{F}(\{\boldsymbol{e}_{k}\})(\theta)\right\rVert_{2}^{2}\mathrm{d}\theta$
+[…]
+[…]
+[…]
+(b) By relevant definitions,
+$\mu_{\mathrm{max}}(\mbox{$\mathscr{A}$}^{\ast})=\max_{|z|=1}\lambda_{\mathrm{max}}(\bar{\mbox{$\mathscr{A}$}}(z)\mbox{$\mathscr{A}$}(z))$ $\geq\left\lVert\boldsymbol{I}_{Q}-\boldsymbol{A}_{1}-\cdots-\boldsymbol{A}_{L}\right\rVert_{\mathrm{op}}^{2}$
+$=\sup_{\left\lVert\boldsymbol{u}\right\rVert=1}\left\lVert\boldsymbol{u}-\boldsymbol{A}_{1}\boldsymbol{u}-\cdots-\boldsymbol{A}_{L}\boldsymbol{u}\right\rVert_{2}^{2}$
+$=\sup_{\left\lVert\boldsymbol{u}\right\rVert=1}\left\lVert(\boldsymbol{I}_{QL}-\boldsymbol{B})(\boldsymbol{u}^{\top},\boldsymbol{u}^{\top},\dots,\boldsymbol{u}^{\top})^{\top}\right\rVert_{2}^{2}$
+$\geq L\lambda_{\min}((\boldsymbol{I}_{QL}-\boldsymbol{B})^{\top}(\boldsymbol{I}_{QL}-\boldsymbol{B}))$
+[…]
+concluding the proof. ∎
+Lemma 7 (Chaining, Theorem 3.5 in Dirksen (2015)).
+Suppose that the random process $\boldsymbol{W}(\boldsymbol{\Delta})_{\boldsymbol{\Delta}\in\mathcal{W}^{\prime}}$ has a mixed tail in the sense that for all $\boldsymbol{\Delta}_{a},\boldsymbol{\Delta}_{b}\in\mathcal{W}^{\prime}$ and $u>0$,
+$\mathbb{P}\left\{\left|\boldsymbol{W}(\boldsymbol{\Delta}_{a})-\boldsymbol{W}(\boldsymbol{\Delta}_{b})\right|\geq u\right\}\leq 2\exp\left\{-\min\left(\frac{u^{2}}{d_{2}(\boldsymbol{\Delta}_{a},\boldsymbol{\Delta}_{b})^{2}},\frac{u}{d_{1}(\boldsymbol{\Delta}_{a},\boldsymbol{\Delta}_{b})}\right)\right\}$
+where $d_{1}$ and $d_{2}$ are two semi-metrics on $\mathcal{W}^{\prime}$.
+Assume that $\mathcal{W}^{\prime}$ is bounded with respect to both $d_{1}$ and $d_{2}$, and that there exists a zero point $\boldsymbol{\Delta}_{0}\in\mathcal{W}^{\prime}$ such that $\boldsymbol{W}(\boldsymbol{\Delta}_{0})=0$ with probability 1.
+Then, there exists a constant $C>0$ such that for any $u\geq 1$,
+$\mathbb{P}\left\{\sup_{\boldsymbol{\Delta}\in\mathcal{W}^{\prime}}\left|\boldsymbol{W}(\boldsymbol{\Delta})\right|>C\left(\gamma_{2}(\mathcal{W}^{\prime},d_{2})+\gamma_{1}(\mathcal{W}^{\prime},d_{1})+u\mathrm{Diam}_{2}({\mathcal{W}^{\prime}})+u^{2}\mathrm{Diam}_{1}(\mathcal{W}^{\prime})\right)\right\}$
+$\leq 2\exp(-u^{2}),$
+where $\mathrm{Diam}_{2}({\mathcal{W}^{\prime}})$ and $\mathrm{Diam}_{1}(\mathcal{W}^{\prime})$ are the diameters of $\mathcal{W}^{\prime}$ with respect to the semi-metric $d_{2}$ and $d_{1}$, respectively.
+Proof.
+Substituting $\sqrt{u}d_{2}(\boldsymbol{\Delta}_{a},\boldsymbol{\Delta}_{b})+ud_{1}(\boldsymbol{\Delta}_{a},\boldsymbol{\Delta}_{b})$ for $u$ in the condition of this lemma, the mixed tail condition set out in Equation (3.8) of Dirksen (2015) is satisfied.
+Then the conclusion follows directly from Theorem 3.5 of Dirksen (2015) by setting the initial point $\boldsymbol{\Delta}_{0}$ to be the one such that $\boldsymbol{W}(\boldsymbol{\Delta}_{0})=0$ with probability 1.
+∎
+Lemma 8 (Martingale inequality).
+Let $\{\mathcal{F}_{t},t\in\mathbb{Z}\}$ be a filtration.
+Suppose that $\{\boldsymbol{w}_{t}\}$ and $\{\boldsymbol{e}_{t}\}$ are processes taking values in $\mathbb{R}^{n}$, and for each integer $t$, $\boldsymbol{w}_{t}$ is $\mathcal{F}_{t-1}$-measurable, $\boldsymbol{e}_{t}$ is $\mathcal{F}_{t}$-measurable, and $\boldsymbol{e}_{t}\mid\mathcal{F}_{t-1}$ is mean-zero and $\kappa^{2}$-sub-Gaussian distributed.
+Then for any $\lambda>0$, we have
+$\mathbb{E}\left[\exp\left(\lambda\sum_{t=1}^{T}\langle\boldsymbol{w}_{t},\boldsymbol{e}_{t}\rangle\right)\right]\leq\mathbb{E}\left[\exp\left(C\lambda^{2}\kappa^{2}\sum_{t=1}^{T}\|\boldsymbol{w}_{t}\|_{2}^{2}\right)\right],$
+where $C$ is some positive constants.
+Proof.
+By the tower rule, we have
+$\mathbb{E}\left[\exp\left(\lambda\sum_{t=1}^{T}\langle\boldsymbol{w}_{t},\boldsymbol{e}_{t}\rangle\right)\right]$ $=\mathbb{E}\left[\mathbb{E}\left[\exp\left(\lambda\sum_{t=1}^{T}\langle\boldsymbol{w}_{t},\boldsymbol{e}_{t}\rangle\right)\mid\mathcal{F}_{T-1}\right]\right]$
+$=\mathbb{E}\left[\exp\left(\lambda\sum_{t=1}^{T-1}\langle\boldsymbol{w}_{t},\boldsymbol{e}_{t}\rangle\right)\mathbb{E}\left[\exp\left(\lambda\langle\boldsymbol{w}_{T},\boldsymbol{e}_{T}\rangle\right)\mid\mathcal{F}_{T-1}\right]\right]$
+$\leq\mathbb{E}\left[\exp\left(c^{\prime}\lambda^{2}\kappa^{2}\|\boldsymbol{w}_{T}\|_{2}^{2}\right)\mathbb{E}\left[\exp\left(\lambda\sum_{t=1}^{T-1}\langle\boldsymbol{w}_{t},\boldsymbol{e}_{t}\rangle\right)\right]\right],$
+where the above inequality follows from the the fact that $\langle\boldsymbol{w}_{T},\boldsymbol{e}_{T}\rangle\mid\mathcal{F}_{T-1}$ is mean-zero and $\kappa^{2}\|\boldsymbol{w}_{T}\|^{2}_{2}$-sub-Gaussian, since $\boldsymbol{e}_{T}\mid\mathcal{F}_{t-1}$ is mean-zero and $\kappa^{2}$-sub-Gaussian.
+Thus we have $\mathbb{E}\left[\exp\left(\lambda\langle\boldsymbol{w}_{T},\boldsymbol{e}_{T}\rangle\right)\mid\mathcal{F}_{T-1}\right]\leq\exp\left(c^{\prime}\lambda^{2}\kappa^{2}\|\boldsymbol{w}_{T}\|_{2}^{2}\right)$.
+Then induction on $t$ gives the conclusion.
+∎
+Appendix E Setups and Results of the Third Simulation Studies in Section 5
+The third simulation compares our feature extracting methods with two traditional ones, hierarchical factor models (HFM) and higher-order factor models (HOFM), under the new model-designing framework.
+We consider two scenarios for the coefficient tensor $\mbox{$\mathscr{A}$}_{\mathrm{AR}}$: (i) it follows HOFM with both $\boldsymbol{\Lambda}_{x}$ and $\boldsymbol{\Lambda}_{y}$ being the matrix product of three different block diagonal matrices; and (ii) it follows our method with two action orders $\{(1,2,3),(2,1,3)\}$ for both $\boldsymbol{\Lambda}_{x}$ and $\boldsymbol{\Lambda}_{y}$.
+For both scenarios, we set the ranks as $r_{n,k}=4$ and $s_{n,k^{\prime}}=3$ for all $n$, $k$ and $k^{\prime}$, and all component and coefficient matrices are generated as in the second experiment; specifically, the construction of the loading matrices $\boldsymbol{\Lambda}_{y}$ and $\boldsymbol{\Lambda}_{x}$ for both our method and HOFM is given as follows:
+  • For our model, we set $(K_{y},K_{x})=(2,2)$, with action sets $\mathcal{P}_{y}(3,2)=\mathcal{P}_{x}(3,2)=\{(1,2,3),(2,1,3)\}$, and generate loading components $\{\{\boldsymbol{G}_{n,k}\}_{n=1}^{3}\}_{k=1}^{3}$ and $\{\{\boldsymbol{H}_{n,k^{\prime}}\}_{n=1}^{3}\}_{k^{\prime}=1}^{3}$ by applying QR decomposition to random matrices with i.i.d.
+standard normal entries.
+Then $\boldsymbol{\Lambda}_{y}$ and $\boldsymbol{\Lambda}_{x}$ for our model are computed according to (2.10).
+  • For HOFM, since it perform feature extraction in a sequential manner and only considers one action order $(1,2,3)$, we simply generate the loading components $\{\boldsymbol{G}_{n}\}_{n=1}^{3}$ such that $\boldsymbol{G}_{1}=\text{diag}(\boldsymbol{G}_{1}^{(1)},\boldsymbol{G}_{1}^{(2)},\ldots,\boldsymbol{G}_{1}^{(p_{2}p_{3})})$, and $\boldsymbol{G}_{2}=\text{diag}(\boldsymbol{G}_{2}^{(1)},\boldsymbol{G}_{2}^{(2)},\ldots,\boldsymbol{G}_{2}^{(p_{3})})$.
+Here, $\boldsymbol{G}_{1}^{(i)}\in\mathbb{R}^{r\times p_{1}}$, $\boldsymbol{G}_{2}^{(j)}\in\mathbb{R}^{r\times p_{2}r}$, and $\boldsymbol{G}_{3}\in\mathbb{R}^{r\times p_{3}r}$ are orthogonal matrices for $1\leq i\leq q_{2}q_{3}$ and $1\leq j\leq q_{3}$, generated by QR decomposition of random matrices with i.i.d.
+standard normal entries.
+Similarly, we generate the loading components $\{\boldsymbol{H}_{n}\}_{n=1}^{3}$ and compute $\boldsymbol{\Lambda}_{y}=\boldsymbol{H}_{1}\boldsymbol{H}_{2}\boldsymbol{H}_{3}$ and $\boldsymbol{\Lambda}_{x}=\boldsymbol{G}_{1}\boldsymbol{G}_{2}\boldsymbol{G}_{3}$ accordingly.
+[…]
+Figure 5: Average rolling forecast MSEs for HOFM, HFM, and our proposed method under our new supervised modeling framework.
+Algorithm 1 is employed to estimate our models, while an alternating least squares algorithm is used for models with HFM and HOFM being feature extracting tools.
+One-step ahead forecasting is used for the evaluation, and Figure 5 gives the MSEs, averaged over 500 replications.
+Our model consistently outperforms both HFM and HOFM, regardless of whether the data are generated by HOFM or our method.
+It perhaps is due to the fact that both HFM and HOFM involve large number of parameters.
+As a result, the efficiency of our model in handling hierarchical structures can be well demonstrated by capturing shared components within each mode under high-dimensional settings.
+Appendix F Additional Analysis for the Personality‑120 Data
+We first offer an interpretation of Item 58 for the first action order $\alpha_{1}=(3,2,1)$ as introduced in Section 6, following the way Item 46 is analyzed in the main text.
+For elderly females, the important questions include Question 3 and 9 of Neuroticism trait, as well as Question 4, 7, and 13 of Openness trait.
+These questions are summarized as follows: “Am afraid of many things” and “Often feel blue” exhibit positive loadings, signifying a tendency of anxiety and depression, while “Like to get lost in thought,” “Prefer variety to routine,” and “Do not like poetry” exhibit negative loadings, signifying a loss of imagination and adventurousness.
+These questions, despite being scattered across different traits, can be collectively interpreted as a concern of control and security among elderly females.
+Figure 6: Heatmaps of obliquely rotated estimated component matrices and the coefficient matrix: (a) $\widehat{\boldsymbol{G}}_{1,2}^{\top}$, (b) $\widehat{\boldsymbol{G}}_{2,2}^{\top}$, (c) $\widehat{\boldsymbol{G}}_{3,2}^{\top}$.
+In (c), A, C, E, N, and O stand for Agreeableness, Conscientiousness, Extraversion, Neuroticism, and Openness, respectively.
+For the second action order $\alpha_{2}=(1,3,2)$, we present the obliquely rotated component matrices for both predictors and responses to illustrate the interpretability of our proposed model.
+Figure 6 shows the heatmaps of the estimated component matrices $\widehat{\boldsymbol{G}}_{1,2}^{\top}$, $\widehat{\boldsymbol{G}}_{2,2}^{\top}$, and $\widehat{\boldsymbol{G}}_{3,2}^{\top}$, whose interpretation follows the same logic as Section 6.
+In Figure 6(a), Group 1–6 represent males and Group 7–12 females.
+Within each group, a larger index indicates a higher age, and therefore the elderly tend to have a different attitude than the younger, especially for females.